### v2

# Index theory for singular Lagrangian systems and Bessel-type differential operators

Abstract
The aim of the present manuscript is to develop an index theory for singular Lagrangian systems, with a particular focus on the important class of singular operators given by Bessel-type differential operators. The main motivation is to address several challenges posed by singular operators, which appear in a wide range of applications: celestial mechanics (for instance, perturbations in planetary motion), oscillatory systems with time-dependent forcing, electromagnetism (such as wave equations in nonuniform media), and quantum mechanics (notably certain Schrodinger equations with periodic potentials).
We pursue two principal objectives. First, we establish a spectral flow formula and a Morse Index Theorem for gap-continuous paths of singular Sturm–Liouville operators. By means of these index formulas, we construct a Morse index theory for a broad class of Bessel-type differential operators and apply it to a family of asymptotic solutions of the gravitational $N$-body problem.
Finally, our new index theory provides new insight into a phenomenon first observed by Rellich concerning the spectrum of one-parameter families of Sturm–Liouville operators with varying domains.
AMS Subject Classification: 34B24, 58J30, 53D12, 34B30.
Keywords: Singular Sturm-Liouville differential operators, Spectral flow, Maslov index, Asymptotic solutions of the N-body problem, Bessel operators, Rellich counterexample.
1 Introduction
In its classical form, the Morse Index Theorem asserts that the number of negative eigenvalues of the Hessian of the geodesic action functional at a critical point equals the number of conjugate points, counted with multiplicity, along the corresponding geodesic. Through standard spectral flow techniques, this integer can be expressed as the sum of the dimension of the intersections between the graph of the fundamental solution with the Lagrangian subspace corresponding to the boundary condition. For singular Sturm–Liouville operators, however, this classical framework fails, and the usual spectral flow identities no longer apply.
The present paper has two main objectives:
  1. (O1)
    To construct a Lagrangian intersection theory that provides the natural framework for both regular and (one-sided) singular Sturm–Liouville (SL) boundary value problems, and within this setting to establish a Spectral Flow Formula and a Morse Index Theorem with general Lagrangian boundary conditions.
  2. (O2)
    To develop a Morse Index Theory for Bessel-type differential operators and to apply to some classes of motions in the gravitational $n$-body problem.
As a direct consequence of this index theory, we rigorously capture the Rellich phenomenon—the “disappearance” of eigenvalues at -$\infty$—thereby refining the relation between spectral flow and the difference of Morse indices in this broader functional-analytic framework. This spectral effect arises when considering unbounded perturbations of closed, self-adjoint, lower-bounded Fredholm operators, which frequently appear in applications, especially in families of operators with varying domains. Finally, we compute the Morse index for a class of unbounded motions in the gravitational $N$-body problem.
The motivations for this theory are twofold:
  • (M1)
    To classify, via the Morse index, all possible motions in the gravitational $N$-body problem—whether partially or totally colliding, or unbounded.
  • (M2)
    To use this classification to a priori exclude collisions along solutions obtained through topological methods in the Calculus of Variations.
To the best of authors knowledge the index theory in this singular context has been investigated in [HS22]. In this paper the authors investigated a general class of linear Hamiltonian systems on intervals with at least one singular endpoint which can be limit-point, limit-circle, or limit-intermediate. In particular, they proved some renormalized oscillation results by using a Lagrangian intersection theory in the classical ${\mathbb{C}}^{2n}$ symplectic space.
1.1 Highlights and Main Results
After setting the stage in Section 2, we establish in Section 3 a spectral flow formula for a family $s\mapsto l_{s}$ of one-sided singular Sturm–Liouville operators defined by the differential expression
$l_{s}:=-\frac{d}{dt}\Bigl(P(t)\frac{d}{dt}+Q(t)\Bigr)+Q(t)^{T}\frac{d}{dt}+R(t)+C_{s}(t),\qquad t\in(a,b),$ (1.1)
with $a,b\in\overline{\mathbb{R}}$ and $s\in[0,1]$, where
$P^{-1}\in\mathscr{C}^{1}\big((a,b),\mathrm{S}_{n}(\mathbb{R})\big),\qquad Q\in\mathscr{C}^{1}\big((a,b),\mathrm{M}_{n}(\mathbb{R})\big),\qquad R\in\mathscr{C}^{0}\big((a,b),\mathrm{S}_{n}(\mathbb{R})\big),\\[3.0pt] \text{and }\ s\mapsto C_{s}\in\mathscr{C}^{0}_{b}\big((a,b),\mathrm{S}_{n}(\mathbb{R})\big)\ \text{is continuous},$ (1.2)
where $\mathscr{C}^{0}_{b}$ denotes the set of bounded continuous functions, and $C_{0}(t)=0$ for all $t\in(a,b)$.
Under the regularity assumptions in Equation (1.2), the formal differential expression $l_{s}$ in Equation (1.1) defines an eventually singular second-order Sturm–Liouville differential operator (abbreviated as SL-operator).
In the Hilbert space $H=L^{2}((a,b),\mathbb{R}^{n})$, let $L_{s}$ (respectively $L_{s}^{*}$) denote the minimal (respectively maximal) operator with domains $D(L_{s})$ and $D(L_{s}^{*})$ induced by $l_{s}$, and introduce the following hypotheses:
  • (H1)
    The endpoint $t=b$ is finite, and for every $s\in[0,1]$, it is a regular endpoint for the operator $\ell_{s}$, meaning that the paths $t\mapsto P(t)$, $t\mapsto Q(t)$, $t\mapsto R(t)$, and $t\mapsto C_{s}(t)$ are continuous up to $t=b$.
  • (H2)
    The path $0,1]\ni s\mapsto L_{s}\in\operatorname{\mathcal{CF}^{s}}(H)$ is gap-continuous and consists of symmetric Fredholm operators.
We note that $D(L_{s})$ is independent of $s$ and coincides with $D(L)$, where $L$ denotes the unperturbed Sturm–Liouville operator. On the space $D(L^{*})$, there is a well-defined skew-symmetric bilinear form given by
$\omega(x,y)=\langle x,L^{*}y\rangle-\langle L^{*}x,y\rangle,$ (1.3)
which, in general, is not symplectic since it is typically degenerate.
One of the main contributions of this first part of the paper is the explicit construction of a symplectic subspace $(W,\omega|_{W})$ through the $\ker L^{*}$, that is, the space of solutions to the Sturm–Liouville equation (see Section 4 for further details), and of a trace map that we now describe.
Under assumptions (H1) and (H2), the following crucial decomposition formula holds:
$D(L^{*})=D(L)\oplus W,$
where $(W,\omega|_{W})$ is a $(2k)$-dimensional symplectic space with $n\leq k\leq 2n$.
Example 1.1.
In the special case of regular endpoints, we have
$D(L_{s}^{*})=W^{2,2}([a,b],\mathbb{R}^{n}),\qquad D(L_{s})=W^{2,2}_{0}([a,b],\mathbb{R}^{n}),\\[3.0pt] \text{and}\qquad W\cong T^{*}\mathbb{R}^{n}\oplus T^{*}\mathbb{R}^{n},$ (1.4)
where the isomorphism between $W$ and $T^{*}\mathbb{R}^{n}\oplus T^{*}\mathbb{R}^{n}$ is defined by
$W\ni x\longmapsto\phi(x)=\big(x^{[1]}(a),x(a),x^{[1]}(b),x(b)\big),\\[3.0pt] \text{where }\quad x^{[1]}(c)=P(c)\,x^{\prime}(c)+Q(c)\,x(c),\quad\text{for }c=a,b.$ (1.5)
The vector $x^{[1]}(c)$ is referred to as the quasi-derivative at $c$.
Integrating by parts in Equation (1.3), we obtain
$\omega(f,g)=[f,g](b)-[f,g](a),\qquad\text{where}\\[3.0pt] [f,g](t):=\langle P(t)f^{\prime}(t)+Q(t)f(t),g(t)\rangle-\langle f(t),P(t)g^{\prime}(t)+Q(t)g(t)\rangle.$ (1.6)
Let $Z=\operatorname{span}\{z_{i}\mid 1\leq i\leq 2n\}$, where each $z_{i}$ is a smooth function satisfying $z_{i}=0$ on $(a,a+\varepsilon]$ and such that the family $\{(z_{i}^{[1]}(b),z_{i}(b))\}_{i=1}^{2n}$ forms a basis of $\mathbb{R}^{2n}$.
  • If $k=n$, on the symplectic space $(W,\omega|_{W})$ we define the trace map $\operatorname{\mathcal{T}}$ by
    $W\ni f\longmapsto\operatorname{\mathcal{T}}(f):=\big(\omega(f,z_{1}),\ldots,\omega(f,z_{2n})\big)=\big([f,z_{1}](b),\ldots,[f,z_{2n}](b)\big)\in\mathbb{R}^{2n}.$
    This case corresponds to the classical limit point case in the Weyl classification of the boundary conditions where no boundary condition is required at $a$.
  • If $k\neq n$, since $\omega|_{Z}$ is nondegenerate, we obtain the following symplectic decomposition:
    $W=Z^{\omega|_{W}}\oplus Z.$ (1.7)
    Setting
    $f,g](a^{+}):=\lim_{t\to a^{+}}[f,g](t),$
    we define the (linear) trace map $\operatorname{\mathcal{T}}:W\to\mathbb{R}^{2k-2n}\oplus\mathbb{R}^{2n}$ by
    $\operatorname{\mathcal{T}}(f):=\big(-[f,y_{1}](a^{+}),\ldots,-[f,y_{2k-2n}](a^{+}),[f,z_{1}](b),\ldots,[f,z_{2n}](b)\big),$
    where $y_{i}\in\ker L^{*}$ for $i=1,\ldots,2k-2n$. The existence of this subset $\widetilde{y}_{1},\ldots,\widetilde{y}_{2k-2n}$ of functions in $\ker L^{*}$ is proved at Lemma 4.4 for details. We refer to Remark 4.7 where is fully explained how to concretely select such a subset of functions. This case corresponds to the limit circle case or regular case in the Weyl classification of the boundary conditions.
Since $\omega|_{W}$ is nondegenerate, it follows that $\operatorname{\mathcal{T}}:Z^{\omega|_{W}}\oplus Z\to\mathbb{R}^{2k-2n}\oplus\mathbb{R}^{2n}$ is a bijection, and the decomposition is $\operatorname{\mathcal{T}}$-invariant. Moreover, the space
$\mathbb{R}^{2k-2n}\oplus\mathbb{R}^{2n}$
is the space of boundary data, where $\mathbb{R}^{2k-2n}$ (respectively $\mathbb{R}^{2n}$) contains the boundary data at $a$ (respectively $b$).
With all these ingredients, we are now in a position to construct a proper Lagrangian intersection theory and to prove a general spectral flow formula for Sturm–Liouville operators.
Let $\operatorname{\mathrm{Lag}}(W,\omega|_{W})$ denote the Lagrangian Grassmannian of $(W,\omega|_{W})$. For each $s$, consider the self-adjoint extension of $L_{s}$ defined by
$L_{s,\Lambda_{s}}:=L_{s}^{*}|_{D(L)\oplus\Lambda_{s}}.$
We observe that if $s\mapsto\Lambda_{s}\in\operatorname{\mathrm{Lag}}(W,\omega|_{W})$ is gap-continuous, then $s\mapsto L_{s,\Lambda_{s}}:=L_{s}^{*}|_{D(L)\oplus\Lambda_{s}}$ is a gap-continuous path of operators in $\operatorname{\mathcal{CF}^{sa}}(H)$. Setting $V_{s}=\ker L_{s}^{*}$, the main result of Section 3 reads as follows.
Theorem 1.
Under conditions (H1)(H2) and assuming that $s\mapsto\Lambda_{s}\in\operatorname{\mathrm{Lag}}(W,\omega|_{W})$ is gap-continuous, the following spectral flow formula holds:
$\operatorname{sf}(L_{s,\Lambda_{s}},s\in[0,1])=-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\operatorname{\mathcal{T}}(\Lambda_{s}),\operatorname{\mathcal{T}}(V_{s}),\rho^{\operatorname{\mathcal{T}}},s\in[0,1]),$
where $\rho^{\operatorname{\mathcal{T}}}$ denotes the push-forward symplectic form.
In the spectral flow formula stated in Theorem 1, the decomposition $D(L^{*})=D(L)\oplus W$ plays an essential role. It is worth mentioning that there exists a more abstract approach to proving a spectral flow formula. In a fairly general setting, this was developed by Furutani in [Fur04], by Booss-Bavnbek, Lesch, and Phillips in [BLP05], and, in the very general context of symplectic Banach spaces, by Booss-Bavnbek and Zhu in [BZ18]. In the aforementioned works, the intersection theory is established in the Gelfand–Robbin quotient space
$\beta(L):=D(L^{*})/D(L).$
By comparing the method developed in this manuscript with those presented in the above papers, our approach offers several advantages:
  1. 1.
    The definition of the Maslov index is concrete and expressed in terms of the paths $s\mapsto V_{s}$ and $s\mapsto\Lambda_{s}$, both of which are contained in $D(L^{*})$.
  2. 2.
    This approach can be extended to the case of varying domains $s\mapsto D(L_{s})$, a generalization that will be addressed in a forthcoming paper.
  3. 3.
    The space $W$ is defined explicitly by using a basis of $\ker L^{*}$. Moreover, since the entire construction depends only on the domains of $L^{*}$ and $L$, the definition of $W$ remains valid if $L$ is replaced by another operator $\widetilde{L}$ (not necessarily belonging to the same family) provided that $D(\widetilde{L})=D(L)$.
The third remark is particularly useful, since in most cases it is extremely difficult—if not impossible—to determine an explicit basis of $\ker L^{*}$. Replacing $L$ by an operator $\widetilde{L}$ having the same domain resolves this difficulty: if $\widetilde{L}^{*}$ is suitably chosen, then $\ker\widetilde{L}^{*}$ and hence $W$ can be computed explicitly. This idea has been applied, for instance, in Subsection 8.1.
In Subsection 4.2, we prove that the spectral flow formula stated in Theorem 1 generalizes several classical results available in the literature. In particular, it encompasses the following cases:
  1. 1.
    Regular Sturm–Liouville operators on bounded domains;
  2. 2.
    Sturm–Liouville operators on $\mathbb{R}$$^{+}$
  3. 3.
    Bessel-type operators on $(0,1]$ and $1,\infty)$.
The regular case is well understood, and several precise and comprehensive results are available in the literature; see, for instance, [CLM94, HP17, LZ00, RS95, Zhu06] and the references therein. In applications, this first case typically arises in the study of trajectories of Lagrangian systems defined on bounded time intervals.
The second case, however, appears naturally in the study of trajectories of Lagrangian systems defined on unbounded time intervals. This is the case, for example, of half-clinic solutions of Lagrangian systems, that is, solutions parametrized by the half-line. Recently, the first two authors established a general spectral flow formula in [HP17], which became a key ingredient in several subsequent results concerning the Morse index theorem in the gravitational $N$-body problem (cf. [HP17, HO16, HOY21] and references therein).
Using the above spectral flow formula, we prove in Section 5 the Morse index theorem for singular Sturm–Liouville operators. For the regular case or the heteroclinic and half-clinic cases, we refer the interested reader to [HWY20] , [HPWX20], [BCJLMS18] and [HP17].
We start with a single Sturm–Liouville operator $l$ defined on the function space $\mathscr{C}_{0}^{\infty}((a,b],\mathbb{R}^{n})$, where the endpoint $t=b$ is regular according to hypothesis (H1). In this context, by restricting the operator $l$ to the interval $(a,\sigma)$ with $\sigma\in(a,b]$, we obtain a one-parameter family of operators $\sigma\mapsto l_{(a,\sigma)}$. This construction is necessary for the application of the spectral flow formula.
For $\sigma\in(a,b)$, we consider $l_{(a,\sigma)}$ to be the one-sided singular Sturm–Liouville differential operator obtained by restricting $l$ to $\mathscr{C}_{0}^{\infty}((a,\sigma),{\mathbb{R}}^{n})$, and we denote by $L_{(a,\sigma)}$ the minimal operator associated with $l_{(a,\sigma)}$. We observe that $L_{(a,\sigma)}$ is singular only at the initial point $t=a$ (hence the final endpoint $\sigma$ is regular). We consider the Lagrangian subspaces
$\Lambda_{0}\in\operatorname{\mathrm{Lag}}({\mathbb{R}}^{2k-2n}),\qquad\Lambda_{D}:={\mathbb{R}}^{n}\times(0),\qquad\Lambda=\Lambda_{0}\oplus\Lambda_{D},$
where $\Lambda_{D}$ is the standard Dirichlet Lagrangian.
We introduce the following conditions:
    (H3) The quadratic form on $\mathscr{C}^{\infty}_{0}((a,b),{\mathbb{R}}^{n})$ associated with the SL-operator $l$ is bounded from below in $L^{2}$.
    (H4) The minimal operator $L_{b}\in\operatorname{\mathcal{CF}^{s}}(L^{2}((a,b),{\mathbb{R}}^{n}))$.
Remark 1.2.
It is worth observing that, under (H3), for every $\sigma\in(a,b]$, the quadratic form induced by $l_{(a,\sigma)}$ is bounded from below. In particular, there exists a well-defined self-adjoint Friedrichs extension $L_{(a,\sigma)}$ (cf. Section C for further details).
Given $\Lambda\in\operatorname{\mathrm{Lag}}({\mathbb{R}}^{2k-2n}\oplus{\mathbb{R}}^{2n},-\Omega\oplus\Omega)$, we set
$L_{(a,\sigma),\Lambda}=L_{(a,\sigma)}^{*}|_{D(L_{(a,\sigma)})\oplus\operatorname{\mathcal{T}}_{\sigma}^{-1}(\Lambda)},$
where $\operatorname{\mathcal{T}}_{\sigma}:W_{\sigma}\to{\mathbb{R}}^{2k-2n}\oplus{\mathbb{R}}^{2n}$ is the trace map defined on the subspace $W_{\sigma}$ associated with the operator $l_{(a,\sigma)}$ and constructed above.
Notation 1.3.
In what follows we denote by
  • $L_{(a,\sigma),F}$ the self-adjoint Friedrichs extension associated with $L_{(a,\sigma)}$;
  • $L_{(a,\sigma),\Lambda}$ the self-adjoint extension of $L_{(a,\sigma)}$ with boundary condition $\Lambda:=\Lambda_{0}\oplus\Lambda_{D}$.
We denote by $\mathrm{n_{-}}(\#)$ the Morse index of the self-adjoint operator #\#, namely, the maximal dimension of the negative spectral subspace of #\#.
Thus, the Morse index theorem reads as follows.
Theorem 2.
[Singular Morse Index Theorem] Assume that the SL-operator $l$ satisfies conditions (H1), (H3), and (H4), and that Dirichlet boundary conditions are imposed at $b$. Then the following equality holds:
$\mathrm{n_{-}}(L_{b,\Lambda})=\sum_{\sigma\in(a,b)}\dim\ker L_{(a,\sigma),\Lambda}.$
The Theorem 2 works even if $\mathrm{n_{-}}(L_{b})=+\infty$ and it is obtained by combining Theorem 5.1 and Proposition 6.3. It is worth noting that, even in this more general setting, Theorem 2 retains, at least formally, a structure closely resembling the classical result in the regular case (cf. [BOPW21, HP17] and the references therein). However, the traditional method of proving such a theorem—namely, by computing the local contributions to the spectral flow and to the Maslov index via crossing forms—fails in this singular context.
The next result relates the Morse index of the Friedrichs extension operator $L_{F}$ with the limit, as $\sigma$ tends to the endpoint, of the Friedrichs operators $L_{(a,\sigma),F}$ and $L_{(\sigma,b),F}$, under the sole assumption that the operator $L$ is bounded from below. In particular, in Theorem 3, neither Fredholmness nor the finiteness of the Morse index has been assumed.
Theorem 3.
Assume condition (H3). If $l$ is regular at $a$ and Dirichlet boundary conditions are imposed, then
$\mathrm{n_{-}}(L_{F})=\lim_{\sigma\to b^{-}}\mathrm{n_{-}}(L_{(a,\sigma),F})=\sum_{t\in(a,b)}\dim\big(\gamma(t)\Lambda_{D}\cap\Lambda_{D}\big),$
where $\gamma$ is the fundamental matrix solution satisfying $\gamma(a)=I$.
Analogously, if $l$ is regular at $b$ and Dirichlet boundary conditions are imposed, then
$\mathrm{n_{-}}(L_{F})=\lim_{\sigma\to a^{+}}\mathrm{n_{-}}(L_{(\sigma,b),F})=\sum_{t\in(a,b)}\dim\big(\gamma(t)\Lambda_{D}\cap\Lambda_{D}\big),$
where $\gamma$ is the fundamental matrix solution satisfying $\gamma(b)=I$.
In [Zet05, Theorem 10.12.1] the author collect the spectral properties and well-known index results for singular scalar Sturm-Liouville operators. The singular Morse Index Theorem can be seen as a multi-dimensional generalization of the aforementioned result.
The case of general Lagrangian boundary conditions is treated in detail in Section 6. In particular, we prove that even in this singular framework, the Morse index with general boundary conditions differs from the Dirichlet case by a correction term computed via the triple index of the Cauchy data space and the projections of the domains of $L_{\Lambda}$ and of the Friedrichs extension $L_{F}$.
Section 7 is devoted to study a rigorous quantification of a phenomenon first observed by Rellich in the case of singular Sturm-Liouville equations. In the abstract terms the stage is the following. If $T$ is a closed self-adjoint operator bounded from below, and $A$ is a relatively bounded symmetric perturbation with sufficiently small relative bound, then $S=T+A$ remains self-adjoint and bounded from below (see [Kat80, Chapter 5, Section 4, Remark 4.13]).
However, this property does not necessarily hold for unbounded perturbations. As Kato explains in [Kat80, Chapter 5, Section 4, Remark 4.13], based on Rellich’s example, when a sequence of self-adjoint operators $T_{n}$ converges in the gap topology to a self-adjoint operator bounded from below, the lower bounds of $T_{n}$ may diverge to -$\infty$, and eigenvalues may “disappear” at -$\infty$. Rellich’s scalar counterexample in the context of Sturm–Liouville operators illustrates this effect, although the precise number of disappearing eigenvalues remains an open problem. Section 7 provides a precise quantification of the multiplicity of eigenvalues that escape to -$\infty$. In the regular case, this phenomenon has been already studied at [HLWZ19, HLWZ22].
This phenomenon occurs, for instance, when the domains of the operators vary with the parameter. Consider $A\in\operatorname{\mathcal{CF}^{s}}(H)$ and a continuous path $s\mapsto\Lambda_{s}$ of Lagrangian subspaces of $(U,\rho)$. The corresponding path $s\mapsto A_{\Lambda_{s}}$ of self-adjoint extensions is gap-continuous, but the operators are no longer relatively bounded perturbations of a fixed operator. Such situations arise naturally in applications. The main result of Section 7 (Theorem 7.2) computes the total number of such limiting ghost eigenvalues that disappear to -$\infty$ through the Morse Index Theorem for singular Sturm–Liouville operators with general boundary conditions.
Section 8 is devoted to a detailed study of regular perturbations of the unperturbed Bessel operator
$l_{q}:=-\frac{d^{2}}{dt^{2}}+\frac{q}{t^{2}},$
acting on $C_{0}^{\infty}((0,1],{\mathbb{R}}^{n})$, which is particularly relevant in the analysis of the variational and index properties of unbounded motions in the gravitational $N$-body problem. Denoting by $L_{q}:D(L_{q})\subset H\to H$ the corresponding minimal operator, where $H=L^{2}((0,1],{\mathbb{R}}^{n})$, we investigate its functional-analytic and spectral properties, with particular emphasis on Fredholmness and the Morse index. The main results of this section are the following.
Theorem 4.
Let $R\in\mathscr{C}^{0}([0,1],\mathrm{S}(n))$, and let $L_{R}$ be the minimal operator induced by the Bessel-type operator
$l_{R}:=-\frac{d^{2}}{dt^{2}}+\frac{R(t)}{t^{2}},\qquad t\in(0,1].$
If R(t)>1/4R(t)>-1/4, then every self-adjoint extension of $L_{R}$ is a Fredholm operator.
The final main results of Section 8 concerns the Morse index properties of Bessel-type operators on $\mathscr{C}_{0}^{\infty}([1,+\infty),{\mathbb{R}}^{n})$ and on $\mathscr{C}_{0}^{\infty}((0,1],{\mathbb{R}}^{n})$, respectively.
Theorem 5.
Let $l_{R}$ be the operator defined above, acting on $\mathscr{C}_{0}^{\infty}([1,+\infty),{\mathbb{R}}^{n})$.
    Case 1. If $\liminf_{t\to+\infty}R(t)>-1/4$, then every self-adjoint extension of $l_{R}$ has finite Morse index.
    Case 2. If $\limsup_{t\to+\infty}R(t)<-1/4$, then every self-adjoint extension of $l_{R}$ has infinite Morse index.
Theorem 6.
Let $l_{R}$ be the operator defined above on $\mathscr{C}_{0}^{\infty}((0,1],{\mathbb{R}}^{n})$.
  • Case 1. If $\liminf_{t\to 0^{+}}R(t)>-1/4$, then any self-adjoint extension has finite Morse index.
  • Case 2. If $\limsup_{t\to 0^{+}}R(t)<-1/4$, then any self-adjoint extension has infinite Morse index.
Using these properties together with the developed index theory, we derive new results on the Morse index of totally colliding and unbounded motions in the $N$-body problem (Section 9). In particular, we provide both a new proof of the main results of [BHPT20, HOY21] and a significant generalization to more general self-adjoint boundary conditions. The main result of Section 9 is the following.
Theorem 7.
Let $q$ be a total collision or a parabolic solution at infinity having limiting normalized central configuration $a$, and set
$\bar{B}(a):=\frac{2}{9}\,\frac{M^{-1}D^{2}U(a)}{U(a)}.$
If
$\bar{B}(a)>-\dfrac{1}{4},$
then the Morse index $\mathrm{n_{-}}(q)$ of $q$ is finite, and the following identity holds:
$\mathrm{n_{-}}(q)=\sum_{t\in(a,b)}\dim\big(\gamma(t)\Lambda_{D}\cap\Lambda_{D}\big),$
where $\gamma$ is the fundamental matrix solution satisfying $\gamma(a)=I$.
If
$\bar{B}(a)<-\dfrac{1}{4},$
the Morse index is $\infty$.
If $q$ is hyperbolic, then the Morse index $\mathrm{n_{-}}(q)$ of $q$ is always finite.
The Appendix collects the most technical material. We first gather definitions, notation, and key results on symplectic indices and spectral flow used throughout the paper. Section C establishes the fundamental properties of Friedrichs extensions, which play a central role in our analysis, while Section D contains postponed technical proofs in order to streamline the main exposition.
List of Notation
  • ${,+}\overline{{\mathbb{R}}}:={\mathbb{R}}\cup\{$-$\infty$$\infty\}$. $J=[a,b]\subset\overline{{\mathbb{R}}}$, $I=[0,1]\subset{\mathbb{R}}$.
  • $\mathrm{M}(n)$ and $\mathrm{S}(n)$ the vector space of all (real) matrices and symmetric matrices, respectively.
  • << denotes inclusion between linear subspaces to distinguish by the set-theoretical inclusion $\subset$
  • $({\mathbb{R}}^{2n},\Omega)$ be the standard symplectic space. With abuse of notation we denote by the same symbol $\Omega$ the standard symplectic form of ${\mathbb{R}}^{2k-2n}$ for $n\leq k\leq 2n$.
  • $H=L^{2}((a,b),\mathbb{R}^{n})$.
  • $\operatorname{\mathcal{S}}(H)$ the set of all closed subspaces of $H$
  • $\mathcal{B}(H)$, $\mathcal{Cl}(H)$ (resp. $\mathcal{Cl}^{s}(H)$) be respectively the set of all linear bounded, closed (resp. closed and symmetric) densely defined operators in $H$. $\mathcal{B}^{sa}(H)$ denotes the set of all bounded and self-adjoint operators on $H$
  • $D(A)$ (resp. $D(A^{*})$) denotes the domain of $A$ (resp. of $A^{*}$, the adjoint of $A$). $\langle\cdot,\cdot\rangle^{G}$ graph inner product of $A$.
  • $p:D(A^{*})\to U$ denotes the quotient projection. Let $\omega$ denote the symplectic form induced by $A$ on $D(A^{*})$
  • $\operatorname{\mathrm{Lag}}$ denotes the space of all Lagrangian subspaces
  • $\operatorname{\mathcal{CF}^{s}}(H)$ denotes the set of all closed densely defined and symmetric Fredholm operators on $H$. $\operatorname{\mathcal{BF}^{sa}}(H)$ denotes the set of all bounded and self-adjoint Fredholm operators on $H$
  • For $s\in I$, let $l_{s}$ denotes a family of Sturm-Liouville operators. We denote by $L_{s}$ (resp. $L_{s}^{*}$) the minimal (resp. maximal) operator associated to $l_{s}$.
  • $\Lambda_{D}={\mathbb{R}}^{n}\times(0)$ denotes the Dirichlet and $\Lambda_{N}=(0)\times{\mathbb{R}}^{n}$ the Neumann Lagrangian subspaces of $({\mathbb{R}}^{2n},\Omega)$. With abuse of notation we denote by the same symbols $\Lambda_{D}$ (resp. $\Lambda_{N}$) the Dirichlet (resp. Neumann) Lagrangian of the symplectic space $({\mathbb{R}}^{2k-2n},\Omega)$, for $n\leq k\leq 2n$.
  • $\mathrm{n_{-}}(\#)$ denotes the Morse index (i.e. the maximal dimension of the negative spectral space) of the operator #\#
  • $\operatorname{sf}(\cdot)$: the spectral flow of a path of closed self-adjoint Fredholm operators;
  • $\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\cdot)$: the Maslov index as defined by Cappell–Lee–Miller .
  • $\iota(\cdot,\cdot,\cdot)$: the triple index, quantifying the difference of Morse indices between different boundary conditions.
List of Assumptions
    (H0) Given $A\in\mathcal{Cl}^{s}(H)$, we assume that $\ker A^{*}\cap D(A)=\{0\}$ (Unique Continuation Property)
    (H1) The instant $t=b$ is finite and for every $s\in[0,1]$ it is a regular endpoint for the operator $\ell_{s}$ meaning that the paths $t\mapsto P^{\prime}(t)$, $t\mapsto Q(t)$, $t\mapsto R(t)$ and $t\mapsto C_{s}(t)$ are continuous up to the instant $t=b$ (Regular endpoint)
    (H2) We assume that $0,1]\ni s\mapsto L^{*}_{s}\in\operatorname{\mathcal{CF}^{s}}(H)$ is a gap-continuous path of symmetric Fredholm operators (gap-continuity and Fredholmenss)
    (H3) We assume that the quadratic form $k$ in $H=L^{2}((a,b),{\mathbb{R}}^{n})$ associated to the SL-operator $l$ is bounded from below
    (H4) The minimal operator $L_{b}$ is a closed symmetric and Fredholm operator in $H$, so $L_{b}\in\operatorname{\mathcal{CF}^{s}}(H)$.
Acknowledgements
The second author is grateful to New York University Abu Dhabi for the opportunity to conduct his research in such an inspiring and supportive environment. He also extends his sincere thanks to the Institute’s administrative and technical staff for their invaluable assistance and the excellent working conditions they provide.
2 An abstract spectral flow formula
In this section, we introduce an abstract decomposition formula for an operator pencil of densely defined, closed self-adjoint Fredholm operators and we finally prove a spectral flow formula. In a different context of Cauchy data space, our main result of this section can be recast in the Gelfand-Robbin quotients. Our primary references are [Fur04, BZ18, Kat80] and the sources cited therein.
2.1 Symplectic Hilbert spaces and a decomposition formula
Let $(H,\langle\cdot,\cdot\rangle)$ be a real and separable Hilbert space with an inner product $\langle\cdot,\cdot\rangle$ and associated norm $\|\cdot\|$. We denote by $\operatorname{\mathcal{S}}(H)$ the set of all closed linear subspaces of $H$ and we set
$\operatorname{dist}(u,V):=\inf_{v\in V}\|u-v\|.$
Given $U,V\in\operatorname{\mathcal{S}}(H)$, we denote by $S_{U}$ the unit sphere of $U$ and we set
$\delta(U,V):=\begin{cases}\sup_{u\in S_{U}}\operatorname{dist}(u,V)&\textrm{ if }U\neq(0)\\[3.0pt] 0&\textrm{ if }U=(0)\end{cases}\quad\textrm{ and }\quad\widehat{\delta}(U,V):=\max\{\delta(U,V),\delta(V,U)\}.$ (2.1)
$\widehat{\delta}(U,V)$ is called the gap between $U$ and $V$ and defines a metric on $\operatorname{\mathcal{S}}(H)$ whose associated metric topology is called gap topology. Given $U\in\operatorname{\mathcal{S}}(H)$, there exists a unique orthogonal projection $P_{U}$ onto $U$ which is a bounded operator on $H$. Given $U,V\in\operatorname{\mathcal{S}}(H)$, we consider the norm of the difference of the corresponding orthogonal projectors:
$d_{G}(U,V):=\lVert P_{U}-P_{V}\rVert.$ (2.2)
It can be shown that metric topology induced by $d_{G}$ is equivalent to the gap topology.
Notation 2.1.
We denote by $\mathcal{Cl}(H)$ (resp. $\mathcal{Cl}^{s}(H)$) the set of all closed (resp. closed and symmetric) and densely defined operators [^1] Since the adjoint of a densely defined linear operator is closed then self-adjoint operators on $H$ are contained in $\mathcal{Cl}(H)$. and we denote by $\mathcal{G}({\#})$ the graph of the operator #\#.
We observe that the gap metric induces in a natural way a metric on $\mathcal{Cl}(H)$ (resp. $\mathcal{Cl}^{s}(H)$) given by
$d(T,S):=\widehat{\delta}(\mathcal{G}({T}),\mathcal{G}({S}))\qquad S,T\in\mathcal{Cl}(H).$
Let $A\in\mathcal{Cl}^{s}(H)$, $A^{*}\in\mathcal{Cl}^{s}(H)$ be its adjoint and we denote by $D(A)$ and $D(A^{*})$ the domain of $A$ and $A^{*}$, respectively. We define the graph inner product on $D(A^{*})$ as follows
$\langle x,y\rangle^{G}:=\langle x,y\rangle+\langle A^{*}x,A^{*}y\rangle\qquad x,y\in D(A^{*})$
and we observe that $D(A^{*})$ equipped with the graph norm becomes a Hilbert space and $D(A)$ is a closed subspace. (Cf. [Fur04, Example 2.2] for further details). Let us introduce on $D(A^{*})$ the following skew-symmetric bilinear form $\omega$ given by
$\omega\big(x,y\big)=\langle x,A^{*}y\rangle-\langle A^{*}\,x,y\rangle$ (2.3)
and we observe that $\omega$ is degenerate being $\ker\omega=D(A)$. Let $A\in\operatorname{\mathcal{CF}^{s}}(H)$ be a densely defined closed, symmetric and Fredholm operator on $H$. Under this assumption the following splitting formula holds
$D(A^{*})=D(A)\oplus U$
where $U$ is a finite dimensional (closed) subspace of $D(A^{*})$. It is worth observing that the restriction $\rho:=\omega|_{U}$ is non-degenerate. So the pair $(U,\rho)$ defines a symplectic space. We denote by $p:D(A^{*})\to U$ the canonical projection on the second factor.
For example we can choose $U$ be the orthogonal complement (wrt the graph inner product) $D(A)^{\perp_{G}}$ of $D(A)$ in $D(A^{*})$ with respect to the graph-inner product; namely with
[…]
From this characterization we get that $A^{*}$ restricted to $D(A)^{\perp_{G}}$ is an orthogonal transformation into itself and defines a complex structure on $D(A)^{\perp_{G}}$ since for every $x,y\in D(A)^{\perp_{G}}$, we have
$\omega(x,y)=\langle x,A^{*}(y)\rangle-\langle A^{*}(x),y\rangle=\langle-A^{*}(A^{*}(x)),A^{*}(y)\rangle-\langle A^{*}(x),y\rangle=\langle-A^{*}(x),y\rangle^{G}.$
This equality shows that the symplectic space $U$ together with $\rho$ defined above and together with the almost complex structure $A^{*}$ after the identification with the orthogonal complement $D(A)^{\perp_{G}}$ is a symplectic space with a compatible symplectic form, inner product and the almost complex structure.
2.2 An abstract spectral flow formula
Let $A,A^{*}\in\mathcal{Cl}^{s}(H)$ with domains respectively given by $D(A),D(A^{*})$, and we consider the splitting
$D(A^{*})=D(A)\oplus U$
Let $p:D(A^{*})\to U$ be projection map and we observe that the subspace $p(\ker A^{*})$ is an isotropic subspace of the symplectic space $U$. We refer to $p(\ker A^{*})$ as the Cauchy data space of the operator $A$. The next well-known result gives a characterization of the self-adjointness of an operator $A$ on $D$ in terms of the symplectic properties of $p(D)$.
Lemma 2.2.
Let $A\in\mathcal{Cl}^{s}(H)$ and let $D$ be denote a linear subspace such that $D(A)<D<D(A^{*})$. Then the restriction of the adjoint operator $A^{*}$ to the subspace $D$ is self-adjoint, if and only if, the subspace $p(D)$ is a Lagrangian subspace in $U$.
Proof.
We have $p(D)=D\cap U$. Let $A_{D}=A^{*}|_{D}:D\to H$. Since $A$ is symmetric then $D(A)<D(A^{*})$ and so $D(A_{D}^{*})<D(A^{*})$. Then we have
[…]
since $A_{D}^{*}x=A^{*}x$ for every $x\in D(A^{*})$ and $A_{D}y=Ay$ for every $y\in D(A)$. So, we get $D(A_{D}^{*})=D^{\omega}$. To conclude that $A_{D}^{*}=A_{D}$, by the above arguments, it’s equivalent to prove that $D^{\omega}=D$. Since $D(A)\subset D$ and $D(A)<D^{\omega}$, we have to show that
$D\cap U=D^{\omega}\cap U$
and being $(D\cap U)^{\rho}=D^{\omega}\cap U$, we just need to show that $(D\cap U)^{\rho}=D\cap U$. The conclusion follows by observing that $D\cap U=p(D)$.
Lemma 2.3.
Let $A\in\operatorname{\mathcal{CF}^{s}}(H)$. Then
  • (a)
    $U$ is a finite dimensional symplectic vector space
  • (b)
    AA has at least one self-adjoint Fredholm extension; that is, there exists a subspace $D$ (closed in the graph norm topology) such that $A_{D}:=A^{*}\big|_{D}$ is self-adjoint and Fredholm.
Proof.
[…]
$\mathrm{ind}\,A=\mathrm{ind}\,A^{*}-\mathrm{ind}\,j.$
Being $j$ an injection, it follows that $\mathrm{ind}\,j=-\dim\operatorname{coker}j=-\dim U$. So, we get that
$\dim U=\mathrm{ind}\,A^{*}-\mathrm{ind}\,A$
and since both integers on the (RHS) of the previous equality are finite, also their difference is so. This concludes the proof of (a).
For proving $(b)$ we start by observing that since by (a), the symplectic space $(U,\rho)$ is finite dimensional, then there exists a Lagrangian subspace $L<U$. We let $D=p^{-1}(L)$ and we observe that $p(D)=L$ since $\gamma$ is surjective. The conclusion directly follows by Lemma 2.2. ∎
Remark 2.4.
The minimal operators $A$ arising from the linear ordinary differential operators considered in this work are closed, self-adjoint, and Fredholm. As a consequence of Lemma 2.3, the associated quotient (or factor) spaces are finite-dimensional.
This stands in contrast to the case of elliptic partial differential operators, for which the minimal operator is generally not Fredholm, as it typically has an infinite-dimensional kernel.
We introduce the following Unique Continuation Property.
  • (H0) Let $A\in\mathcal{Cl}^{s}(H)$ and we assume that $\ker A^{*}\cap D(A)=(0)$.
Remark 2.5.
In the case of regular linear ordinary differential operators, the unique‐continuation property holds automatically. However, for general elliptic partial differential operators this property can fail, a phenomenon that is intimately connected to Carleman estimates and the Fefferman–Phong inequality.
The first basic consequence of (H0) is the following result.
Lemma 2.6.
Let $A\in\operatorname{\mathcal{CF}^{s}}(H)$ and we assume that (H0) holds. Then
$p\big(\ker A^{*}\big)$
is a Lagrangian subspace of $U$. Moreover this property doesn’t depend on the chosen decomposition.
Proof.
By (H0) we infer that $\dim p(\ker A^{*})=\dim\ker A^{*}$. We start showing that this property is independent on the chosen decomposition. Let $D(A^{*})=D(A)\oplus U^{\prime}$ be another decomposition. Since $D(A)=\ker\omega$, then the map $\Phi:(U^{\prime},\rho^{\prime})\mapsto(U,\rho)$ is a symplectic isomorphism. So, by this argument we get that the property for $p(\ker A^{*})$ to be Lagrangian is independent on the chosen decomposition. Without loss of generality, we can assume that $\ker A^{*}<U$ . We have
$(\ker A^{*})^{\rho}=(\ker A^{*})^{\omega}\cap U.$
We observe that
[…]
It follows that $(\ker A^{*})^{\omega}\cap U=(D(A)+\ker A^{*})\cap U=\ker A^{*}+(D(A)\cap U)=\ker A^{*}$. ∎
At this point, we introduce two technical results that will be used frequently throughout the manuscript.
Lemma 2.7.
Let $A\in\mathcal{Cl}^{s}(H)$ be a closed and symmetric operator on $H$ and let $B\in\mathcal{B}^{sa}(H)$ be a bounded self-adjoint operator on $H$. Then
  • $A+B\in\mathcal{Cl}^{s}(D(A),H)$
  • $D(A^{*})=D((A+B)^{*})$.
Proof.
By [Kat80, Chapter 4, Theorem 1.1] we get that $A+B$ is a closed operator. For proving the symmetry we start observing that the operator-matrix $\begin{pmatrix}0&C\\ D&0\end{pmatrix}$ defined on $H\times H$ is self-adjoint if and only if $D=C^{*}$. By [Kat80, Chapter 5, Theorem 4.3], we get that $\begin{pmatrix}0&A+B\\ A^{*}+B&0\end{pmatrix}$ is self-adjoint since it is a bounded self-adjoint perturbation of $\begin{pmatrix}0&A\\ A^{*}&0\end{pmatrix}$. Then we have
$(A+B)^{*}=A^{*}+B$
and by this equality the result then follows. ∎
Lemma 2.8 which is frequently used along the paper asserts that the perturbation of a closed operator on $H$ by a continuous family of bounded linear operators, parametrized by an open subset of $\mathbb{R}^{m}$, defines a gap-continuous family in $H$. In particular, a one-parameter family resulting from such a perturbation is gap-continuous.
Lemma 2.8.
Let $L\in\mathcal{Cl}(H)$ be a closed operator. Let $\Omega\subset{\mathbb{R}}^{m}$ be an open subset and we assume that the mapping
$\Omega\ni v\mapsto K_{v}\in\mathcal{B}(H)$
is continuous. Then the map $v\mapsto L+K_{v}$ is gap-continuous in $H$.
Proof.
Let $u\in\Omega$ and let $\varepsilon>0$. Since the mapping $v\mapsto K_{v}$ is continuous, up to choosing a smaller neighborhood $U$ of $u$ we can also assume that $\|K_{u}-K_{v}\|<\varepsilon$ for $v\in U$. By the assumption, there exists $M>0$ such that $\|K_{u}\|\leq M$ for $u\in U$. In the product space $H\times H$, let us consider the following norm:
$\|(p,q)\|_{H\times H}=\|p\|_{H}+\|q\|_{H}.$
Let $(w,Lw+K_{u}w)\in\operatorname{gr}(L+K_{u})$ such that $\|w\|+\|Lw+K_{u}w\|=1$. So, we get
$d(\operatorname{gr}(L+K_{u}),\operatorname{gr}(L+K_{v}))\leq\|(w,Lw+K_{u}w)-(w,Lw+K_{v}w)\|\leq\|K_{u}w-K_{v}w\|\\ \leq\|K_{u}-K_{v}\|\|w\|\leq\varepsilon.$ (2.6)
Moreover, we observe that
$\|(w,Lw+K_{v}w)\|\geq\|(w,Lw+K_{u}w)\|-\|K_{u}-K_{v}\|\|w\|\geq 1-\varepsilon.$
Summing together and by the true definition of the gap metric, we get:
$d(\operatorname{gr}(L+K_{v}),\operatorname{gr}(L+K_{u}))\leq(1-\varepsilon)^{-1}\|(w,Lw+K_{v}w)-(w,Lw+K_{u}w)\|\leq\varepsilon\,(1-\varepsilon)^{-1}.$
So, $v\mapsto L+K_{v}$ is gap-continuous concluding the proof. ∎
Let $0,1]\ni s\longmapsto A_{s}:=A+sI\in\operatorname{\mathcal{CF}^{s}}(H)$ be a family of closed symmetric and Fredholm operators. Then, by Lemma 2.7, we get that
$D(A_{s})=D(A)\quad\textrm{ and }\quad D(A^{*}_{s})=D(A^{*})\qquad s\in[0,1].$ (2.7)
As a direct consequence of Equation (2.7), is that both the factor space $U_{s}$ and the associated form $\rho_{s}$ induced by $A_{s}$ are independent on $s$ and so, they coincide with the factor space of $(U,\rho)$. Moreover, by invoking Lemma 2.8, we get that the path $s\mapsto A_{s}$ is gap-continuous.
Remark 2.9.
This choice greatly simplifies the construction of the entire index theory by permitting us to work in a fixed symplectic space. The situation becomes markedly more complicated if $D(A_{s}^{*})$—and hence the quotient space UU—depends on $s$, or if $\beta$ remains ss‑independent while the symplectic form $\omega$ varies with $s$. The latter scenario arises, for instance, when the operator path exhibits a dependence on $s$ in its zero‑order term.
Lemma 2.10.
We let $V_{s}:=\ker A^{*}_{s}$ and we assume that condition (H0) holds. Then the map $s\mapsto p(V_{s})$ is gap-continuous in $\operatorname{\mathrm{Lag}}(U,\rho)$.
Proof.
By Corollary D.2, the map $s\mapsto V_{s}$ is gap continuous in $\operatorname{\mathcal{S}}(D(A^{*}))$ with respect to the graph norm. We have
$p(V_{s})=(V_{s}+D(A))\cap U.$
By assumption (H0), we get that $V_{s}\cap D(A)=(0)$. So, by Lemma D.1, $V_{s}+D(A)$ is gap continuous in $\operatorname{\mathcal{S}}(D(A^{*}))$. Since $V_{s}+D(A)+U=D(A^{*})$, by Lemma D.1, also $p(V_{s})$ is gap-continuous in $D(A^{*})$ and then it is gap-continuous in $\operatorname{\mathrm{Lag}}(U,\rho)$. ∎
The next result shows that the Lagrangian path $s\mapsto V_{s}$ is a negative curve meaning that each crossing instant is nondegenerate and gives a negative contribution to the Maslov index.
Lemma 2.11.
The path $0,1]\ni s\longmapsto p(V_{s})\in\operatorname{\mathrm{Lag}}(U,\rho)$ is negative.
Proof.
We assume that $s_{0}\in[0,1]$ is a crossing instant. Let $x(s)\in p(V_{s})$ such that $\lim_{s\to s_{0}}x(s)=x$ and we set $z(s)=p^{-1}(x(s))$. By a direct calculation, it follows that
$\omega(x,x(s))=\langle z(s_{0}),A_{s_{0}}^{*}\,z(s)\rangle-\langle A_{s_{0}}^{*}\,z(s_{0}),z(s)\rangle=\langle z(s_{0}),A_{s_{0}}^{*}z(s)\rangle\\ =\langle z(s_{0}),A_{s}^{*}z(s)\rangle+\langle z(s_{0}),(s_{0}-s)z(s)\rangle=(s_{0}-s)\langle z(s_{0}),z(s)\rangle$ (2.8)
since $z(s)\in\ker A_{s}^{*}$ and $z(s_{0})\in\ker A_{s_{0}}^{*}$. So, we get that
$\Gamma:V_{s_{0}}\ni x\longmapsto\left.\dfrac{d}{ds}\right|_{s=s_{0}}\omega(x,x(s))=-\langle z(s_{0}),z(s_{0})\rangle=-\|z(s_{0})\|^{2}$
which is a negative definite quadratic form. This concludes the proof. ∎
Summing up all the previous arguments we get the following spectral flow formula for an operator pencil in $\operatorname{\mathcal{CF}^{s}}(H)$.
Proposition 2.12.
Let $s\mapsto A_{s}=A+sI\in\operatorname{\mathcal{CF}^{s}}(H)$ and we assume that assumption (H0) holds. Given $L\in\operatorname{\mathrm{Lag}}(U,\rho)$, let us consider the self-adjoint extension $A_{L}$. Then we get
$\operatorname{sf}(A_{L}+sI,s\in[0,1])=-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(L,p(V_{s}),\rho,s\in[0,1]).$ (2.9)
Remark 2.13.
Some remarks are in order. First, by Lemma 2.7, the symplectic space $(U_{s},\rho_{s})=(U,\rho)$ is independent of $s$. By Lemma 2.3, since $A\in\operatorname{\mathcal{CF}^{s}}(H)$, there exists a self-adjoint extension, denoted $A_{L}$. By Lemma 2.2, this extension corresponds to a Lagrangian subspace $L$. Moreover, by Lemma 2.8, the path $s\mapsto A_{L}+sI$ is gap-continuous, and hence the spectral flow is well-defined. Finally, by the Unique Continuation Property, the path $s\mapsto p(V_{s})$ is also gap-continuous and so the $\mu^{\scriptscriptstyle{\mathrm{CLM}}}$-index is well-defined, too. Thus, both sides of Equation (2.9) are well-defined.
Proof.
Crossing instants are isolated being zeros of an analytic function and so on the compact interval $I$ are in a finite number. To prove the theorem, it is enough to show that the local contributions to the Maslov index and to the spectral flow are equal. For, we assume that $s_{0}\in[0,1]$ is a crossing instant and we observe that
$\dim\ker(A_{L}+s_{0}I)=\dim\big(L\cap p(V_{s_{0}})\big).$ (2.10)
Moreover $s\mapsto A_{L}+sI$ is a positive curve, $s\mapsto p(V_{s})$ is a negative curve as proved in Lemma 2.11 and the crossing instants for these two paths are in 1-1 correspondence. This, in particular implies the following two facts:
  • The local contribution to the spectral flow which is given by the signature of the crossing form at $s_{0}$ coincides with the (LHS) of Equation (2.10)
  • The local contribution to the Maslov index which is given by the signature of the crossing form at $s_{0}$ coincides with the negative of the (RHS) of Equation (2.10).
The conclusion follows by summing all over the crossings. ∎
Remark 2.14.
We briefly indicate a geometric interpretation of Proposition 2.12 in the spirit of the (infinite-dimensional) Maslov index, without developing that theory in detail.
Consider the pair of subspaces
$H\times\{0\},\qquad\operatorname{gr}(A_{L}+sI),$
and equip $H\oplus H$ with the standard symplectic form
$\widetilde{\omega}\big((u_{1},v_{1}),(u_{2},v_{2})\big)\;=\;\langle u_{1},v_{2}\rangle-\langle v_{1},u_{2}\rangle,$
where $\langle\cdot,\cdot\rangle$ denotes the inner product on $H$. For $s\in{\mathbb{R}}$ set $A_{s,L}:=A_{L}+sI$ and take a path
$z(s):=(u,A_{s,L}u)\in\operatorname{gr}(A_{s,L}).$
Then, at a fixed parameter $s_{0}$,
$\frac{d}{ds}\,\tilde{\omega}\!\big(z(s_{0}),z(s)\big)=\frac{d}{ds}\,\big(\langle u,A_{s,L}u\rangle-\langle A_{s_{0},L}u,u\rangle\big)=\left\langle u,\dfrac{d}{ds}A_{s,L}\,u\right\rangle=\langle u,u\rangle,$
since $\tfrac{d}{ds}A_{s,L}=I$. By Definition A.2, the crossing form
$\Gamma\big(\operatorname{gr}(A_{s,L}),\,H\times\{0\}\big)$
is therefore positive.
Fix $s_{0}$ and choose a closed complement $V$ to $\ker A_{s_{0}}^{*}$ so that, in a neighborhood of $s_{0}$,
$D(A_{s}^{*})\;=\;D(A_{s})\oplus\big(\ker A_{s}^{*}\oplus V\big).$
Set $W_{s}:=D(A_{s})\oplus\big(\ker A_{s}^{*}\oplus V\big)$. Then
[…]
is a symplectic subspace. Let
$\Lambda_{s}\;=\;W_{s}\cap D(A_{s,L}).$
Inside the symplectic subspace $\operatorname{gr}(A_{s}^{*}|_{W_{s}})$, the subspaces
$\ker(A_{s}^{*})\times\{0\}\ <\ H\times\{0\},\qquad\operatorname{gr}\big(A_{s}^{*}|_{\Lambda_{s}}\big)\ <\ \operatorname{gr}(A_{s,L})$
are Lagrangian. Since $\Gamma\big(\operatorname{gr}(A_{s,L}),H\times\{0\}\big)$ is positive, its restriction to the Lagrangian pair
$\big(\operatorname{gr}(A_{s}^{*}|_{\Lambda_{s}}),\ \ker(A_{s}^{*})\times\{0\}\big)$
is also positive:
$\Gamma\big(\operatorname{gr}(A_{s}^{*}|_{\Lambda_{s}}),\ \ker A_{s}^{*}\times\{0\},\ \widetilde{\omega}\big)\;>\;0.$
Finally, using the natural identifications
$\operatorname{gr}\big(A_{s}^{*}|_{U_{s}}\big)\ \cong\ U_{s}\ \xrightarrow{\;\;p\;\;}\ U,$
(where $p$ denotes the projection onto $U$ associated with $A_{s}$), we obtain that the induced crossing form
$\Gamma\big(\Lambda,\ p(\ker A_{s}^{*}),\ \rho\big)$
is positive, where $\Lambda:=p(\Lambda_{s})$ is the boundary data space and $\rho$ is the induced symplectic form on $U$. This identifies the positivity in Proposition 2.12 with a positivity statement for the Maslov crossing in the boundary phase space.
3 An Abstract Spectral Flow Formula for SL-Operators
The primary objective of this section is to prove Theorem 3.3, which establishes a spectral flow formula for a class of one‑sided singular Sturm–Liouville operators. We begin by deriving several preliminary results on gap-continuity that are crucial for the proof of the main theorem. For a comprehensive treatment of regular and singular Sturm–Liouville differential operators, see [Zet05] and the references therein.
3.1 A splitting formula for gap-continuous paths
Let $a,b\in\overline{{\mathbb{R}}}$. We define
$l_{s}:=-\dfrac{d}{dt}\left(P(t)\dfrac{d}{dt}+Q(t)\right)+{Q}^{\mathsf{T}}(t)\dfrac{d}{dt}+R(t)+C_{s}(t)\qquad\textrm{ for }\quad(t,s)\in(a,b)\times[0,1]$ (3.1)
where
$P^{-1}\in\mathscr{C}^{1}\big((a,b),\mathrm{S}_{n}({\mathbb{R}})\big)\qquad Q\in\mathscr{C}^{1}\big((a,b),\mathrm{M}_{n}({\mathbb{R}})\big)\qquad R\in\mathscr{C}^{0}\big((a,b),\mathrm{S}_{n}({\mathbb{R}})\big)\\[3.0pt] \textrm{ and }s\mapsto C_{s}\in\mathscr{C}^{0}_{b}\big((a,b),\mathrm{S}_{n}({\mathbb{R}})\big)\textrm{ is continuous}$ (3.2)
where $\mathscr{C}^{0}_{b}$ denotes the set of bounded continuous functions and where $C_{0}(t)=0$ for every $t\in(a,b)$.
In the Hilbert space $H=L^{2}((a,b),\mathbb{R}^{n})$, we refer to the minimal operator $L_{s}$ defined by $l_{s}$ as the perturbed SL-operator and to the minimal operator $L$ corresponding to $s=0$ as the unperturbed SL-operator. We denote by $L_{s}^{*}$ and by $L^{*}$ the maximal operator associated to $l_{s}$ and to $l$ respectively and we observe that both domains $D(L_{s})$ and $D(L_{s}^{*})$ are independent on $s$.
We now prove that under the assumption (H1), the unique continuation property holds for the Sturm-Liouville operator $l$.
Lemma 3.1.
We assume that condition (H1) holds and let $L$ be the minimal operator associated to the Sturm-Liouville operator $\ell$ defined at Equation (3.1). Then we get that the equation
Lu=0L\,u=0
has no nontrivial solutions in $D(L)$.
Proof.
Arguing by contradiction, we assume that there exists a nontrivial solution $0\neq u$ in $H$ and by (H1) the operator $L$ is only one-sided singular at $t=a$. We note that for each $u\in D(L)$ it holds that $u(b)=u^{\prime}(b)=0$. For every $d\in(a,b)$ the equation Lu=0L\,u=0 is a regular Sturm-Liouville equation on $[d,b]$ and by this we get that $u(t)=0$ for every $t\in(d,b]$. This argument implies that $u(t)=0$ for every $t\in(a,b]$, concluding the proof. ∎
Before proving the gap-continuity for a family of self-adjoint Fredholm extension operators, we start with the following well-known preliminary result.
Lemma 3.2.
Under assumptions (H1) and (H2), then we get:
  1. 1.
    The path of minimal operators $s\mapsto L_{s}$ is gap-continuous
  2. 2.
    $s\mapsto p(V_{s})$ is gap-continuous in $\operatorname{\mathcal{S}}(U)$
  3. 3.
    $s\mapsto\Lambda_{s}\in\operatorname{\mathrm{Lag}}(U,\rho)$ is a gap-continuous Lagrangian path if and only if
    $s\mapsto L_{s,\Lambda_{s}}:=L_{s}^{*}|_{D(L)\oplus\Lambda_{s}}$
    is a gap-continuous path of operators in $\operatorname{\mathcal{CF}^{sa}}(H)$.
Proof.
We start by observing that item (1) is a direct consequence of Lemma 2.8, since $s\mapsto L_{s}$ is a bounded perturbation of the closed operator $L$. Item (2) comes from Lemma 2.10 and Lemma 3.1.
We are now in a position to prove item (3). Assume that $s\mapsto\Lambda_{s}$ is a gap-continuous path of Lagrangian subspaces in $U$. Then it also defines a gap-continuous path in $\operatorname{\mathcal{S}}(H)$. Since
$\operatorname{gr}(L_{s}^{*}|_{\Lambda_{s}})=\operatorname{gr}(L_{s}^{*})\cap(\Lambda_{s}\oplus H),$
by Lemma D.1, we get that the path $s\mapsto\operatorname{gr}(L_{s}^{*}|_{\Lambda_{s}})$ is a gap-continuous path in $\operatorname{\mathcal{S}}(H\times H)$, and hence
$\operatorname{gr}(L_{s,\Lambda_{s}})=\operatorname{gr}(L_{s})\oplus\operatorname{gr}(L_{s}^{*}|_{\Lambda_{s}})$
is gap-continuous in $\operatorname{\mathcal{S}}(H\times H)$.
Conversely, suppose that $s\mapsto\operatorname{gr}(L_{s,\Lambda_{s}})$ is a gap-continuous path in $\operatorname{\mathcal{S}}(H\times H)$. Then
$s\mapsto\operatorname{gr}(L_{s}^{*}|_{\Lambda_{s}})=\operatorname{gr}(L_{s,\Lambda_{s}})\cap(U\times H)$
defines a gap-continuous path in $\operatorname{\mathcal{S}}(H\times H)$. By invoking once again Lemma D.1, we get that
$\Lambda_{s}\times H=\operatorname{gr}(L_{s}^{*}|_{\Lambda_{s}})+(\{0\}\times H)$
is a gap-continuous path in $\operatorname{\mathcal{S}}(H\times H)$. In conclusion, $s\mapsto\Lambda_{s}$ is gap-continuous in $\operatorname{\mathcal{S}}(H)$, and hence also in $\operatorname{\mathcal{S}}(U)$.
The main result of this section is the following spectral flow formula.
Theorem 3.3.
Under assumptions (H1) and (H2) and if $s\mapsto\Lambda_{s}\in\operatorname{\mathrm{Lag}}(U,\rho)$ is a gap-continuous Lagrangian path, then we get:
$\operatorname{sf}(L_{s,\Lambda_{s}},s\in[0,1])=-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{s},p(V_{s}),\rho,s\in[0,1]).$
Proof.
As a direct consequence of Lemma 3.2, we obtain that the map $s\mapsto L_{s,\Lambda_{s}}$ is gap-continuous. Moreover, by the previous discussion, the two Lagrangian maps $s\mapsto\Lambda_{s}$ and $s\mapsto p(V_{s})$ are continuous.
We also observe that $L_{s}+rI=L+C_{s}+rI$, and by Lemma 2.8, it follows that the map $(s,r)\mapsto L_{s}+rI$ is continuous with respect to the gap topology.
Using the localization properties of the Maslov index and the spectral flow, to conclude the proof, it suffices to verify that the formula holds in a sufficiently small neighborhood of any fixed instant $s\in[0,1]$. Without loss of generality, we may localize near 0.
Let us choose $\varepsilon>0$ such that $L_{0,\Lambda_{0}}+\varepsilon I$ is invertible. Since invertibility is an open condition, there exists $\delta>0$ such that $L_{s,\Lambda_{s}}+\varepsilon I$ remains invertible for all $s\in[0,\delta]$.
By the homotopy invariance properties of the spectral flow and the Maslov index, we obtain:
$\operatorname{sf}(L_{s,\Lambda_{s}},\,s\in[0,\delta])=\operatorname{sf}(L_{0,\Lambda_{0}}+rI,\,r\in[0,\varepsilon])-\operatorname{sf}(L_{\delta,\Lambda_{\delta}}+rI,\,r\in[0,\varepsilon])\quad\text{and}\\[5.0pt] \mu^{\scriptscriptstyle{\mathrm{CLM}}}\big(\Lambda_{s},p(V_{s}),\rho,\,s\in[0,\delta]\big)=\mu^{\scriptscriptstyle{\mathrm{CLM}}}\big(\Lambda_{0},p(V_{0,r}),\rho,\,r\in[0,\varepsilon]\big)\\ -\mu^{\scriptscriptstyle{\mathrm{CLM}}}\big(\Lambda_{\delta},p(V_{\delta,r}),\rho,\,r\in[0,\varepsilon]\big).$ (3.3)
From this computation, and by invoking once again Proposition 2.12, we immediately deduce that
$\operatorname{sf}(L_{s,\Lambda_{s}},\,s\in[0,\delta])=-\mu^{\scriptscriptstyle{\mathrm{CLM}}}\big(\Lambda_{s},p(V_{s}),\rho,\,s\in[0,\delta]\big).$
The conclusion then follows by summing up all local contributions to both the Maslov index and the spectral flow.
4 An explicit decomposition and a trace map
The spectral flow formula for one‐sided singular Sturm–Liouville operators provided at Theorem 3.3 is given, in terms of a splitting of the maximal domain and of a coordinate map $O$. In this section we implement $U,O$ and $p$ concretely by means of a trace map, built from the fundamental solutions of the Sturm–Liouville equation. We then show that, in the regular (nonsingular) case, our abstract formula specializes exactly to the well‐known classical spectral flow formulas (see, e.g., [BOPW21, HP17]). Finally, we give a brief, self‐contained derivation of all self‐adjoint boundary conditions for a Sturm–Liouville operator.
Remark 4.1.
For the sake of the reader we observe that Shi and Sun [SS10] have been already characterized all self‐adjoint boundary conditions for Hamiltonian systems via the Glazman–Krein–Naimark (GKN) theory.
4.1 The trace map
We are now ready to give a concrete splitting decomposition
$D(L^{*})=D(L)\oplus U$
where $U$ is a $2k$-dimensional subspaces.
[…]
Lemma 4.2.
Let $L\in\operatorname{\mathcal{CF}^{s}}(H)$ and we assume condition (H1). Then the following formula holds:
$D(L^{*})=D(L)+\ker L^{*}+Z.$
Proof.
Since $\dim(\ker L^{*}+Z)<+\infty$, then we have $(\ker L^{*}+Z)^{\omega\omega}=\ker\omega+(\ker L^{*}+Z)=D(L)+\ker L^{*}+Z$ . Since $D(L)]^{\omega}=D(L^{*})$, we only need to show that $(\ker L^{*}+Z)^{\omega}=D(L)$.
By arguing as in the proof of Lemma 2.6, we get that
$(\ker L^{*})^{\omega}=D(L)+\ker L^{*}$
and so
$(\ker L^{*})^{\omega}\cap Z^{\omega}=D(L)+(\ker L^{*}\cap Z^{\omega}).$
Let now $v\in Z^{\omega}$ and we observe that for every $i=1,\ldots,2n$ it follows that $\omega(v,z_{i})=[v,z_{i}](b)=0$. Since $Z(b)$ is a basis of ${\mathbb{R}}^{2n}$, then we get that $v^{\prime}(b)=v(b)=0$. So, $\ker L^{*}\cap Z^{\omega}=\set{0}$ by the Cauchy-Lipschitz Theorem, concluding the proof. ∎
The next result provides a characterization of the $D(L)$ in terms of the $\ker L^{*}$ and of the boundary values at the regular endpoint.
Corollary 4.3.
Let $L\in\operatorname{\mathcal{CF}^{s}}(H)$, let $f\in D(L^{*})$ and we assume condition (H1). Then $f\in D(L)$ if and only if $\omega(f,v)=0$ for all $v\in\ker L^{*}$ and $f^{\prime}(b)=f(b)=0$.
Proof.
[…]
This concludes the proof. ∎
Lemma 4.4.
Let $L\in\operatorname{\mathcal{CF}^{s}}(H)$ and we assume condition (H1). Let $Y:=\set{y_{1},\cdots,y_{k}}$ be a basis of $\ker L^{*}$. Then we have that $n\leq k\leq 2n$. Moreover, there exists a $2k$-dimensional subspace $W$ such that
$D(L^{*})=D(L)\oplus W$
where
  • $W:=\operatorname{span}(\overline{Y})\oplus Z$ and where $\overline{Y}<Y$ such that $\dim\overline{Y}=2k-2n$ if $k\neq n$
  • $W:=Z$ for $k=n$.
Remark 4.5.
We observe that the case $k=2n$ corresponds to either limit circle endpoints or regular conditions, whereas the case $k=n$ corresponds to the limit point. (Cfr. [Zet05] and references therein for the endpoints classification of the Sturm-Liouville BVP).
Proof.
By Corollary 4.3, $D(L^{*})\cap Z=\set{0}$. Then by Lemma 4.2, we can choose $\overline{Y}<Y$ such that
$D(L^{*})=D(L)\oplus(\operatorname{span}(\overline{Y})\oplus Z).$
By Lemma 2.6, $\dim(\operatorname{span}(\overline{Y})\oplus Z)=2\dim(\ker L^{*})=2k$. It follows that
$Y=\dim(\operatorname{span}(\overline{Y}))=\dim(\operatorname{span}(\overline{Y})\oplus Z)-\dim Z=2k-2n.$
This concludes the proof.
Trace map $\operatorname{\mathcal{T}}$ and the subspace $W$
Summing up the above discussion we finally get the decomposition
[…]
  • If $k=n$, we consider the symplectic space $(W,\omega|_{W})$ and we define the trace map $\operatorname{\mathcal{T}}:W\to{\mathbb{R}}^{2n}$ as:
    • $f\longmapsto\operatorname{\mathcal{T}}(f):=(\omega(f,z_{1}),\cdots,\omega(f,z_{2n}))=([f,z_{1}](b),[f,z_{2n}](b))\in{\mathbb{R}}^{2n}.$
  • If $k\neq n$, we observe that since $\omega|_{Z}$ is nondegenerate, then $Z^{\omega|_{W}}=Z^{\omega}\cap W$ is a symplectic subspace of $(W,\omega|_{W})$. Then $Z^{\omega|_{W}}$ has a basis $\set{\widetilde{y}_{1},\cdots,\widetilde{y}_{2k-2n}}$ such that $\widetilde{y}_{i}-y_{i}\in W$ and $\widetilde{y}_{i}^{\prime}(b)=\widetilde{y}_{i}(b)=0$ for every $1\leq i\leq 2k-2n$. Then $W$ has the following symplectic decomposition:
    $W=Z^{\omega|_{W}}\oplus Z.$ (4.4)
    Notation 4.6.
    We set
    $f,g](a^{+}):=\lim_{t\to a^{+}}[f,g](t).$
    In this case, we define the (linear) trace map $\operatorname{\mathcal{T}}:W\to{\mathbb{R}}^{2k-2n}\oplus{\mathbb{R}}^{2n}$ as:
    $f\longmapsto\operatorname{\mathcal{T}}(f):=(\omega(f,\widetilde{y}_{1}),\cdots,\omega(f,\widetilde{y}_{2k-2n}),\omega(f,z_{1}),\cdots,\omega(f,z_{2n}))\\ =(-[f,y_{1}](a^{+}),\cdots,-[f,y_{2k-2n}](a^{+}),[f,z_{1}](b),\ldots,[f,z_{2n}](b))\in{\mathbb{R}}^{2k-2n}\oplus{\mathbb{R}}^{2n}$ (4.5)
Since $\omega|_{W}$ is nondegenerate, it follows that
  • $\operatorname{\mathcal{T}}:Z^{\omega|_{W}}\oplus Z\to{\mathbb{R}}^{2k-2n}\oplus{\mathbb{R}}^{2n}$ is a bijection and moreover the decomposition is $\operatorname{\mathcal{T}}$-invariant.
Remark 4.7.
In the construction above we defined the trace map, but we did not specify how to select the 2k2n2k-2n functions $\{y_{1},\dots,y_{2k-2n}\}\subset\ker L^{*}$ so that
$W=\operatorname{span}\{y_{1},\dots,y_{2k-2n},\,z_{1},\dots,z_{2n}\}.$
We choose them to make $\omega|_{W}$ nondegenerate. It suffices to ensure that the restriction
$\omega\big|_{\operatorname{span}\{\tilde{y}_{1},\dots,\tilde{y}_{2k-2n}\}}$
is nondegenerate, where $\{\tilde{y}_{i}\}$ is an appropriate reordering of a basis of $\ker L^{*}$. To this end, start with any basis $\{y_{1},\dots,y_{k}\}$ of $\ker L^{*}$ and consider the $k\times k$ Gram matrix $G=(G_{ij})$ of the boundary pairing
$G_{ij}\;=\;[y_{i},y_{j}](a),\qquad 1\leq i,j\leq k.$
Reorder the basis so that the principal $(2k-2n)\times(2k-2n)$ block $\big([y_{i},y_{j}](a)\big)_{1\leq i,j\leq 2k-2n}$ is invertible (this is always possible as soon as $\operatorname{rank}G\geq 2k-2n$). Then take
$\{y_{1},\dots,y_{2k-2n}\}$
to be the first 2k2n2k-2n vectors in this reordered basis. By construction, $\omega$ is nondegenerate on $\operatorname{span}\{y_{1},\dots,y_{2k-2n}\}$, and hence $\omega|_{W}$ is nondegenerate as claimed.
Remark 4.8.
We observe that the space
${\mathbb{R}}^{2k-2n}\oplus{\mathbb{R}}^{2n}$
is the space of boundary data and actually ${\mathbb{R}}^{2k-2n}$ (resp. ${\mathbb{R}}^{2n}$) contains the boundary data at $a$ (resp. $b$).
Summing up the previous discussion, it is possible to characterize $D(L)$ in terms of $D(L^{*})$ and the trace map $\operatorname{\mathcal{T}}$.
Corollary 4.9.
We assume condition (H1) holds. Let
  1. 1.
    $f\in D(L^{*})$. Then $f\in D(L)$ if and only if $\operatorname{\mathcal{T}}(f)=0$
  2. 2.
    Let $f,y_{1}](a^{+}),\cdots,[f,y_{2k-2n}](a^{+})=0$. Then $f,g](a^{+})=0$ for each $g\in D(L^{*})$.
Proof.
We start by proving item 1. By the very definition of $\operatorname{\mathcal{T}}$, we get that $\operatorname{\mathcal{T}}(D(L))=\set{0}$ since $D(L)=\ker\omega$. Viceversa, we observe that $\operatorname{\mathcal{T}}:W\to{\mathbb{R}}^{2k-2n}\oplus{\mathbb{R}}^{2k}$ is a bijection and $D(L^{*})=D(L)\oplus W$. So, we get that $\operatorname{\mathcal{T}}(f)=0$ implies $f\in D(L)$.
We prove item 2. Since $f,y_{1}](a^{+}),\cdots,[f,y_{2k-2n}](a^{+})=0$, then we have $f\in\operatorname{\mathcal{T}}^{-1}(\set{0}\times{\mathbb{R}}^{2n})=D(L)\oplus Z$. By setting $f=u+v$ with $v\in Z$, we get that $f,g](a^{+})=[v,g](a^{+})=0$.
By using the trace map defined above, the spectral flow formula provided in Theorem 3.3 reduces to the following.
Theorem 4.10.
Under conditions (H1)-(H2) and assuming that $s\mapsto\Lambda_{s}\in\operatorname{\mathrm{Lag}}(W,\omega|_{W})$ is gap-continuous, then the following spectral flow formula holds:
$\operatorname{sf}(L_{s,\Lambda_{s}},s\in[0,1])=-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\operatorname{\mathcal{T}}(\Lambda_{s}),\operatorname{\mathcal{T}}(V_{s}),\rho^{\operatorname{\mathcal{T}}},s\in[0,1]),$
where $\rho^{\operatorname{\mathcal{T}}}$ denotes the push-forward symplectic form.
Proof.
The proof of this result is a direct consequence of Theorem 3.3, Lemma 4.2, Corollary 4.3, Lemma 4.4, and Corollary 4.9. ∎
4.2 Two classical examples
We now are ready to show that the spectral flow formula given at Theorem 4.10 is a generalization of the well-known spectral flow formulas. Actually in the already known
  1. 1.
    Case of regular SL-operators on bounded domains
  2. 2.
    Case of SL-operators on $\mathbb{R}$$^{+}$.
it reduces to the already known formulas.
Regular SL-operators on bounded intervals
For $s\in[0,1]$, we consider the Sturm-Liouville operator $l_{s}$ given at Equation (3.1) and we let
$X=\operatorname{span}\set{x_{1},\cdots,x_{2n}}\qquad\textrm{ and }\qquad Y=\set{y_{1},\cdots,y_{2n}}$
where
  • $x_{i}$ are smooth functions defined on $[a,b]$ and vanishing on the interval $[(a+b)/2,b]$ .
  • $y_{i}$ are smooth functions defined on $[a,b]$ and vanishing on the interval $[a,(a+b)/2]$.
We set
$U=X\oplus Y$
and we let $\mathcal{E}_{a}$ and $\mathcal{E}_{b}$ be the $(2n)$-dimensional subspaces isomorphic to ${\mathbb{R}}^{2n}$ and defined by
[…]
We denote by $O:U\to{\mathbb{R}}^{2n}\oplus{\mathbb{R}}^{2n}$ be the coordinate map defined by
$O(u)=\big(P(a)u^{\prime}(a)+Q(a)u(a),u(a),P(b)u^{\prime}(b)+Q(b)u(b),u(b)\big)$ (4.7)
and we consider the decomposition $D(L^{*})=D(L)\oplus U$ and let $p:D(L^{*})\to U$ be the projection onto the second factor.
Denoting by $M_{s}$ the monodromy matrix of the Hamiltonian system induced by the Sturm-Liouville operator $l_{s}$, then we get that
$O(p(V_{s}))=\operatorname{gr}(M_{s})\qquad\textrm{ for every }s\in[0,1]\qquad\textrm{ and }\qquad V_{s}=\ker A_{s}^{*}.$
This equality is a direct consequence of the definition of the monodromy matrix $M_{s}$ being the projection of the solution space onto the space of boundary conditions. Let $s\mapsto\Lambda_{s}\in\operatorname{\mathrm{Lag}}(U,\rho)$ be a continuous Lagrangian path. Then we get that
[…]
and the following spectral flow formula holds
$\operatorname{sf}(L_{s,\Lambda_{s}},s\in[0,1])=-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(O(\Lambda_{s}),\operatorname{gr}(M_{s}),-\Omega\oplus\Omega,s\in[0,1])$ (4.8)
where $\Omega$ denotes the standard symplectic form of ${\mathbb{R}}^{2n}$.
Example 4.11.
The spectral flow formula provided at Equation (4.8) coincides with the well-known spectral flow formula for the case of regular Sturm-Liouville operator. Actually, the standard framework is the following. Let $s\mapsto Y_{s}$ be a path in $\operatorname{\mathrm{Lag}}({\mathbb{R}}^{2n}\oplus{\mathbb{R}}^{2n},-\Omega\oplus\Omega)$ and not in the functional space $D(L^{*})$. The path $s\mapsto Y_{s}$ induces a path $s\mapsto L_{s,Y_{s}}$ of self-adjoint operators with domain
$D(L_{s,Y_{s}}):=\left\{u\in W^{2,2}((a,b),{\mathbb{R}}^{n})|\right.\\ \left.(P(a)u^{\prime}(a)+Q(a)u(a),u(a),P(b)u^{\prime}(b)+Q(b)u(b),u(b))\in Y_{s}\right\}$ (4.9)
and the spectral flow formula reduces to the usual one
$\operatorname{sf}(L_{s,Y_{s}},s\in[0,1])=-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(Y_{s},\operatorname{gr}(M_{s}),-\Omega\oplus\Omega,s\in[0,1]).$ (4.10)
By comparing the Equation (4.8) and Equation (4.10), the only difference is the presence of the mapping $O$. This is a direct consequence of the fact that in former case the paths are in the functional space $D(L^{*})$ whilst in the latter case are directly in ${\mathbb{R}}^{2n}\oplus{\mathbb{R}}^{2n}$. So, the two formulas coincide once definining $O(\Lambda_{s})=Y_{s}$ for every $s\in[0,1]$.
Remark 4.12.
The last remark is that in Theorem 3.3 the symplectic form appearing in the computation of the Maslov index is $\rho^{O}$ and not just $\Omega\oplus\Omega$. This is because the aforementioned theorem works also in the singular case in which the explicit coordinate map given at Equation (4.7) is not well-defined anymore.
SL-operators on the half-line
We consider a Sturm-Liouville operator $l_{s}$ acting on $C_{0}^{\infty}({\mathbb{R}}^{+},{\mathbb{R}}^{n})$ and we assume that
$P^{-1}\in\mathscr{C}^{1}([0,+\infty),\mathrm{S}_{n}({\mathbb{R}}))\qquad Q\in\mathscr{C}^{1}([0,+\infty),\mathrm{M}_{n}({\mathbb{R}}))\\ \quad R\in\mathscr{C}^{0}\big([0,1],\mathrm{S}_{n}(\mathscr{C}^{0}([0,+\infty),{\mathbb{R}}))\big)\quad\textrm{ are bounded on ${\mathbb{R}}$}.$ (4.11)
In this case, it’s easy to check that $D(L_{s})=D(L)$ is independent on $s$; moreover we have
$D(L)=W^{2,2}_{0}({\mathbb{R}}^{+},{\mathbb{R}}^{n})\qquad D(L^{*})=W^{2,2}({\mathbb{R}}^{+},{\mathbb{R}}^{n})$
Then the mapping $s\mapsto L_{s}$ is a gap-continuous path of self-adjoint operators.
Let us assume that $s\mapsto L_{s}$ is a path in $\operatorname{\mathcal{CF}^{sa}}(H)$ [^3] This is, for instance, the case if $\lim_{t\to+\infty}\begin{pmatrix}P(t)&Q(t)\\ Q^{T}(t)&R(t)+C_{s}(t)\end{pmatrix}$ exists and it is positive definite matrix. and let us consider $U=\operatorname{span}\{x_{1},\ldots,x_{2n}\}$ where
  • $x_{i}$ are smooth functions defined on $0,+\infty)$ and vanishing on $(1,+\infty)$.
  • $\mathcal{E}_{0}$ is the linear subspace defined by
[…]
Let us now consider the decomposition $D(L^{*})=D(L)\oplus U$ where $(U,\rho)$ is the symplectic space with respect to the symplectic form defined by $\rho=\omega|_{U}$. We define the coordinate map
$O:U\longrightarrow{\mathbb{R}}^{2n}\quad\textrm{ defined by }\quad O(u)=\big(P(0)u^{\prime}(0)+Q(0)u(0),u(0)\big)$
and let $p:D(L^{*})\to U$ be the projection onto the second factor. We let $s\mapsto\Lambda_{s}\in\operatorname{\mathrm{Lag}}(U,\rho)$. Then the following spectral flow formula holds
$\operatorname{sf}(L_{s,\Lambda_{s}},s\in[0,1])=-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(O(\Lambda_{s}),O(p(V_{s}),-\Omega,s\in[0,1])$ (4.13)
where $V_{s}:=\ker(L_{s}^{*})$ is the linear subspace of all $L^{2}$ solutions of the system on the half-line $0,+\infty)$. So, in particular, $O(p(V_{s}))$ is just the boundary value of these solutions or which is the same the stable subspace of the system at 0.
Example 4.13.
The spectral flow formula provided at Equation (4.13) coincides with the well-known spectral flow formula for the case of Sturm-Liouville operator on the half-line as proved by authors in [HP17]. The framework in that paper is to start with a path $s\mapsto Z_{s}$ is in $\operatorname{\mathrm{Lag}}({\mathbb{R}}^{2n},\Omega)$ and not in the functional space $D(L^{*})$. The path $s\mapsto Z_{s}$ induces a path $s\mapsto L_{s,Z_{s}}$ of self-adjoint operators with domain
$D(L_{s,Z_{s}}):=\left\{u\in W^{2,2}([0,+\infty),{\mathbb{R}}^{n})|\right.\left.(P(0)u^{\prime}(0)+Q(0)u(0),u(0))\in Z_{s}\right\}$ (4.14)
and the spectral flow formula reduces to the usual one
$\operatorname{sf}(L_{s,Z_{s}},s\in[0,1])=-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(Z_{s},W_{s}^{st}(0),-\Omega,s\in[0,1]).$
Here the $W_{s}^{st}(0)$ is the stable subspace of at 0 and it is defined by
[…]
5 The Morse Index Theorem for Singular Sturm-Liouville
Theorem 4.10 provides an abstract and general spectral flow formula for a family of Sturm–Liouville differential operators. It serves as a foundational step toward establishing a Morse Index Theorem within this singular framework.
The aim of this section is to apply the spectral flow formula described above to the specific case of a single Sturm–Liouville operator $l$ defined on the function space $\mathscr{C}_{0}^{\infty}((a,b],\mathbb{R}^{n})$, where the endpoint $t=b$ is regular according to hypothesis (H1).
In this context, by restricting the operator $l$ to the interval $(a,\sigma)$ with $\sigma\in(a,b]$, we obtain a one-parameter family of operators $\sigma\mapsto l_{(a,\sigma)}$. This construction is necessary for the application of the spectral flow formula.
Let $\sigma\in(a,b)$. We define $l_{(a,\sigma)}$ be the one-sided singular Sturm-Liouville differential operator obtained by restricting $l$ on $\mathscr{C}_{0}^{\infty}((a,\sigma),{\mathbb{R}}^{n})$. We denote by $L_{(a,\sigma)}$ be its minimal operator associated to $l_{(a,\sigma)}$ and we assume that $L_{(a,\sigma)}$ is singular only at the starting instant $t=a$ (and so the final instant $\sigma$ is a regular endpoint). We consider the Lagrangian subspaces
$\Lambda_{0}\in\operatorname{\mathrm{Lag}}({\mathbb{R}}^{2k-2n})\quad\textrm{ and }\Lambda_{D}:={\mathbb{R}}^{n}\times(0)\qquad\textrm{ and }\Lambda=\Lambda_{0}\oplus\Lambda_{D}$
where the $\Lambda_{D}$ is standard Dirichlet Lagrangian. We introduce the following two conditions.
    (H3) The quadratic form on $\mathscr{C}^{\infty}_{0}((a,b),{\mathbb{R}}^{n})$ associated to the SL-operator $l$ is bounded from below in $L^{2}$.
    (H4) The minimal operator $L_{b}\in\operatorname{\mathcal{CF}^{s}}(L^{2}((a,b),{\mathbb{R}}^{n}))$.
Given $\Lambda\in\operatorname{\mathrm{Lag}}({\mathbb{R}}^{2k-2n}\oplus{\mathbb{R}}^{2n},-\Omega\oplus\Omega)$, we set
$L_{(a,\sigma),\Lambda}=L_{(a,\sigma)}^{*}|_{D(L_{(a,\sigma)})\oplus\operatorname{\mathcal{T}}_{\sigma}^{-1}(\Lambda)}$
where $\operatorname{\mathcal{T}}_{\sigma}:W_{\sigma}\to{\mathbb{R}}^{2k-2n}\oplus{\mathbb{R}}^{2n}$ is the trace map defined on the subspace $W_{\sigma}$ associated to the operator $l_{(a,\sigma)}$ and constructed as in Subsection 4.1.
Theorem 5.1.
[Morse Index Theorem] We assume that the SL-operator $l$ satisfies conditions (H1)-(H3)-(H4) and we assume Dirichlet boundary condition at $b$. If $\mathrm{n_{-}}(L_{b})<\infty$, then the following equality holds:
$\mathrm{n_{-}}(L_{b,\Lambda})=\sum_{\sigma\in(a,b)}\dim\ker L_{(a,\sigma),\Lambda}.$
The next result relates the Morse index of the Friedrich extension operator $L_{F}$ with the limit on $\sigma$ of the Friedrich operators $L_{(a,\sigma),F}$ and $L_{(\sigma,b),F}$ under the only assumption that the operator $L$ is lower bounded. In particular, neither Fredholmness nor the finiteness of the Morse index is assumed.
Theorem 5.2.
We assume condition (H3). If $l$ is regular at $a$ and if we assume Dirichlet boundary condition at this endpoint, then we get
$\mathrm{n_{-}}(L_{F})=\lim_{\sigma\to b^{-}}\mathrm{n_{-}}(L_{(a,\sigma),F})=\sum_{t\in(a,b)}\dim\big(\gamma(t)\Lambda_{D}\cap\Lambda_{D}\big),$
where $\gamma$ is the fundamental matrix solution such that $\gamma(a)=I$.
If $l$ is regular at $b$ and we assume Dirichlet boundary condition at this endpoint, then we have
$\mathrm{n_{-}}(L_{F})=\lim_{\sigma\to a^{+}}\mathrm{n_{-}}(L_{(\sigma,b),F})=\sum_{t\in(a,b)}\dim\big(\gamma(t)\Lambda_{D}\cap\Lambda_{D}\big),$
where $\gamma$ is the fundamental matrix solution such that $\gamma(b)=I$.
Proof.
We establish the result in the case where $a$ is a regular endpoint, since the other case is entirely analogous.
For, we start by observing that $\mathrm{n_{-}}(L_{b,F})=\mathrm{n_{-}}(t_{L_{b}})$ where $t_{L_{b}}$ is the closure of the quadratic form $u\mapsto(l_{b},u,u),u\in C_{0}^{\infty}((a,b),{\mathbb{R}}^{n})$. Let $W$ be a maximum negative subspace of $t_{L_{b}}$ and let $\set{w_{1},\cdots,w_{m}}$ be a basis of $W$. Then there exists $\set{w_{1}^{\prime},\cdots,w_{m}^{\prime}}\subset C_{0}^{\infty}((a,b),{\mathbb{R}}^{n})$ such that for every $\delta>0$ the following holds:
$|t_{L_{b}}(w_{i}^{\prime},w_{j}^{\prime})-t_{L_{b}}(w_{i},w_{j})|<\delta\qquad 1\leq i,j\leq m.$
In particular, if $\delta>0$ is sufficiently small, also $\widetilde{W}=\operatorname{span}\set{w_{1}^{\prime},\cdots,w_{m}^{\prime}}$ is a negative subspace of $t_{L_{b}}$. By choosing $\sigma\in(a,b)$ such that $\operatorname{supp}w_{i}^{\prime}\subset(a,\sigma),1\leq i\leq m$, then we get
$\mathrm{n_{-}}(L_{[a,\sigma],F})\geq m=\mathrm{n_{-}}(t_{L_{b}})=\mathrm{n_{-}}(L_{b,F}).$
Moreover $D(t_{L_{[a,\sigma]}})<D(t_{L_{b}})$; so, we have
$\mathrm{n_{-}}(L_{[a,\sigma],F})\leq\mathrm{n_{-}}(t_{L_{b}}).$
Putting all together, we conclude that
$\mathrm{n_{-}}(L_{b,F})=\mathrm{n_{-}}(L_{[a,\sigma],F})$
Since $L_{[a,\sigma],F}$ is a regular Sturm-Liouville operator with Dirichlet boundary condition, then we get that
$\mathrm{n_{-}}(t_{L_{[a,\sigma],F}})=\sum_{t\in(a,\sigma)}\dim\big(\Lambda_{D}\cap\operatorname{gr}(\gamma(t))\big)=\sum_{t\in(a,b)}\dim\big(\Lambda_{D}\cap\operatorname{gr}(\gamma(t))\big).$
This concludes the proof. ∎
Remark 5.3.
In the Theorem 5.2, we only assume that $l_{b}$ is lower bounded. No assumption of Fredholmness of $L_{b}$ is made.
It is worth noting that, if $\mathrm{n_{-}}(L_{b})$ is infinite, then the theorem is still true since also the limit and the sum appearing in the thesis are all infinite.
5.1 Preliminary Results on Lagrangian Intersection Theory
In this subsection, we prove some foundational results from Lagrangian intersection theory that will be used throughout the proofs. These include algebraic and topological properties of Lagrangian subspaces in symplectic vector spaces, as well as basic facts concerning the behavior of their intersections under perturbations.
We let $V_{\sigma,s}=\ker(L_{(a,\sigma)}^{*}+sI)$ and for $M>0$, we set
$m_{\sigma,s}:=\operatorname{\mathcal{T}}_{\sigma}(p_{\sigma}(V_{\sigma,s}))\qquad(\sigma,s)\in(a,b)\times[0,M]$
where $p_{\sigma}:D(L_{(a,\sigma)}^{*})\to W$. We notice that $m_{\sigma,s}\in\operatorname{\mathrm{Lag}}({\mathbb{R}}^{2n}\oplus{\mathbb{R}}^{2n},-\Omega\oplus\Omega)$ for every $(\sigma,s)\in(a,b)\times[0,M]$.
Lemma 5.4.
Under assumptions (H1), (H3) & (H4), the two-parameters family $(\sigma,s)\mapsto m_{\sigma,s}$ is continuous in $\operatorname{\mathrm{Lag}}({\mathbb{R}}^{2k-2n}\oplus{\mathbb{R}}^{2n},-\Omega\oplus\Omega)$.
Proof.
Let $\psi_{s}(\sigma)$ be the fundamental (matrix) solution of the Hamiltonian system defined by the Sturm-Liouville operator $l_{b}+sI$ such that $\psi_{s}(b)=I$ and we let
$G_{\sigma,s}:=\begin{pmatrix}I_{2k-2n}&0\\ 0&\psi_{s}(\sigma)\end{pmatrix}$.
Since the bijection
$V_{b,s}\ni v_{s}\longmapsto v_{s}|_{[a,\sigma]}\in V_{\sigma,s}$
we get that
$m_{\sigma,s}=G_{\sigma,s}\,m_{b,s}.$
By Lemma 2.10, the map $s\mapsto m_{b,s}$ is continuous. Moreover, since $s\mapsto\psi_{s}(b)$ is continuous (being $t=b$ a regular endpoint) so the map $(\sigma,s)\mapsto G_{\sigma,s}$ is continuous too. In conclusionn we get that the map $(\sigma,s)\mapsto m_{\sigma,s}$ is continuous. ∎
Notation 5.5.
From now on, we drop the index $s$ in the case of $s=0$. So, in particular $m_{\sigma,0}=m_{\sigma}$ and $V_{\sigma,0}=V_{\sigma}$.
We are going to prove that the path $\sigma\mapsto m(\sigma)$ is a plus curve with respect to the Lagrangian subspace $\Lambda_{0}\oplus\Lambda_{D}$ meaning that all crossing instants are positive. A direct consequence of this fact is that the local contribution to the Maslov index of a crossing instant is provided by the its multiplicity.
Lemma 5.6.
Under the assumptions of Lemma 5.4, we get that the path $\sigma\mapsto m(\sigma)$ is positive with respect to the Lagrangian subspace $\Lambda_{0}\oplus\Lambda_{D}$.
Proof.
By definition, we have
[…]
and $\mathcal{J}=\begin{pmatrix}-J_{2n}&0\\ 0&J_{2n}\end{pmatrix}$. Given a crossing instant $\sigma\in(a,b)$ and setting $M_{\sigma}=\begin{pmatrix}I&0\\ 0&\psi(\sigma)\end{pmatrix}$, we get that the crossing form
$C_{\sigma}:=-\mathcal{J}\dfrac{d}{d\sigma}G_{\sigma}\,G_{\sigma}^{-1}|_{m_{\sigma}\cap(\Lambda_{0}\oplus\Lambda_{D})}$
is a positive definite quadratic form. Actually, by a direct calculation, we get
$\mathcal{J}\left(\dfrac{d\,G_{\sigma}}{d\sigma}\right)G_{\sigma}^{-1}=\begin{pmatrix}0_{2n}&0\\ 0&-J_{2n}\dfrac{d\,\psi}{d\sigma}\psi(\sigma)^{-1}\end{pmatrix}$.
Let $u=\begin{pmatrix}u_{a}\\ u_{b}\end{pmatrix}\in m_{\sigma}\cap(\Lambda_{0}\oplus\Lambda_{D})$. By the uniqueness theorem for linear ODEs, we get that the following holds:
$m_{\sigma}\cap({\mathbb{R}}^{2n}\oplus(0))=\operatorname{\mathcal{T}}_{\sigma}(p_{\sigma}(V_{\sigma}))\cap({\mathbb{R}}^{2n}\oplus(0))=(0).$
Then $u_{b}\in\Lambda_{D}$ and $u\neq 0$ if and only if $u_{b}\neq 0$. For concluding the proof we only need to prove that
$J_{2n}\left(\dfrac{d\,\psi}{d\sigma}\right)\psi(\sigma)^{-1}\Big|_{\Lambda_{D}}$
is positive definite. To do so, we just observe that $J_{2n}\left(\dfrac{d\,\psi}{d\sigma}\right)\psi(\sigma)^{-1}=\begin{pmatrix}P^{-1}(\sigma)&0\\ 0&-R(\sigma)\end{pmatrix}$. Since $P^{-1}(\sigma)$ is positive definite, $J_{2n}\dfrac{d\,\psi}{d\sigma}\psi(\sigma)^{-1}$ is a positive definite quadratic form on $\Lambda_{D}$. Finally, $C_{\sigma}$ is a positive quadratic form on $m_{\sigma}\cap(\Lambda_{0}\oplus\Lambda_{D})$ concluding the proof. ∎
Proposition 5.7.
Under assumptions (H1), (H3) & (H4) and for every $c\in(a,b)$, we get
$\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{0}\oplus\Lambda_{D},m_{\sigma},-\Omega\oplus\Omega,\sigma\in[c,b])=\sum_{s\in[c,b)}\dim\ \Big((\Lambda_{0}\oplus\Lambda_{D})\cap m_{\sigma}\Big).$
Proof.
The proof follows directly upon observing that the crossing instants are regular, and hence isolated. Since the interval is compact, these crossing instants are finite in number. By Lemma 5.6, each local contribution to the Maslov index is determined by the multiplicity of the crossing, namely, the dimension of the intersection of the path $\sigma\mapsto m_{\sigma}$ with the reference Lagrangian $\Lambda_{0}\oplus\Lambda_{D}$. The desired conclusion is then obtained by summing these finitely many local contributions. ∎
We now prove that the Morse index of the self-adjoint extensions $L_{\sigma,\Lambda}$ varies monotonically as a function of the parameter $\sigma$.
Lemma 5.8.
Let $\Lambda=L_{0}\oplus L_{D}$. Under the assumptions (H1),(H3) & (H4) and for every $a\leq c\leq b$, the following inequality holds:
$\mathrm{n_{-}}(L_{c,\Lambda})\leq\mathrm{n_{-}}(L_{b,\Lambda}).$
Moreover there exists $c_{0}\in(a,b)$ such that $\mathrm{n_{-}}(L_{c_{0},\Lambda})=0$.
Proof.
Let $t_{L_{c,\Lambda}}$ be the closure of the quadratic form defined by $\langle L_{c,\Lambda}\cdot,\cdot\rangle_{L^{2}}$ having domain
$D(t_{L_{c,\Lambda}})=\overline{D(L_{c,\Lambda})}=\overline{D(L_{c})+\operatorname{\mathcal{T}}_{c}^{-1}(\Lambda)}$
where the closure is meant in the $D(t_{L_{c,\Lambda}})$. By Lemma D.11, we get that
$\mathrm{n_{-}}(t_{L_{c,\Lambda}})=\mathrm{n_{-}}(L_{c,\Lambda}).$
Moreover
[…]
So, $D(t_{L_{c,\Lambda}})=D(t_{L_{c}})+W_{c}$ where $W_{c}$ can be any subspace of $D(L_{c}^{*})$ such that $\operatorname{\mathcal{T}}(W_{c})=\Lambda_{0}\oplus\set{0}$.
For $d<c\leq b$, we choose $W_{d}$ in such a way that $D(t_{L_{c,\Lambda}})=D(t_{L_{c}})+W_{d}$ and $D(t_{L_{b,\Lambda}})=D(t_{L_{b}})+W_{d}$ .
The natural embedding $\iota:D(t_{L_{c}})\hookrightarrow D(t_{L_{b}})$ induces the embedding between the domains $D(t_{L_{c,\Lambda}})\hookrightarrow D(t_{L_{b,\Lambda}})$ pointwise defined by
$u\mapsto\widetilde{u}(x):=\begin{cases}u(x)&x\in[a,c]\\[3.0pt] 0&x\in[c,b].\end{cases}$
By this arguments and by invoking Lemma D.3, we immediately get that $D(t_{L_{c,\Lambda}})<D(t_{L_{b,\Lambda}})$ and $t_{L_{c,\Lambda}}=t_{L_{b,\Lambda}}|_{D(t_{L_{c,\Lambda}})}$. In particular, the following inequality holds
$\mathrm{n_{-}}(L_{c,\Lambda})\leq\mathrm{n_{-}}(L_{b,\Lambda}).$
Since $\bigcap_{a<\sigma<b}D(t_{L_{(a,\sigma),\Lambda}})=\set{0}$, then by invoking Lemma D.15 and Lemma C.2, there exists $c_{0}\in(a,b)$ such that $\mathrm{n_{-}}(L_{c_{0},\Lambda})=0$, concluding the proof.
Proof of Theorem 5.1
For any $c\in(a,b)$, there exists $M>0$ sufficiently large such that
$\mathrm{n_{-}}(L_{b,\Lambda})-\mathrm{n_{-}}(L_{c,\Lambda})=-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda,m_{b,s},-\Omega\oplus\Omega,s\in[0,M])+\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda,m_{c,s},-\Omega\oplus\Omega,s\in[0,M]).$
By the stratum homotopy invariance property of Maslov index, we have
$\mathrm{n_{-}}(L_{b,\Lambda})-\mathrm{n_{-}}(L_{c,\Lambda})\\ =-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda,m_{b,s},-\Omega\oplus\Omega,s\in[0,M])+\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda,m_{c,s},-\Omega\oplus\Omega,s\in[0,M])\\ =-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda,m_{\sigma,M},-\Omega\oplus\Omega,\sigma\in[c,b])+\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda,m_{\sigma,0},-\Omega\oplus\Omega,\sigma\in[c,b]).$ (5.1)
By Equation (5.1) and by Proposition 5.7, we get that for $M$ large enough,
$\mathrm{n_{-}}(L_{b,\Lambda})-\mathrm{n_{-}}(L_{c,\Lambda})=-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda,m_{\sigma,M},-\Omega\oplus\Omega,\sigma\in[c,b])+\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda,m_{\sigma,0},\sigma\in[c,b])\\ =-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda,m_{\sigma,M},-\Omega\oplus\Omega,\sigma\in[c,b])+\sum_{\sigma\in(c,b)}\dim\Big((\Lambda_{0}\oplus\Lambda_{D})\cap m_{\sigma}\Big).$ (5.2)
Since $L_{b,\Lambda}$ is bounded from below, then for $M>0$ large enough, we get that $\mathrm{n_{-}}(L_{b,\Lambda}+MI)=0$. By Proposition 5.7, then we get that
$\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{0}\oplus\Lambda_{D},m_{\sigma,M},\sigma\in[c,b])=\sum_{\sigma\in(c,b)}\dim\big((\Lambda_{0}\oplus\Lambda_{D})\cap m_{\sigma,M}\big)=\sum_{\sigma\in(c,b)}\dim\ker(L_{(a,\sigma),\Lambda}+MI).$
We also observe that for $M$ large enough, we can assume that $\mathrm{n_{-}}(L_{b,\Lambda}+(M-1)I)=0$. By Lemma 5.8, we get that
$\mathrm{n_{-}}(L_{(a,\sigma),\Lambda}+(M-1)I)\leq\mathrm{n_{-}}(L_{b,\Lambda}+(M-1)I)=0\qquad\sigma\in[c,b].$ (5.3)
This implies that $\dim\ker(L_{(a,\sigma),\Lambda}+MI)=\{0\}$, (if not, the corresponding non-trivial vector in the kernel contributes to the number $\mathrm{n_{-}}(L_{(a,\sigma),\Lambda}+(M-1)I)$). So, we get that for $M$ sufficiently large,
$\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{0}\oplus\Lambda_{D},m_{\sigma,M},-\Omega\oplus\Omega,\sigma\in[c,b])=\sum_{\sigma\in(c,b)}\dim\ker(L_{(a,\sigma),\Lambda}+MI)=0.$ (5.4)
By Equation (5.2) and Equation (5.4) we get that
$\mathrm{n_{-}}(L_{b,\Lambda})-\mathrm{n_{-}}(L_{c,\Lambda})=\sum_{\sigma\in(c,b)}\dim\big((\Lambda_{0}\oplus\Lambda_{D})\cap m_{\sigma}\big).$ (5.5)
Passing to the limit for $c$ converging to $a$ at Equation (5.5), we get
$\mathrm{n_{-}}(L_{b,\Lambda})\geq\sum_{\sigma\in(a,b)}\dim\big((\Lambda_{0}\oplus\Lambda_{D})\cap m_{\sigma}\big).$ (5.6)
By Lemma 5.8, there exists $d>a$ such that $\mathrm{n_{-}}(L_{d,\Lambda})=0$ and so, we get
$\mathrm{n_{-}}(L_{b,\Lambda})=\sum_{\sigma\in(d,b)}\dim\big((\Lambda_{0}\oplus\Lambda_{D})\cap m_{\sigma}\big)\leq\sum_{\sigma\in(a,b)}\dim\big((\Lambda_{0}\oplus\Lambda_{D})\cap m_{\sigma}\big).$ (5.7)
Summing up the Equation (5.6) and Equation (5.7), Theorem 5.1 is proved. ∎
6 Morse Index Theorem for General Boundary Conditions
The purpose of this section is twofold. First, we extend Theorem 5.1 to the setting of general Lagrangian boundary conditions. The main result, stated in Theorem 6.1, expresses the difference between the Morse indices of two self-adjoint Fredholm operators—obtained as the Friedrichs extensions of the corresponding quadratic forms—in terms of the triple index.
Second, we establish that a perturbation of a two-sided singular Sturm–Liouville operator $L$, which is bounded from below, may fail to remain bounded from below when the perturbation is given by a symmetric but not relatively bounded operator. This analysis sheds light on a phenomenon already noted by Rellich, wherein eigenvalues may “vanish” by escaping to -$\infty$; see Kato [Kat80, Ch. V, §4, Thm. 4.1] for a precise formulation.
The difference between two operators subject to different boundary conditions has been investigated in [HWY20] and [HPWX20], both for regular Sturm–Liouville operators and for Sturm–Liouville operators defined on the entire real line. In [BCLS24, Theorem 1.2], this result has been further generalized to arbitrary symmetric operators.
Here, by employing an approach analogous to that developed in [HWY20], we establish a difference formula for general Sturm–Liouville operators subject to distinct boundary conditions.
For the reader’s convenience, the proofs of the foundational results concerning Friedrichs extensions are deferred to Appendix C.
Theorem 6.1.
Let $L$ be a one-sided singular the Sturm–Liouville operator defined at Equation (3.1), and suppose that conditions (H3)–(H4) hold. Then we have:
$\mathrm{n_{-}}(L_{\Lambda})-\mathrm{n_{-}}(L_{F})=\iota\bigl(p(\ker L^{*}),\,p(D(L_{\Lambda})),\,p(D(L_{F}))\bigr),$
where $\Lambda\in\operatorname{\mathrm{Lag}}(W,\omega|_{W})$, $\iota$ denotes the triple index (Cfr. Appendix A and references therein).
Proof.
Without loss of generality. We can assume that $D(L^{*})=D(L)\oplus W$ such that $\ker L^{*}<U$, $\Lambda\in\operatorname{\mathrm{Lag}}(W,\rho)$, $F:=D(L_{F})\cap W\in\operatorname{\mathrm{Lag}}(W,\rho)$ and $L_{\Lambda}=L^{*}|_{D(L)\oplus\Lambda}$ . We start to observe that by Lemma D.11, we get that $\mathrm{n_{-}}(L_{\Lambda})=\mathrm{n_{-}}(t_{L_{\Lambda}})$ and $\mathrm{n_{-}}(L_{F})=\mathrm{n_{-}}(t_{L})$. By Lemma 2.6, we get that $\ker L^{*}$ is a Lagrangian subspace of $W$ and by invoking Lemma D.10 for $G=D(t_{L})=D(t_{L_{F}})$, $V=D(t_{L_{\Lambda}})$ and $Q=t_{L_{\Lambda}}$, to compute the difference of the Morse indices, it’s enough to compute
$\dim(D(t_{L})\cap(D(t_{L}))^{t_{L_{\Lambda}}}+\ker t_{L_{\Lambda}})/\ker t_{L_{\Lambda}}.$
By Lemma C.6 and Corollary C.4, we get that $\ker t_{L_{\Lambda}}=\ker L_{\Lambda}$. Then we have
$D(t_{L})\cap(D(t_{L}))^{t_{L_{\Lambda}}}=(D(t_{L}))^{t_{L}}=\ker L_{F}$ (6.1)
We get
$\dim\left[(D(t_{L})\cap(D(t_{L}))^{t_{L_{\Lambda}}}+\ker t_{L_{\Lambda}})/\ker t_{L_{\Lambda}}\right]=\dim\left[(\ker L_{F}+\ker L_{\Lambda})/\ker L_{\Lambda}\right]$
$=\dim\left[(\ker L_{F})/(\ker L_{F}\cap\ker L_{\Lambda})\right]=\dim(\ker L^{*}\cap F)-\dim(\ker L^{*}\cap F\cap\Lambda).$
Then by Lemma D.12 and Lemma D.10, we finally get
$\mathrm{n_{-}}(L_{\Lambda})-\mathrm{n_{-}}(L_{F})=\iota(\ker L^{*},\Lambda,F).$
This concludes the proof. ∎
Corollary 6.2.
Consider a decomposition $D(L^{*})=D(L)\oplus U$ with $\ker L^{*}<U$. Let $\Lambda_{0},\Lambda_{1}\in\operatorname{\mathrm{Lag}}(U,\rho)$. Let $F=U\cap D(L_{F})$ where $L_{F}$ is the Friedrich extension of $L$. Then we have
$\mathrm{n_{-}}(L_{\Lambda_{1}})-\mathrm{n_{-}}(L_{\Lambda_{0}})=\iota(\ker L^{*},\Lambda_{1},\Lambda_{0})-\iota(\Lambda_{1},\Lambda_{0},F).$
Proof.
By Theorem 6.1, we get that
$\mathrm{n_{-}}(L_{\Lambda_{i}})-\mathrm{n_{-}}(L_{F})=\iota(W,\Lambda_{i},F)\qquad\textrm{ for }\ i=0,1.$
By this identity and by using Lemma A.12 and Proposition A.9, we get that
$\mathrm{n_{-}}(L_{\Lambda_{1}})-\mathrm{n_{-}}(L_{\Lambda_{0}})$ $=\iota(W,\Lambda_{1},F)-\iota(\ker L^{*},\Lambda_{0},F)$ (6.2)
$=[\iota(F,\ker L^{*},\Lambda_{1})-\dim(F\cap\Lambda_{1})]-[\iota(F,\ker L^{*},\Lambda_{0})-\dim(F\cap\Lambda_{0})]$ (6.3)
$=s(F,\ker L^{*},\Lambda_{0},\Lambda_{1})-\dim(F\cap\Lambda_{1})+\dim(F\cap\Lambda_{0})$ (6.4)
By invoking once again Proposition A.9, we get that
$s$ $(F,\ker L^{*},\Lambda_{0},\Lambda_{1})-\dim(F\cap\Lambda_{1})+\dim(F\cap\Lambda_{0})$
$=-s(F,\ker L^{*},\Lambda_{1},\Lambda_{0})-\dim(F\cap\Lambda_{1})+\dim(F\cap\Lambda_{0})$
$=-\iota(F,\Lambda_{1},\Lambda_{0})+\iota(\ker L^{*},\Lambda_{1},\Lambda_{0})-\dim(F\cap\Lambda_{1})+\dim(F\cap\Lambda_{0})$
$=\iota(\ker L^{*},\Lambda_{1},\Lambda_{0})-\iota(\Lambda_{1},\Lambda_{0},F).$
The last equation comes from Lemma A.12. ∎
We close this section with an additional discussion about infinite Morse index.
Proposition 6.3.
Let $\Lambda=\Lambda_{0}\oplus\Lambda_{D}$ and we assume conditions (H3) & (H4). Then $\mathrm{n_{-}}(L_{\Lambda})=+\infty$ if and only if
$\sum_{d\in(a,b)}\dim\ker L_{(a,d),\Lambda_{0}\oplus\Lambda_{D}}=+\infty.$
Proof.
We start by observing that the map $\operatorname{\mathcal{T}}_{a}:\ker L_{b}^{*}\to{\mathbb{R}}^{2k-2n}$ is a surjection. Then we have
$\dim\big[\operatorname{\mathcal{T}}^{-1}(\Lambda_{0}\oplus{\mathbb{R}}^{2n})\cap\ker L_{b}^{*}\big]=\dim\big[\operatorname{\mathcal{T}}_{a}^{-1}(\Lambda_{0})\cap\ker L_{b}^{*}\big]=\dim\ker L_{b}^{*}-\operatorname{codim}(\Lambda_{0})\\ =\dim\ker L_{b}^{*}-(2k-2n-\dim\Lambda_{0})=k-(2k-2n-(k-n))=n.$ (6.5)
By the existence and uniqueness theorem for solution of linear odes, it follows that $\operatorname{\mathcal{T}}_{b}:\ker L_{b}^{*}\to{\mathbb{R}}^{2n}$ is an injection and in particular $\Gamma:=\operatorname{\mathcal{T}}_{b}(\operatorname{\mathcal{T}}_{a}^{-1}(\Lambda)\cap\ker L_{b}^{*})$ is a $n$-dimensional subspace of $({\mathbb{R}}^{2n},\omega_{b})$.
We are now ready to prove that $U$ is a Lagrangian subspace of $({\mathbb{R}}^{2n},\omega_{b})$. We only need to show that it is isotropic. Let $u_{1},u_{2}\in\Gamma$ with $u_{1}=f_{1}(b),u_{2}=f_{2}(b)$ where $f_{1},f_{2}\in\operatorname{\mathcal{T}}_{a}^{-1}(\Lambda)\cap\ker L_{b}^{*}$. We have
$0=\omega(f_{1},f_{2})=[f_{1},f_{2}](b)-[f_{1},f_{2}](a+).$
Since $\operatorname{\mathcal{T}}_{a}(f_{1}),\operatorname{\mathcal{T}}_{a}(f_{2})\in\Lambda$ which is a Lagrangian subspace, we have $f_{1},f_{2}](a+)=0$. Then we get $f_{1},f_{2}](b)=0$ which imply that $\omega_{b}(u_{1},u_{2})=0$. Then $\Gamma$ is isotropic and being $n$-dimensional then $\Gamma\in\operatorname{\mathrm{Lag}}({\mathbb{R}}^{2n},\omega_{b})$. With these notation, we have
$\dim\ker L_{(a,c),\Lambda_{0}\oplus\Lambda_{D}}=\dim\ker L_{(c,b),\Lambda_{D}\oplus\Gamma}$
Then we can conclude that
$\sum_{d\in(c,b)}\dim\ker L_{(a,d),\Lambda_{0}\oplus\Lambda_{D}}=\sum_{d\in(c,b)}\dim\ker L_{(d,b),\Lambda_{D}\oplus\Gamma}.$
By Theorem 5.2, then we have
$\mathrm{n_{-}}(L_{(c,b),\Lambda_{D}\oplus\Gamma})=\sum_{d\in(c,b)}\dim\ker L_{(d,b),\Lambda_{D}\oplus\Gamma}=\sum_{d\in(c,b)}\dim\ker L_{(a,d),\Lambda_{0}\oplus\Lambda_{D}}.$
[…]
$\lim_{c\to a^{+}}\mathrm{n_{-}}(L_{(c,b),\Lambda_{D}\oplus\Lambda_{D}})=+\infty$
By invoking Theorem 6.1, we get that
$\lim_{c\to a}\mathrm{n_{-}}(L_{(c,b),\Lambda_{D}\oplus\Gamma})=+\infty.$
That is
$\lim_{c\to a^{+}}\sum_{d\in(c,b)}\dim\ker L_{(a,d),\Lambda_{0}\oplus\Lambda_{D}}=\sum_{d\in(a,b)}\dim\ker L_{(a,d),\Lambda_{0}\oplus\Lambda_{D}}=+\infty.$
7 About a Rellich counterexample to a spectral flow formula
In this section, we use the abstract index theory to examine Rellich’s classical counterexample as presented in Kato [Kat80, Chapter V, Section 4, Remark 4.13].
It is well known that if $T$ is a closed, self‑adjoint operator bounded below and $A$ is a relatively bounded symmetric perturbation with sufficiently small bound, then the operator
$S=T+A$
remains self‑adjoint and bounded below (see Kato [Kat80, Chapter V, Section 4, Theorem 4.1]). However, this conclusion can fail when the perturbation is not relatively bounded: Rellich’s scalar counterexample shows that, in such cases, certain eigenvalues may “disappear at -$\infty$.” It is still an open problem to determine exactly how many eigenvalues can vanish in this way. The purpose of this section is to address that question and to provide a precise measure of the multiplicity of eigenvalues that escape to -$\infty$.
We start by recalling a well-known fact about a continuous path of self-adjoint Fredholm operators bounded from below, defined on the same domain and parametrized by a compact interval. In this case what the spectral flow does it to measure the difference of the Morse indices of the operators at the boundary points of the interval. More precisely, let $s\mapsto A_{s}$ be a continuous path of self-adjoint Fredholm operators having fixed domain. Then the spectral flow $\operatorname{sf}(A_{s},s\in[0,1])$ is the net number of eigenvalue crossing the 0 when the parameter $s$ runs from 0 to $1$ whilst $\mathrm{n_{-}}(A_{1})-\mathrm{n_{-}}(A_{0})$ measures the difference of the negative spectral spaces between the operator at the instant $1$ and the one at 0. Under the above condition, we get that
$\mathrm{n_{-}}(A_{1})-\mathrm{n_{-}}(A_{0})+\operatorname{sf}(A_{s},s\in[0,1])=0.$ (7.1)
Things drastically change if the domains of the operators are not anymore constant. This is, for instance, the case of a gap-continuous path of self-adjoint extensions $s\mapsto A_{\Lambda_{s}}$ of an operator $A\in\operatorname{\mathcal{CF}^{s}}(H)$ for a gap-continuous path $s\mapsto\Lambda_{s}$ of Lagrangian subspaces of the symplectic space $(U,\rho))$. In this case, the domain is not anymore fixed and so, the operators are not anymore relative bounded perturbation of a fixed one.
In this new situation, the Equation (7.1) doesn’t hold anymore! The lack of the equality is very much related to the phenomenon observed by Rellich through a counterexample to the equality provided at Equation (7.1) in the case of Sturm-Liouville operators on $[0,1]$ with varying domains.
In the next result we provide a sharp formula measuring the difference of the Morse indices of the operators $L_{\Lambda_{1}}$ and $L_{\Lambda_{0}}$ in terms of the spectral flow of the associated path $s\mapsto L_{\Lambda_{s}}$ and of the Maslov index of the Lagrangian path. Without loss of generality, we will consider the decomposition $D(L^{*})=D(L)\oplus W$ with $\ker L^{*}<W$.
We let
$V:=\ker L^{*}\quad\textrm{ and }\quad\Lambda_{F}=D(L_{F})\cap W.$
Theorem 7.1.
Assume conditions (H3) & (H4). If $\mathrm{n_{-}}(L_{\Lambda_{0}})<\infty$, then we get:
$\mathrm{n_{-}}(L_{\Lambda_{1}})-\mathrm{n_{-}}(L_{\Lambda_{0}})+\operatorname{sf}(L_{\Lambda_{s}},s\in[0,1])=\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{F},\Lambda_{s},\rho,s\in[0,1]).$
Proof.
By Theorem 6.1, we get that
$\mathrm{n_{-}}(L_{\Lambda_{s}})-\mathrm{n_{-}}(L_{F})=\iota(p(V),\Lambda_{s},\Lambda_{F}).$
By invoking Lemma A.12 and Proposition A.9, we get that
$\mathrm{n_{-}}(L_{\Lambda_{1}})-\mathrm{n_{-}}(L_{\Lambda_{0}})$ $=\iota(p(V),\Lambda_{1},\Lambda_{F})-\iota(p(V),\Lambda_{0},\Lambda_{F})$ (7.2)
$=[\iota(\Lambda_{F},p(V),\Lambda_{1})-\dim(\Lambda_{F}\cap\Lambda_{1})]-[\iota(\Lambda_{F},p(V),\Lambda_{0})-\dim(\Lambda_{F}\cap\Lambda_{0})]$ (7.3)
$=s(\Lambda_{F},p(V),\Lambda_{0},\Lambda_{1})-\dim(\Lambda_{F}\cap\Lambda_{1})+\dim(\Lambda_{F}\cap\Lambda_{0})$ (7.4)
where we denoted by $s(\Lambda_{F},p(V),\Lambda_{0},\Lambda_{1})$ the Hörmander index of the Lagrangian quadruple. (Cf. Appendix A). Now, by using the spectral flow formula proved at Theorem 3.3:
$\operatorname{sf}(L_{\Lambda_{s}},s\in[0,1])=-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{s},p(V),\rho,s\in[0,1])$
and since of Definition A.3, we get that
$\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{s},p(V),\rho,s\in[0,1])-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{s},\Lambda_{F},\rho,s\in[0,1])=s(\Lambda_{F},p(V),\Lambda_{0},\Lambda_{1}).$
Summing up we conclude that
$\mathrm{n_{-}}(L_{\Lambda_{1}})-\mathrm{n_{-}}(L_{\Lambda_{0}})+\operatorname{sf}(L_{\Lambda_{s}},s\in[0,1])\\ =s(\Lambda_{F},p(V),\Lambda_{0},\Lambda_{1})-\dim(\Lambda_{F}\cap\Lambda_{1})+\dim(\Lambda_{F}\cap\Lambda_{0})-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{s},p(V),\rho,s\in[0,1])\\ =-\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{s},\Lambda_{F},\rho,s\in[0,1])-\dim(\Lambda_{F}\cap\Lambda_{1})+\dim(\Lambda_{F}\cap\Lambda_{0})\\ \Lambda_{F},\Lambda_{s},\rho,s\in[0,1]).$ (7.5)
This concludes the proof. ∎
The reason behind Theorem 7.1 is due to the fact that some negative eigenvalue disappearing at -$\infty$. The total number of such a kind of ghost eigenvalues can be estimated in terms of the Maslov index.
Theorem 7.2.
We Assume condition (H3) and if $\Lambda_{s}\cap\Lambda_{F}=(0)$ for $s\neq 0$, then there exists $K>0$ such that for every $M>K$ there exists $\varepsilon>0$ such that the following holds:
$((-\infty,-M)\cap\mathfrak{sp}(L_{\Lambda_{u}}))=\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{F},\Lambda_{s},\rho,s\in[0,\epsilon])\qquad u\in(0,\varepsilon),$ (7.6)
where $\mathfrak{sp}$ denotes the spectrum.
Remark 7.3.
Before proving this result, some remarks are in order. We assume that for $\varepsilon$ sufficiently small, the (RHS) at Equation (7.6) is not-zero. So, since $\mu^{\scriptscriptstyle{\mathrm{CLM}}}$ is a continuous integer-valued function, it is locally constant. By this we get that for for every $u\in(0,\varepsilon)$ the operator $L_{\Lambda_{u}}$ has $\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{D},\Lambda_{s},s\in[0,u])=l\neq 0$ number of negative eigenvalues less that M-M. By choosing $u$ arbitrarily close to 0, we get that the $l$ eigenvalues are arbitrarily negative and in the limit for $u\to 0^{+}$, they diverge to -$\infty$. So, the constant $K$ at the Theorem 7.2 decompose the spectrum of $L_{\Lambda_{u}}$ into two parts:
  1. 1.
    The upper part of the spectrum given by the eigenvalues greater than $K$
  2. 2.
    The lower part of the spectrum given by the eigenvalues less than $K$.
The eigenvalues belonging to the lower part of the spectrum, in the limit, diverges to -$\infty$. The eigenvalues belonging to the upper part of the spectrum, stay uniformly bounded by $K$.
Proof.
Choose $K>0$ such that $L_{\Lambda_{0}}+KI$ is a positive and Fredholm operator. Since $s\mapsto L_{\Lambda_{s}}+KI$ is continuous, then there exists $\varepsilon>0$ such that $\ker(L_{\Lambda_{u}}+KI)=\set{0}$ for every $u\in[0,\varepsilon]$. By this we get that
$\operatorname{sf}(L_{\Lambda_{r}}+KI,r\in[0,u])=0\qquad\textrm{ for every }u\in[0,\varepsilon].$
By Theorem 7.1, we get
$\mathrm{n_{-}}(L_{\Lambda_{u}}+KI)=\mathrm{n_{-}}(L_{\Lambda_{u}}+KI)-\mathrm{n_{-}}(L_{\Lambda_{0}}+KI)=\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{F},\Lambda_{s},\rho,s\in[0,u]).$
Since $\Lambda_{F}\cap\Lambda_{s}=(0)$ and since $s\neq 0$, we have
$\mathrm{n_{-}}(L_{\Lambda_{u}}+KI)=\mathrm{n_{-}}(L_{\Lambda_{u}}+KI)-\mathrm{n_{-}}(L_{\Lambda_{0}}+KI)=\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{F},\Lambda_{s},\rho,s\in[0,\epsilon])\qquad u\in(0,\varepsilon).$
So, the number of negative eigenvalues of $L_{\Lambda_{u}}$ is $\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{F},\Lambda_{s},\rho,s\in[0,\varepsilon])$ for every $u\in(0,\varepsilon)$.
For $M>K$, the operator $L_{\Lambda_{0}}+MI>0$. So, with the same reason, we can choose $\epsilon$ and get the same equation. Then, the result follows. ∎
Corollary 7.4.
Assume condition (H3) and if $\Lambda_{s}\cap\Lambda_{F}=(0)$ for $s\neq 0$. Then there is $K>0$ and $a>0$ such that
$((-\infty,-K)\cap\mathfrak{sp}(L_{\Lambda_{u}}))=\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{F},\Lambda_{s},\rho,s\in[0,a])\qquad\forall\,u\in(0,a].$
Let $k=\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{F},\Lambda_{s},\rho,s\in[0,a])$. Then for each $u\in(0,a]$, there are precisely $k$ eigenvalues of $L_{\Lambda_{u}}$ less than K-K, namely $\lambda_{1}(u)\leq,\cdots,\leq\lambda_{k}(u)$. Then we have
$\lim_{u\to 0^{+}}\lambda_{i}(u)=-\infty\qquad 1\leq i\leq k.$
Proof.
By Theorem 7.2, there is $K>0,a>0$
$((-\infty,-K)\cap\mathfrak{sp}(L_{\Lambda_{u}}))=\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{F},\Lambda_{s},\rho,s\in[0,a])\qquad\forall\,u\in(0,a].$
For any $M>K$, it is possible to choose $0<a^{\prime}<a$ such that
$((-\infty,-M)\cap\mathfrak{sp}(L_{\Lambda_{u}}))=\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\Lambda_{F},\Lambda_{s},\rho,s\in[0,a])\qquad\forall\,u\in(0,a^{\prime}].$
So, for $u\in(0,a^{\prime})$ we have
$(-\infty,-K)\cap\mathfrak{sp}(L_{\Lambda_{u}})=(-\infty,-M)\cap\mathfrak{sp}(L_{\Lambda_{u}}).$
By this it follows that $\lambda_{i}(u)<-M$ for $1\leq i\leq k$ and for $u\in(0,a^{\prime})$. Then we get
$\lim_{u\to 0^{+}}\lambda_{i}(u)=-\infty\qquad 1\leq i\leq k.$
This concludes the proof. ∎
8 Bessel-type differential operators
The aim of this section is to establish the functional-analytic and spectral properties of Bessel-type differential operators, and to provide sufficient conditions under which the index theory developed below applies.
We consider the unperturbed Bessel-type differential operator $h_{q}$ on $\mathscr{C}_{0}^{\infty}((0,1],{\mathbb{R}}^{n})$ defined by
$h_{q}:=-\dfrac{d^{2}}{dt^{2}}+\dfrac{q}{t^{2}}\quad\textrm{ for }\quad t\in(0,1)\textrm{ and }\quad q<3/4$
and we are interested in finding the solution space of $h_{q}$. A convenient way to do so, is to reparametrize the differential operator $h_{q}$ through the following function $r\mapsto q(r)$ where
$q(r)=\begin{cases}-1/4+r^{2}&r\geq 0\\[3.0pt] -1/4-r^{2}&r<0.\end{cases}$ (8.1)
So, in this case the operator $h$ can be written as
$h_{r}=-\dfrac{d^{2}}{dt^{2}}+\dfrac{q(r)}{t^{2}},\qquad t\in(0,1).$
and we observe that the function $q$ is a bijection from $(-\infty,1)$ to $(-\infty,3/4)$:
2-21.5-1.51-10.5-0.50.50.5111.51.5222-222rrq(r)q(r)$q(r)$
Remark 8.1.
The function $r\mapsto q(r)$ controls the strength of the singularity in the operator $L_{r}$. For $r\geq 0$, the potential is repulsive near $t=0$; for $r<0$, it becomes increasingly attractive. The threshold q(r)=1/4q(r)=-1/4 marks a transition in the limit-point/limit-circle classification, leading to qualitative changes in the spectrum. In particular, for large negative $r$, the operator may lose semi-boundedness and develop diverging negative eigenvalues. (Cf. Section 7).
By a direct computation we get that this equation has two linearly independent solutions in $H=L^{2}([0,1],{\mathbb{R}}^{n})$ given by
$y_{1,r}(t)=\dfrac{1}{2}(t^{1/2-r}+t^{1/2+r})$ $y_{2,r}(t)=\dfrac{1}{2\,r}(t^{1/2+r}-t^{1/2-r})$ $\textrm{ for }r\in(0,1)$ (8.2)
$y_{1,r}(t)=t^{1/2}$ $y_{2,r}(t)=t^{1/2}\ln t$ $\textrm{ for }r=0$ (8.3)
$y_{1,r}(t)=t^{1/2}\cos(r\ln t)$ $y_{2,r}(t)=\dfrac{1}{r}t^{1/2}\sin(r\ln t)$ $\textrm{ for }r\in(-\infty,0).$ (8.4)
By a direct calculation, it readily follows that
$y_{1,r},y_{2,r}](0)=-1.$
8.1 Functional properties of the Bessel-type operator
The aim of this paragraph is to prove the Fredholmness of the Bessel differential operator and to study its stability properties under bounded perturbation. We start with two abstract results. The first result gives a sufficient condition for any self-adjoint extension of a SL-operator to have compact resolvent. The second is an abstract functional analytic result about the stability property of an operator to remaining a compact resolvent operator once perturbed by a bounded one.
Lemma 8.2.
Let $l$ be a Sturm–Liouville operator defined on $\mathscr{\mathscr{}}C^{\infty}([0,1],{\mathbb{R}}^{n})$ and we assume that the operator $l$ is regular at $t=1$.
If $\dim\ker L^{*}=2n$, then $L$ is a symmetric Fredholm operators. Furthermore, any self-adjoint extension of $L$ has compact resolvent.
Proof.
For a Sturm Liouville equation $\dfrac{d}{dt}(P(t)\dfrac{d}{dt}+Q(t))u+{Q}^{\mathsf{T}}\dfrac{d}{dt}u+R(t)u=0$. Let $M(t)$ be the monodromy matrix of the Hamiltonian system induced by the Sturm-Liouville operator $l$ such that $M(1)=I$. Let $M(t)=\begin{bmatrix}U(t)\\ V(t)\end{bmatrix}$. Then $V(t)=(v_{1}(t),\cdots,v_{2n}(t))$ where $\set{v_{1},\cdots,v_{2n}}$ is a basis of $\ker L^{*}$. Now we solve the equation $Lv=f$ with $f\in\operatorname{im}(L)$. Since that $M(t)$ is a symplectic matrix, then we have
$\begin{pmatrix}v^{[1]}\\ v\end{pmatrix}(t)=M(t)c+M(t)\int_{1}^{t}M(s)^{-1}J\begin{pmatrix}0\\ f\end{pmatrix}ds=M(t)c+M(t)\int_{1}^{t}J{M}^{\mathsf{T}}(s)\begin{pmatrix}0\\ f\end{pmatrix}ds$
for some constant vector $c$. Since $v\in D(L)$ so $c=0$. We finally get
$v(t)=(L^{-1}f)(t)=\int_{1}^{t}V(t)J{V}^{\mathsf{T}}(s)f(s)\,ds\qquad\text{for }f\in\operatorname{im}L.$
Since $V\in L^{2}([0,1],{\mathbb{R}}^{n\times 2n})$, $L^{-1}$ is bounded on $\operatorname{im}(L)<L^{2}([0,1],{\mathbb{R}}^{n})$. Since $L$ is closed, $L^{-1}$ is also closed. So, $\operatorname{im}(L)=D(L^{-1})$ is closed since $L^{-1}$ is bounded. Furthermore, $L^{-1}$ is an integral operator with $L^{2}$ kernel, so $L^{-1}$ is compact. This concludes the proof. ∎
Lemma 8.3.
Let $A$ be a self-adjoint operator having compact resolvent. For each bounded operator $B$, $A+B$ also has compact resolvent.
Proof.
Note that $\|(A+i\lambda I)^{-1}\|\leq 1/\lambda$ for $\lambda>0$. Choose $\lambda$ large enough. We can assume that $\|B(A+i\lambda I)^{-1}\|<1$. We have
$(A+B+i\lambda I)^{-1}=(A+i\lambda I)^{-1}(I+B(A+i\lambda I)^{-1})^{-1},$
which is compact. ∎
Corollary 8.4.
Let us consider the Bessel operator
$l_{q}:=-\dfrac{d^{2}}{dt^{2}}+\dfrac{q}{t^{2}}$
acting on $C_{0}^{\infty}((0,1],{\mathbb{R}}^{n})$ and we denote by $L_{q}:D(L_{q})<H\longrightarrow H$ the corresponding minimal operator where $H=L^{2}([0,1],{\mathbb{R}}^{n})$.
Assume that $q<3/4$ and let $B\in\mathcal{B}^{sa}(H)$. Then any self-adjoint extension of $L_{q}+B$ is a compact resolvent operator.
Proof.
By Lemma 8.2 and Lemma 8.3, we only need to compute the $\dim\ker L_{q}^{*}$. Such a computation can be reduced to the scalar case. So, by the Equation (8.2), we get that the kernel dimension in the scalar case is 2. By this, we conclude that $\dim\ker L_{q}^{*}=2n$. ∎
The next step is to study the Fredholmness properties of the operator $L_{q}$. For $q\geq 3/4$, this will be done by using a result of Naimark (cf. [Nai68]) which is known in literature as splitting method. Here for the sake of the reader, we write it in the case of Sturm-Liouville operators, we are interested in.
Lemma 8.5 (Naimark, splitting method).
Let $l$ be the SL-operator defined on $C_{0}^{\infty}((a,b),{\mathbb{R}}^{n})$. Let $c\in(a,b)$. Let $l_{1},l_{2}$ be the restriction of $l$ on $C_{0}^{\infty}((a,c),{\mathbb{R}}^{n})$ and $C_{0}^{\infty}((c,b),{\mathbb{R}}^{n})$ respectively.
Let $L$ be the corresponding minimal operator associated to $l$. Let $L_{1}$ $L_{2}$ be the corresponding minimal operators associated to $l_{1},l_{2}$, respectively. The continuous part of the spectrum of every self-adjoint extension of the operator $L$ is the union of the continuous parts of the spectra of any self-adjoint extensions of $L_{1}$ and $L_{2}$
Proof.
For the proof we refer the interested reader to [Nai68, Theorem 24.1.1]
Proposition 8.6.
Given $R\in\mathscr{C}^{0}([0,1],\mathrm{S}(n))$, we define $l_{R}$ to be the differential operator acting on $\mathscr{C}_{0}^{\infty}((0,1),{\mathbb{R}}^{n})$ and given by
$l_{R}:=-\dfrac{d^{2}}{dt^{2}}+\dfrac{R(t)}{t^{2}}\qquad t\in(0,1)$
and we assume that $\sigma(R(0))\subset(-1/4,+\infty)$. Denoting by $L_{R}$ (resp. $L_{R}^{*}$) the minimal (resp. maximal) operator induced by $l_{R}$, then, for every $D$ such that $D(L_{R})<D<D(L_{R}^{*})$ and for every $\lambda\in{\mathbb{C}}$ we get that the $(L_{R}+\lambda I)|_{D}$ is a Fredholm operator.
Proof.
The proof of this result readily follows once we prove that
$\sigma_{\mathrm{ess}}(L_{R})=\emptyset.$
To do so, we apply Lemma 8.5 to the interval $(0,1)=(0,\delta)\cup(\delta,1)$. We denote by $L_{R}^{1}$ (resp. $L_{R}^{2}$) the minimal operator corresponding to the differential operator $l_{R}^{1}$ (resp. $l_{R}^{2}$) obtained by restricting $l_{R}$ on $\mathscr{C}_{0}^{\infty}((0,\delta),{\mathbb{R}}^{n})$ (resp. $\mathscr{C}_{0}^{\infty}((\delta,1),{\mathbb{R}}^{n})$). Let $\phi\in C^{\infty}_{0}((0,1),{\mathbb{R}}^{n})$. By Hardy’s inequality, we have
$\langle L_{R}\phi,\phi\rangle=\langle\dot{\phi},\dot{\phi}\rangle+\left\langle\dfrac{R(t)}{t^{2}}\phi,\phi\right\rangle\geq\left\langle\dfrac{1/4+R(t)}{t^{2}}\phi,\phi\right\rangle.$
We note that $\min(\sigma(R(t))$ is continuous on $[0,1]$. So, there exists $\delta>0$ such that $\min(\sigma(R(t))\geq-1/4+c$ on $0,\delta]$, for some $c>0$. It follows that
$\langle L_{R}\phi,\phi\rangle\geq c\left\|\dfrac{\phi}{t}\right\|^{2}\geq\dfrac{c}{\delta^{2}}\|\phi\|^{2}\qquad\textrm{ for any }\qquad\phi\in\mathscr{C}^{\infty}_{0}((0,\delta),{\mathbb{R}}^{n}).$
Let $L^{i}_{R,F}$ be the Friedrich extensions of $L_{R}^{i}$ for $i=1,2$. It follows that
$\sigma_{\mathrm{ess}}(L^{1}_{R,F})\subset\sigma(L^{1}_{R,F})\subset\left[\dfrac{c}{\delta^{2}},+\infty\right).$
Since $l_{R}$ is regular on $\delta,1]$, we have $\sigma_{\mathrm{ess}}(L^{2}_{R,F})=\emptyset$. Therefore,
$\sigma_{\mathrm{ess}}(L^{1}_{R,F}\oplus L^{2}_{R,F})\subset\left[\dfrac{c}{\delta^{2}},+\infty\right).$
By Lemma 8.5, we conclude that
$\sigma_{\mathrm{ess}}(L_{R})\subset\left[\dfrac{c}{\delta^{2}},+\infty\right)\qquad\forall\,\delta>0.$
Then $\sigma_{\mathrm{ess}}(L_{R})=\emptyset$. The proof now follows by the arbitrariness of $\delta$. ∎
Theorem 8.7.
Let $b\in\mathscr{C}^{0}([0,1],\mathrm{S}(n))$ and let $L_{q}$ be the minimal operator of the Bessel operator defined at Corollary 8.4. Denoting by $L_{b}$ the operator pointwise defined by $b$, then for any $q\in{\mathbb{R}}$ we get that any self-adjoint extension of $L_{q}+L_{b}$ is a Fredholm operator.
Proof.
We let $R(t)=q\,I+t^{2}\,b(t)$. If
  • q>1/4q>-1/4, by invoking Proposition 8.6, we get that any self-adjoint extension of $L_{q}+L_{b}$ is Fredholm
  • $q\leq-1/4$, then by invoking Corollary 8.4 (which works for $q<3/4$), we get that $L_{q}+L_{b}$ has a compact resolvent and in particular it is Fredholm.
This concludes the proof. ∎
Let us now consider the Bessel-type operator on $C_{0}^{\infty}((1,+\infty),{\mathbb{R}}^{n})$ defined by
$l_{R}:=-\dfrac{d^{2}}{dt^{2}}+\dfrac{R(t)}{t^{2}}\qquad\textrm{ for }\qquad t\in[1,+\infty).$
We observe that if $\lim_{t\to\infty}t^{-2}R(t)=0$, then $l_{R}$ is not Fredholm. Nevertheless, since the essential spectrum is contained in $0,+\infty)$, it is possible to study its Morse index.
Theorem 8.8.
Let $l_{R}$ be the operator defined above on $\mathscr{C}_{0}^{\infty}((1,+\infty),{\mathbb{R}}^{n})$.
  • Case 1. If $\liminf_{t\to+\infty}R(t)>-1/4$, then any self-adjoint extension of it has finite Morse index.
  • Case 2. If $\limsup_{t\to+\infty}R(t)<-1/4$, then any self-adjoint extension of it has infinite Morse index.
Proof.
Let $M>1$. By Using splitting method, let $l_{1}=l|_{C_{0}^{\infty}((1,M),{\mathbb{R}}^{n})}$ and $l_{2}=l|_{C_{0}^{\infty}((M,+\infty),{\mathbb{R}}^{n})}$ . Let $L,L_{1},L_{2}$ be the minimal operators corresponding to $l,l_{1},l_{2}$, respectively. Let $L_{D},L_{1,D},L_{2,D}$ be the self-adjoint operators associated to $l,l_{1},l_{2}$ with Dirichlet boundary conditions, respectively. We observe that $L_{D}$ and $L_{1,D}\oplus L_{2,D}$ are both finite-dimensional self-ajoint extensions of $L_{1}\oplus L_{2}$.
By Proposition D.9, we have
$|\mathrm{n_{-}}(L_{D})-\mathrm{n_{-}}(L_{1,D}\oplus L_{2,D})|<\infty.$
Moreover $\mathrm{n_{-}}(L_{1,D}\oplus L_{2,D})=\mathrm{n_{-}}(L_{1,D})+\mathrm{n_{-}}(L_{2,D})$ where $\mathrm{n_{-}}(L_{1,D})<\infty$. So, we only need to study $L_{2,D}$.
Case 1.
We assume that $\liminf_{t\to+\infty}R(t)>-1/4I$. Let $M>0$ such that R(t)>1/4R(t)>-1/4 for $t\geq M$. Then, for each $u\in C_{0}^{\infty}((M,+\infty),{\mathbb{R}}^{n})$, by Hardy’s inequality, we have
$\langle l_{2}u,u\rangle>0$
and so $\mathrm{n_{-}}(L_{2,D})=0$. Then we get $\mathrm{n_{-}}(L_{2})<\infty$ .
Case 2.
We assume that $\limsup_{t\to+\infty}R(t)<-1/4I$. Let $M>0$ such that R(t)<1/4IR(t)<-1/4I for $t\geq M$. We now prove that for q<1/4q<-1/4, the operator $\dfrac{d^{2}}{dt^{2}}+\dfrac{q}{t^{2}}$ on $W^{2,2}((1,+\infty),{\mathbb{R}}^{n})$ with Dirichlet boundary condition has infinite Morse index.
By Theorem 5.2 and Remark 5.3, we only need to show that a solution of the equation with Dirichlet boundary condition has infinitely many zeroes. By Equation (8.2), for q<1/4q<-1/4, the SL-equation has the following solution
$\dfrac{1}{r}t^{1/2}\sin(r\ln t)\qquad\textrm{ for }\qquad q=-1/4-r^{2}$
having infinitely many zeros. This concludes the proof. ∎
Theorem 8.9.
Let $l_{R}$ be the operator defined above on $\mathscr{C}_{0}^{\infty}((0,1],{\mathbb{R}}^{n})$.
  • Case 1. If $\liminf_{t\to 0^{+}}R(t)>-1/4$, then any self-adjoint extension has finite Morse index.
  • Case 2. If $\limsup_{t\to 0^{+}}R(t)<-1/4$, then any self-adjoint extension has infinite Morse index.
Proof.
The proof of the Case 1 is similar to the corresponding of Theorem 8.8 and essentially based on splitting the interval $(0,1)$ into $(0,c]\cup[c,1]$ and then on using Hardy’s inequality on $\mathscr{C}_{0}^{\infty}((0,c),{\mathbb{R}}^{n})$.
The proof of Case 2 is also analogous to the corresponding of Theorem 8.8 since $\dfrac{1}{r}t^{1/2}\sin(r\ln t)$ also has infinite zeros on $(0,c)$ for any $c>0$. This concludes the proof. ∎
9 Asymptotic solutions of the N-body problem
The goal of this section is to prove at once and without any involved blow-up techniques the main results recently proved by authors in [BHPT20, HOY21, OP25] with some suitable ad-hoc methods. When a solution of the $n$-body problem does not experience any collision or noncollision singularity in the future or the past, then a natural and important question is about the final motion of the masses as times goes positive or negative infinity. A classification of possible final motions were listed by Chazy. In this section we will focus on total collision singularity as wellas the two of the simplest unbounded motions provided by the total parabolic/hyperbolic motion, which for simplicity will be referred as parabolic/hyperbolic motion.
9.1 A description of the problem
The Newtonian $n$-body problem studies the motion of $n$ point masses, $m_{i}>0$, according to Newton’s law of universal gravitation. Let $M=\operatorname{diag}\left(m_{1}I_{3},\ldots,m_{n}I_{3}\right)$ be the mass matrix, where $I_{3}$ is the $3\times 3$ identity matrix with $d\geq 1$. Then $q=\left(q_{i}\right)_{i=1}^{n}\left(q_{i}\in\mathbb{R}^{d}\right$. represents the position of $\left.m_{i}\right)$ satisfies the Newton’s equation
$M\ddot{q}=\nabla U(q)\quad\textrm{ where }\quad U(q)=\sum_{1\leq i<j\leq n}\frac{m_{i}m_{j}}{\left|q_{i}-q_{j}\right|}$ (9.1)
is the potential function (the negative potential energy) and $\nabla$ is the gradient with respect to the Euclidean metric.
The solutions of Equation (9.1) are invariant under linear translations, so there is no loss of generality to restrict ourselves to the $n^{*}:=d(n-1)$ dimensional subspace
$\mathcal{X}:=\left\{q\in\mathbb{R}^{3n}:\sum_{i=1}^{n}m_{i}q_{i}=0\right\}$
where the center of mass is fixed at the origin. Let $T\mathcal{X}$ be the tangent bundle of $\mathcal{X}$. The Lagrangian $L:T\mathcal{X}\rightarrow[0,+\infty)\cup$ $\{+\infty\}$
$L(q,v)=K(v)+U(q),\text{ where }K(v):=\frac{1}{2}|v|_{M}^{2}:=\frac{1}{2}\langle Mv,v\rangle,$
has singularities at the collision configurations
$\Delta=\bigcup_{1\leq i<j\leq n}\Delta_{ij},\quad\text{ where }\Delta_{ij}=\left\{q\in\mathcal{X}:q_{i}=q_{j}\right\}$
It is well-known that the Lagrangian action functional
$\mathcal{A}\left(q\right):=\int_{a}^{b}L(q(t),\dot{q}(t))\mathrm{d}t$
[…]
Because the Newtonian gravity is a weak force, the action value of a path with collisions could still be finite. This means the critical points obtained using variational methods may contain a subset of collision moments with zero measure and only satisfies Equation (9.1) in the complement of it. Such solutions were named generalized solutions by Bahri and Rabinowitz.
When a solution of the $n$-body problem does not experience any collision or non-collision singularity in the future or the past, then a natural and important question is about the final motion of the masses as times goes positive or negative infinity. A classification of possible final motions was listed by Chazy. In this section we will focus on total colliding motions and on two of the simplest classes of unbounded motions, namely the completely parabolic/hyperbolic motions.
Definition 9.1.
[…]
  • $q$ is a (completely) parabolic motion, if $\lim_{t\to\pm\infty}|q_{i}(t)-q_{j}(t)|=\infty$, for all $i\neq j$, and $\lim_{t\to\pm\infty}\dot{q}_{i}(t)=0$, for all $i$
  • $q$ is a hyperbolic motion if $\lim_{t\to\pm\infty}|q_{i}(t)-q_{j}(t)|=\infty$, for all $i\neq j$, and $\lim_{t\to\pm\infty}\dot{q}_{i}(t)$ exists for all $i$, and the limits are different from each other.
Moreover $q(\pm\infty)$ will be called a parabolic or hyperbolic infinity accordingly.
Under polar coordinates, when $q(\pm\infty)$ is a hyperbolic infinity, by Chazy we get that $s(t)$ converges to some $s^{\pm}\in\mathcal{E}$ as $t\to\pm\infty$. On the other hand if $q(\pm\infty)$ is a parabolic infinity, just like the case of a total collision, $s(t)$ converges to the set of normalized central configurations, as $t\to\pm\infty$. We are now in position to introduce the following.
Definition 9.2.
[…]
  • $q(T)$ is a total collision solution iff $T$ is finite
  • $q(T)$ is a parabolic/hyperbolic infinity if and only if $T=\infty$.
Remark 9.3.
[…]
By linearizing the Equation (9.1) along the solution $q$ we end-up with the following variational equation
$M\,\ddot{x}(t)=B(t)\,x(t)\qquad t\in[0,T)$ (9.2)
where $B(t)=D^{2}U(q(t))$ and where the Hessian matrix $D^{2}U(q)$ is the $3n\times 3n$ block symmetric matrix with $3\times 3$ blocks of the form
[…]
By introducing the polar coordinates
$r=\sqrt{I(q)}$ $s=\left(s_{i}\right)_{i=1}^{n}=q/r=\left(q_{i}/r\right)_{i=1}^{n}$ (9.4)
where $I(q)=\langle Mq,q\rangle$ is the moment of inertia and $\mathcal{E}:=\{q\in\mathcal{X}:I(q)=1\}$ is the set of normalized configurations, one finds that as $t\rightarrow T$, the radial part $r(t)$ satisfies the Sundman-Sperling estimate and $s(t)=q(t)/r(t)$ converges to the set of normalized central configurations, as $t\rightarrow T$ (it is not clear whether there will be a definite limit).
We recall that, a central configuration $q_{0}$ is a solution of the following equation
$\mu Mq_{0}+\nabla U(q_{0})=0\quad\textrm{ where }\quad\mu=\dfrac{U(q_{0})}{I(q_{0})}.$
A normalized central configuration $s\in\mathcal{E}$ is where the gradient of $U$ restricted on $\mathcal{E}$, vanishes. So,
$\left.\nabla U\right|_{\mathcal{E}}(s)=\nabla U(s)+U(s)Ms=0.$
At a normalized central configuration the Hessian can be represente with respect to the Euclidean product by the matrix $D^{2}U(q)+U(q)\cdot M$ or with respect to the mass inner product by
$M^{-1}D^{2}U(q)+U(q).$
It is worth noticing that, since
$\dfrac{D^{2}U(q)\cdot I(q)}{U(q)}=\dfrac{D^{2}U(\mu q)\cdot I(\mu q)}{U(\mu q)}$ (9.5)
then the function doesn’t depend on $\mu$.
Remark 9.4.
We observe that $D_{ij}=D_{ji}$. Moreover, it’s easy to check that the following three vectors
$\widetilde{e}_{1}=(e_{1},\ldots,e_{1})\qquad\widetilde{e}_{2}=(e_{2},\ldots,e_{2})\qquad\widetilde{e}_{3}=(e_{3},\ldots,e_{1})$ (9.6)
belong to $\ker D^{2}U(q)$.
The following result is crucial for establishing the spectral properties of the linearized operator along a total collisions and the parabolic/hyperbolic.
Lemma 9.5.
We assume that
  • (a)
    $q(T)$ is a total collision or a solution parabolic at infinity having limit normalized central configuration $a$. Then the following asymptotic estimates, hold
    $\lim_{t\to T^{\pm}}r(t)\,[\beta(t)]^{-2/3}=\left[\dfrac{9}{2}U(a)\right]^{1/3}\\$ (9.7)
    where
[…]
  • (b)
    $q(T)$ is hyperbolic at infinity, then
    $\lim_{t\to T}r(t)\,|t|^{-1}=\sqrt{2H_{0}}$ $\lim_{t\to T}|\dot{r}(t)|=\sqrt{2H_{0}}$ $\lim_{t\to T}r(t)|\dot{s}(t)|_{M}=0$ (9.8)
Proof.
For the proof we refer the interested reader to [HOY21, Lemma 2.1] and references therein. ∎
Remark 9.6.
It is worth observing that the estimates at item (a) are the well-known Sundman-Sperling estimates. [^4] We observe that there is a mistake in the asymptotic estimates stated in [BHPT20, Lemma 2.3] and [HOY21, Lemma 2.1]. The limit was written in terms of the constant $K=\dfrac{\alpha+2}{\alpha}\sqrt{2b}$, but the correct constant is $K=\dfrac{\alpha+2}{2}\sqrt{2b}$. This mistake does not affect either the results or the proofs contained therein.
By Equation (9.5), we get that
$\dfrac{I(q(t)}{U(q(t))}\cdot D^{2}U(q(t))=\dfrac{D^{2}U(q(t)/r(t))}{U(q(t)/r(t))}$
and so we immediately get that
$D^{2}U(q(t))=\dfrac{1}{r^{3}(t)}\cdot U(q(t)/r(t))\cdot\dfrac{D^{2}U(q(t)/r(t))}{U(q(t)/r(t))}.$
Under the assumption that the asymptotic solution has a limit normalized central configuration $a$, then $q(t)/r(t)$ converges to $a$ for $t\to T^{\pm}$. Moreover, by Lemma 9.5, we get that
$r^{3}(t)\sim_{T^{\pm}}\beta^{2}(t)\cdot\dfrac{9}{2}\,U(a).$
We define the asymptotic variational equation corresponding to the solution $t\mapsto q(t)$ as
$\ddot{q}(t)=\dfrac{1}{\beta^{2}(t)}\,\dfrac{2}{9}\,\dfrac{M^{-1}\,D^{2}U(a)}{U(a)}.$ (9.9)
If
$\dfrac{2}{9}\,\dfrac{M^{-1}\,D^{2}U(a)}{U(a)}>-\dfrac{1}{4}$ (9.10)
then we get that
$M^{-1}D^{2}U(a)+U(a)>-\dfrac{1}{8}\,U(a)$
which is precisely the so-called [BS]-condition at [BHPT20].
Remark 9.7.
As discussed at the Remark 8.1, the constant 1/4-1/4 which marks the transition in the Weyl classification of the limit-point, limit-circle cases in the case of Bessel-type differential operators is precisely the [BS]-condition for these asymptotic solutions.
In conclusion, the spectral properties and in particular the finiteness of the Morse index along an asymptotic solution in the N-body problem is described precisely by the limiting Bessel-type differential operators
[…]
where $\bar{B}(a):=\dfrac{2}{9}\,\dfrac{M^{-1}\,D^{2}U(a)}{U(a)}$.
Summing up all previous arguments we get a new proof of the main result proved by authors at [BHPT20, HOY21].
Theorem 9.8.
Let $q$ be a total collision or a parabolic solution at infinity having limiting normalized central configuration $a$, and set
$\bar{B}(a):=\frac{2}{9}\,\frac{M^{-1}D^{2}U(a)}{U(a)}.$
If
$\bar{B}(a)>-\dfrac{1}{4},$
then the Morse index $\mathrm{n_{-}}(q)$ of $q$ is finite, and the following identity holds:
$\mathrm{n_{-}}(q)=\sum_{t\in(a,b)}\dim\big(\gamma(t)\Lambda_{D}\cap\Lambda_{D}\big),$
where $\gamma$ is the fundamental matrix solution satisfying $\gamma(a)=I$.
If
$\bar{B}(a)<-\dfrac{1}{4},$
the Morse index is $\infty$.
If $q$ is hyperbolic, then the Morse index $\mathrm{n_{-}}(q)$ of $q$ is always finite.
Proof.
The proof follows by the above discussion and by invoking Theorem 8.9 and Theorem 5.2. The second statement is a direct consequence of [BS08]. The last statement readily follows from Theorem 8.8. (Cf. [HOY21] for further details). ∎
Remark $9.$
It’s worth noticing that with the theory constructed in the manuscript we are able to handle at once any self-adjoint boundary condition generalizing the main results proved by authors at [BHPT20] and [HOY21].
Appendix A A Concise Review of the Maslov-Type Indices
The purpose of this section is to present the functional analytic and symplectic preliminaries underlying the spectral flow and the Maslov index. Our primary references are [CLM94, Dui76, RS93, LZ00], from which we adopt some notation and definitions.
We denote by $(V,\omega)$ be the standard symplectic space and let $\mathscr{P}([a,b];V)$ the space of continuous maps
$f:J\to\Set{\textrm{pairs of Lagrangian subspaces in }V}$
equipped with the compact-open topology and we recall the following definition.
Definition A.1.
The CLM-index is the unique integer valued function
$\mu^{\scriptscriptstyle{\mathrm{CLM}}}:\mathscr{P}((a,b);{\mathbb{R}}^{2n})\to{\mathbb{Z}}$
which satisfies the following properties:
  • (I)
    Affine Scale Invariance
  • (II)
    Deformation Invariance relative to the Endpoints
  • (III)
    Path Additivity
  • (IV)
    Symplectic Additivity
  • (V)
    Symplectic Invariance
  • (VI)
    Normalization.
We refer the reader to [CLM94, Theorem 1.1]. Following authors in [LZ00, Section 3] and references therein, let us now introduce the notion of crossing form that gives an efficient way for computing the intersection indices in the Lagrangian Grassmannian context.
Let $\ell$ be a $\mathscr{C}^{1}$-curve of Lagrangian subspaces such that $\ell(0)=L$ and $\dot{\ell}(0)=\widehat{L}$. Now, if $W$ is a fixed Lagrangian subspace transversal to $L$. For $v\in L$ and small enough $t$, let $w(t)\in W$ be such that $v+w(t)\in\ell(t)$. Then the form
$Q(L,\widehat{L})[v]=\dfrac{d}{dt}\Big|_{t=0}\omega\big(v,w(t)\big)$ (A.1)
is independent on the choice of $W$.
Definition A.2.
Let $t\mapsto\ell(t)=(\ell_{1}(t),\ell_{2}(t))$ be a map in $\mathscr{P}([a,b];V)$. For $t\in[a,b]$, the crossing form is a quadratic form defined by
$\Gamma(\ell_{1},\ell_{2},t)=Q(\ell_{1}(t),\dot{\ell}_{1}(t))-Q(\ell_{2}(t),\dot{\ell}_{2}(t))\Big|_{\ell_{1}(t)\cap\ell_{2}(t)}$ (A.2)
A crossing instant for the curve $t\mapsto\ell(t)$ is an instant $t\in[a,b]$ such that $\ell_{1}(t)\cap\ell_{2}(t)\neq\{0\}$ non-trivially. A crossing is termed regular if the $\Gamma(\ell_{1},\ell_{2},t)$ is nondegenerate.
We observe that if $t$ is a crossing instant, then $\Gamma(\ell_{1},\ell_{2},t)=-\Gamma(\ell_{2},\ell_{1},t).$ If $\ell$ is regular meaning that it has only regular crossings, then the $\mu^{\scriptscriptstyle{\mathrm{CLM}}}$-index can be computed through the crossing forms, as follows
$\mu^{\scriptscriptstyle{\mathrm{CLM}}}\big(\ell_{1}(t),\ell_{2}(t),\Omega,t\in[a,b]\big)=\mathrm{n_{+}}\big(\Gamma(\ell_{2},\ell_{1},a)\big)+\sum_{a<t<b}\operatorname{sgn}\big(\Gamma(\ell_{2},\ell_{1},t)\big)-\mathrm{n_{-}}\big(\Gamma(\ell_{2},\ell_{1},b)\big)$ (A.3)
where the summation runs over all crossings $t\in(a,b)$ and $\mathrm{n_{+}},\mathrm{n_{-}}$ are the dimensions of the positive and negative spectral spaces, respectively and $\operatorname{sgn}:=\mathrm{n_{+}}-\mathrm{n_{-}}$ is the signature. (We refer the interested reader to [LZ00] ).
Let $L_{0}$ be a distinguished Lagrangian and we assume that $\ell_{1}(t)\equiv L_{0}$ for every $t\in J$. In this case we get that the crossing form at the instant $t$ provided in Equation (A.2) actually reduce to
$\Gamma\big(\ell_{2}(t),L_{0},t\big)=Q|_{\ell_{2}(t)\cap L_{0}}$ (A.4)
and hence
$\mu^{\scriptscriptstyle{\mathrm{CLM}}}\big(L_{0},\ell_{2}(t),\Omega,t\in[a,b]\big)=\mathrm{n_{+}}\big(\Gamma(\ell_{2},L_{0},a)\big)+\sum_{a<t<b}\operatorname{sgn}\big(\Gamma(\ell_{2},L_{0},t)\big)-\mathrm{n_{-}}\big(\Gamma(\ell_{2},L_{0},b)\big)$ (A.5)
A.1 On the Triple and Hörmander Index
A crucial ingredient that, in some sense, measures the difference in the relative Maslov index with respect to two different Lagrangian subspaces is the Hörmander index. This index is also related to the difference of the triple index and to an interesting generalization recently developed by the last author and his collaborators in [ZWZ18].
We begin with the following definition of the Hörmander index.
Definition A.3.
([ZWZ18, Definition 3.9]) Let $\lambda,\mu\in\mathscr{C}^{0}\big((a,b),\operatorname{\mathrm{Lag}}(V,\omega)\big)$ such that
$\lambda(a)=\lambda_{1},\quad\lambda(b)=\lambda_{2}\quad\textrm{ and }\quad\mu(a)=\mu_{1},\quad\mu(b)=\mu_{2}.$
Then the Hörmander index is the integer given by
$s(\lambda_{1},\lambda_{2};\mu_{1},\mu_{2}):=\mu^{\scriptscriptstyle{\mathrm{CLM}}}\big(\mu_{2},\lambda(t);t\in J\big)-\mu^{\scriptscriptstyle{\mathrm{CLM}}}\big(\mu_{1},\lambda(t);t\in J\big)\\ =\mu^{\scriptscriptstyle{\mathrm{CLM}}}\big(\mu(t),\lambda_{2};t\in J\big)-\mu^{\scriptscriptstyle{\mathrm{CLM}}}\big(\mu(t),\lambda_{1};t\in J\big).$ (A.6)
Compare [ZWZ18, Equation (17), pag. 736] once observing that we observe that $\mu^{\scriptscriptstyle{\mathrm{CLM}}}(\lambda,\mu)$ corresponds to $\textrm{Mas}\{\mu,\lambda\}$ in the notation of [ZWZ18].
Properties of the Hörmander index.
We briefly recall some well-useful properties of the Hörmander index.
  • $s(\lambda_{1},\lambda_{2};\mu_{1},\mu_{2})=-s(\lambda_{1},\lambda_{2};\mu_{2},\mu_{1})$
  • $s(\lambda_{1},\lambda_{2};\mu_{1},\mu_{2})=-s(\mu_{1},\mu_{2};\lambda_{1},\lambda_{2})+\sum_{j,k\in\{1,2\}}(-1)^{j+k+1}\dim(\lambda_{j}\cap\mu_{k})$.
  • If $\lambda_{j}\cap\mu_{k}=\{0\}$ then $s(\lambda_{1},\lambda_{2};\mu_{1},\mu_{2})=-s(\mu_{1},\mu_{2};\lambda_{1},\lambda_{2})$.
The Hörmander index is computable as the difference of the two Maslov indices each one involving three different Lagrangian subspaces. This index is defined in terms of the local chart representation of the atlas of the Lagrangian Grassmannian manifold.
Definition A.4.
$\alpha,\beta,\gamma\in\operatorname{\mathrm{Lag}}(V,\omega)$ and we assume that $\alpha\cap\beta=\gamma\cap\beta=(0)$. Then we define the quadratic form $Q(\alpha,\beta;\gamma):\alpha\to{\mathbb{R}}$ as follows
[…]
Definition A.5.
Let $\alpha,\beta,\gamma\in\operatorname{\mathrm{Lag}}(V,\omega)$, $\varepsilon:=\alpha\cap\beta+\beta\cap\gamma$ and let $\pi:=\pi_{\varepsilon}$ be the projection in the symplectic reduction of $V$ mod $\varepsilon$. We term triple index the integer defined by
$\iota(\alpha,\beta,\gamma):=\mathrm{n}_{+}\,\big[Q(\pi\alpha,\pi\beta;\pi\gamma)\big]+\dim(\alpha\cap\gamma)-\dim(\alpha\cap\beta\cap\gamma)\\[3.0pt] \leq n-\dim(\alpha\cap\beta)-\dim(\beta\cap\gamma)+\dim(\alpha\cap\beta\cap\gamma).$ (A.7)
Remark A.6.
Definition A.5 is well-posed and we refer the interested reader to [Dui76, Lemma 2.4] and [ZWZ18, Corollary 3.12 & Lemma 3.13] for further details). It is worth noticing that $Q(\pi\alpha,\pi\beta;\pi\gamma)$ is a quadratic form on $\pi\alpha$. Being the reduced space $V_{\varepsilon}$ a $2(n-\dim\varepsilon)$ dimensional subspace, it follows that inertial indices of $Q(\pi\alpha,\pi\beta;\pi\gamma)$ are integers between $\{0,\ldots,n-\dim\varepsilon\}$.
Remark A.7.
We also observe that for arbitrary Lagrangian subspaces $\alpha,\beta,\gamma$, the quadratic form $Q(\alpha,\beta,\gamma)$ is well-defined and it is a quadratic form on $\alpha\cap(\beta+\gamma)$. Furthermore, we have
$\mathrm{n}_{+}\,\big[Q(\alpha,\beta,\gamma)\big]=\mathrm{n}_{+}\,\big[Q(\pi\alpha,\pi\beta,\pi\gamma)\big].$
So, we can also define the triple index as
$\iota(\alpha,\beta,\gamma):=\mathrm{n}_{+}\,\big[Q(\alpha,\beta;\gamma)\big]+\dim(\alpha\cap\gamma)-\dim(\alpha\cap\beta\cap\gamma).$
Authors in [ZWZ18, Lemma 3.2] give a useful property for calculating such a quadratic form.
$\mathrm{n}_{+}\,\big[Q(\alpha,\beta,\gamma)\big]=\mathrm{n}_{+}\,\big[Q(\beta,\gamma,\alpha)\big]=\mathrm{n}_{+}\,\big[Q(\gamma,\alpha,\beta)].$ (A.8)
We observe that if $(\alpha,\beta)$ is a Lagrangian decomposition of $(V,\omega)$ and $\beta\cap\gamma=\{0\}$ then $\pi$ reduces to the identity and both terms $\dim(\alpha\cap\gamma)$ and $\dim(\alpha\cap\beta\cap\gamma)$ drop down. In this way the triple index is nothing different from the the quadratic form $Q$ defining the local chart of the atlas of $\operatorname{\mathrm{Lag}}(V,\omega)$. It is possible to prove (cf. [ZWZ18, proof of the Lemma 3.13]) that
$\dim(\alpha\cap\gamma)-\dim(\alpha\cap\beta\cap\gamma)=\mathrm{n}_{0}\,\big[Q(\pi\alpha,\pi\beta;\pi\gamma)\big],$ (A.9)
where we denoted by $\mathrm{n}_{0}\,Q$ the nullity (namely the kernel dimension of the quadratic form $Q$). By summing up Equation (A.7) and Equation (A.9), we finally get
[…]
Lemma A.8.
Let $\lambda\in\mathscr{C}^{1}\big((a,b),\operatorname{\mathrm{Lag}}(V,\omega)\big)$. Then, for every $\mu\in\operatorname{\mathrm{Lag}}(V,\omega)$, we have
  1. (I)
    $s\big(\lambda(a),\lambda(b);\lambda(a),\mu\big)=-\iota\big(\lambda(b),\lambda(a),\mu\big)\leq 0$,
  2. (II)
    $s\big(\lambda(a),\lambda(b);\lambda(b),\mu\big)=\iota\big(\lambda(a),\lambda(b),\mu\big)\geq 0$.
Proof.
For the proof, we refer the interested reader to [ZWZ18, Corollary 3.16]. ∎
The next result, which is the main result of [ZWZ18], allows us to reduce the computation of the Hörmander index to the computation of the triple index.
Proposition A.9.
[ZWZ18, Theorem 1.1] Let $(V,\omega)$ be a $2n$-dimensional symplectic space and let $\lambda_{1},\lambda_{2},\mu_{1},\mu_{2}\in\operatorname{\mathrm{Lag}}(V,\omega)$. Under the above notation, we get
$s(\lambda_{1},\lambda_{2},\mu_{1},\mu_{2})=\iota(\lambda_{1},\lambda_{2},\mu_{2})-\iota(\lambda_{1},\lambda_{2},\mu_{1})\\ =\iota(\lambda_{1},\mu_{1},\mu_{2})-\iota(\lambda_{2},\mu_{1},\mu_{2})$ (A.11)
Remark A.10.
We emphasize that no transversality conditions are assumed on the Lagrangian subspaces in Proposition A.9
Lemma A.11.
Let $(V,\omega)$ be a finite dimensional symplectic space and let $W$ be a coisotropic subspace. Then $W$ is a Lagrangian subspace if and only if it contains a unique Lagrangian subspace of $V$.
Proof.
Assume that $\Lambda\subsetneq W$ is a Lagrangian subspace of $V$. Let $v\in W\setminus\Lambda$. Let $l$ be the one dimension subspace spanned by $v$. Then we have
$((l^{\omega}\cap\Lambda)+l)^{\omega}=(l^{\omega\omega}+\Lambda^{\omega})\cap l^{\omega}=(l+\Lambda)\cap l^{\omega}.$
Since one dimension subspace of $V$ is isotropic, we have $l<l^{\omega}$. It follows that
$((l^{\omega}\cap\Lambda)+l)^{\omega}=(l+\Lambda)\cap l^{\omega}=l+\Lambda\cap l^{\omega}.$
So, $W$ contains a subspace $(l^{\omega}\cap\Lambda)+l$ which is different with $\Lambda$ and it is a Lagrangian subspace of $V$. So if $\Lambda<W$ is the unique Lagrangian subspace contained in $W$, we have $W=\Lambda$ and $W$ is a Lagrangian subspace of $V$.
Lemma A.12.
Let $\alpha,\beta,\gamma\in\operatorname{\mathrm{Lag}}(V,\omega)$. Then the following identities hold:
$\iota(\alpha,\beta,\gamma)-\iota(\beta,\gamma,\alpha)=\dim(\alpha\cap\gamma)-\dim(\beta\cap\alpha)$
or equivalently
$\iota(\alpha,\beta,\gamma)-\dim(\alpha\cap\gamma)=\iota(\beta,\gamma,\alpha)-\dim(\beta\cap\alpha)$
Proof.
By the definition of triple index
$\iota(\alpha,\beta,\gamma)=\mathrm{n_{+}}\big[Q(\alpha,\beta;\gamma)\big]+\dim(\alpha\cap\gamma)-\dim(\alpha\cap\beta\cap\gamma)$
$\iota(\beta,\gamma,\alpha)=\mathrm{n_{+}}\big[Q(\beta,\gamma;\alpha)\big]+\dim(\alpha\cap\beta)-\dim(\alpha\cap\beta\cap\gamma).$
Since $\mathrm{n_{+}}[Q]$ is invariant under circular permutation, the lemma follows. ∎
Appendix B A Quick Overview of the Spectral Flow
The aim of this section is to briefly recall the Definition and the main properties of the spectral flow for a continuous path of closed self-ad[a,b]oint Fredholm operators $\operatorname{\mathcal{CF}^{sa}}(H)$ on the Hilbert space $H$. Our basic reference is [HP17] and references therein.
Let $H$ be a separable complex Hilbert space and let $A:D(A)<H\to H$ be a self-ad[a,b]oint Fredholm operator. By the Spectral decomposition Theorem (cf., for instance, [Kat80, Chapter III, Theorem 6.17]), there is an orthogonal decomposition $H=E_{-}(A)\oplus E_{0}(A)\oplus E_{+}(A),$ that reduces the operator $A$ and has the property that
$\sigma(A)\cap(-\infty,0)=\sigma\big(A_{E_{-}(A)}\big),\quad\sigma(A)\cap\{0\}=\sigma\big(A_{E_{0}(A)}\big),\quad\sigma(A)\cap(0,+\infty)=\sigma\big(A_{E_{+}(A)}\big).$
Definition B.1.
Let $A\in\operatorname{\mathcal{CF}^{sa}}(H)$. We term $A$ essentially positive if $\sigma_{ess}(A)\subset(0,+\infty)$, essentially negative if $\sigma_{ess}(A)\subset(-\infty,0)$ and finally strongly indefinite respectively if $\sigma_{ess}(A)\cap(-\infty,0)\not$$\emptyset$ and $\sigma_{ess}(A)\cap(0,+\infty)\not$$\emptyset$.
If $\dim E_{-}(A)<\infty$, we define its Morse index as the integer denoted by $\mu_{\scriptscriptstyle{\mathrm{Mor}}}\left[A\right]$ and defined as $\mu_{\scriptscriptstyle{\mathrm{Mor}}}\left[A\right]:=\dim E_{-}(A).$ Given $A\in\mathcal{CF}^{sa}(H)$, for $a,b\notin\sigma(A)$ we set
$\mathcal{P}_{[a,b]}(A):=\mathrm{Re}\left(\dfrac{1}{2\pi\,i}\int_{\gamma}(\lambda-A)^{-1}d\,\lambda\right)$
where $\gamma$ is the circle of radius $\frac{b-a}{2}$ around the point $\frac{a+b}{2}$. We recall that if $K\subset\sigma(A)$ consists of isolated eigenvalues of finite type then $\mathrm{Im}\mathcal{P}_{K}(A)=E_{K}(A):=\bigoplus_{\lambda\in(a,b)}\ker(\lambda-A);$ (cf. [GGK90, Section XV.2], for instance) and 0 either belongs in the resolvent set of $A$ or it is an isolated eigenvalue of finite multiplicity. The next result allows us to define the spectral flow for gap continuous paths in $\mathcal{CF}^{sa}(H)$.
Proposition B.2.
Let $A_{0}\in\mathcal{CF}^{sa}(H)$ be fixed.
  1. (i)
    There exists a positive real number $a\notin\sigma(A_{0})$ and an open neighborhood $\mathscr{N}\subset\mathcal{CF}^{sa}(H)$ of $A_{0}$ in the gap topology such that $\pm a\notin\sigma(A)$ for all $A\in\mathscr{N}$ and the map
    $\mathscr{N}\ni A\longmapsto\mathcal{P}_{[-a,a]}(A)\in\mathcal{B}(H)$
    is continuous and the pro[a,b]ection $\mathcal{P}_{[-a,a]}(A)$ has constant finite rank for all $t\in\mathscr{N}$.
  2. (ii)
    If $\mathscr{N}$ is a neighborhood as in (i) and $a\leq c\leq d\leq a$ are such that $c,d\notin\sigma(A)$ for all $A\in\mathscr{N}$, then $A\mapsto\mathcal{P}_{[c,d]}(A)$ is continuous on $\mathscr{N}$. Moreover the rank of $\mathcal{P}_{[c,d]}(A)\in\mathscr{N}$ is finite and constant.
Proof.
For the proof of this result we refer the interested reader to [BLP05, Proposition 2.10]. ∎
Let $A:[a,b]\to\mathcal{CF}^{sa}(H)$ be a gap-continuous path. As consequence of Proposition B.2, for every $t\in[a,b]$ there exists $a>0$ and an open connected neighborhood $\mathscr{N}_{t,a}\subset\mathcal{CF}^{sa}(H)$ of $\mathcal{A}(t)$ such that $\pm a\notin\sigma(A)$ for all $A\in\mathscr{N}_{t,a}$ and the map $\mathscr{N}_{t,a}\in A\longmapsto\mathcal{P}_{[-a,a]}(A)\in\mathcal{B}$ is continuous and hence $\operatorname{rank}\left(\mathcal{P}_{[-a,a]}(A)\right)$ does not depends on $A\in\mathscr{N}_{t,a}$. Let us consider the open covering of the interval $[a,b]$ given by the pre-images of the neighborhoods $\mathcal{N}_{t,a}$ through $A$ and, by choosing a sufficiently fine partition of the interval $[a,b]$ having diameter less than the Lebesgue number of the covering, we can find $a=:t_{0}<t_{1}<\dots<t_{n}:=b$, operators $T_{i}\in\mathcal{CF}^{sa}(H)$ and positive real numbers $a_{i}$, $i=1,\dots,n$ in such a way the restriction of the path $A$ on the interval $t_{i-1},t_{i}]$ lies in the neighborhood $\mathscr{N}_{t_{i},a_{i}}$ and hence the $\dim E_{[-a_{i},a_{i}]}(A_{t})$ is constant for $t\in[t_{i-1},t_{i}]$, $i=1,\dots,n$.
Definition B.3.
The spectral flow of $A$ (on the interval $[a,b]$) is defined by
$\operatorname{sf}(A,[a,b]):=\sum_{i=1}^{N}\dim\,E_{[0,a_{i}]}(A_{t_{i}})-\dim\,E_{[0,a_{i}]}(A_{t_{i-1}})\in{\mathbb{Z}}.$
The spectral flow as given in Definition B.3 is well-defined (in the sense that it is independent either on the partition or on the $a_{i}$) and only depends on the continuous path $A$. Here We list one of the useful properties of the spectral flow.
  • (Path Additivity) If $A_{1},\mathcal{A}_{2}:[a,b]\to\mathcal{CF}^{sa}(H)$ are two continuous path such that $A_{1}(b)=A_{2}(a)$, then $\operatorname{sf}(A_{1}*A_{2})=\operatorname{sf}(A_{1})+\operatorname{sf}(A_{2}).$
As already observed, the spectral flow, in general, depends on the whole path and not [a,b]ust on the ends. However, if the path has a special form, it actually depends on the end-points. More precisely, let $A,B\in\mathcal{CF}^{sa}(H)$ and let $\widetilde{A}:[a,b]\to\mathcal{CF}^{sa}(H)$ be the path pointwise defined by $\widetilde{A}(t):=A+\widetilde{\mathcal{B}}(t)$ where $\widetilde{B}$ is any continuous curve of $A$-compact operators parametrised on $[a,b]$ such that $\widetilde{B}(a):=0$ and $\widetilde{B}(b):=B$. In this case, the spectral flow depends of the path $\widetilde{A}$, only on the endpoints (cf. [ZL99] and reference therein).
Remark B.4.
It is worth noticing that, since every operator $\widetilde{A}(t)$ is a compact perturbation of a a fixed one, the path $\widetilde{A}$ is actually a continuous path into $\mathcal{B}(W;H)$, where $W:=D(A)$.
Definition B.5.
([ZL99, Definition 2.8]). Let $A,B\in\mathcal{CF}^{sa}(H)$ and we assume that $B$ is $A$-compact (in the sense specified above). Then the relative Morse index of the pair $A$, $A+B$ is defined by $I(A,A+B)=-\operatorname{sf}(\widetilde{A};[a,b])$ where $\widetilde{A}:=A+\widetilde{B}(t)$ and where $\widetilde{B}$ is any continuous curve parametrised on $[a,b]$ of $A$-compact operators such that $\widetilde{B}(a):=0$ and $\widetilde{B}(b):=B$.
In the special case in which the Morse index of both operators $A$ and $A+B$ are finite, then
$I(A,A+B)=\mathrm{n_{-}}(A+B)-\mathrm{n_{-}}(A).$ (B.1)
Let $W,H$ be separable Hilbert spaces with a dense and continuous inclusion $W\hookrightarrow H$ and let $A:[a,b]\to\mathcal{CF}^{sa}(H)$ having fixed domain $W$. We assume that $A$ is a continuously differentiable path $A:[a,b]\to\mathcal{CF}^{sa}(H)$ and we denote by $\dot{A}_{\lambda_{0}}$ the derivative of $A_{\lambda}$ with respect to the parameter $\lambda\in[a,b]$ at $\lambda_{0}$.
Definition B.6.
An instant $\lambda_{0}\in[a,b]$ is called a crossing instant if $\ker\,A_{\lambda_{0}}\neq 0$. The crossing form at $\lambda_{0}$ is the quadratic form defined by
$\Gamma(A,\lambda_{0}):\ker A_{\lambda_{0}}\to{\mathbb{R}},\quad\Gamma(A,\lambda_{0})[u]=\langle\dot{A}_{\lambda_{0}}\,u,u\rangle_{H}.$ (B.2)
Moreover a crossing $\lambda_{0}$ is called regular, if $\Gamma(\mathcal{A},\lambda_{0})$ is nondegenerate.
We recall that there exists $\varepsilon>0$ such that $A+\delta\,I_{H}$ has only regular crossings for almost every $\delta\in(-\varepsilon,\varepsilon)$. In the special case in which all crossings are regular, then the spectral flow can be easily computed through the crossing forms. More precisely the following result holds.
Proposition B.7.
If $A:[a,b]\to\mathcal{CF}^{sa}(W,H)$ has only regular crossings then they are in a finite number and
$\operatorname{sf}(A,[a,b])=-\mathrm{n_{-}}{\left[\Gamma(A,a)\right]}+\sum_{t_{0}\in(a,b)}\operatorname{sgn}\left[\Gamma(A,t_{0})\right]+\mathrm{n_{+}}\left[\Gamma(A,b)\right]$
where the sum runs over all the crossing instants.
Proof.
The proof of this result follows by arguing as in [RS95]. This conclude the proof. ∎
Appendix C Friedrichs Extensions: Definition and Basic Results
Friedrich extensions provide a canonical way to associate a self-adjoint operator to a densely defined, symmetric, semi-bounded from below operator in a Hilbert space. This construction is particularly relevant in the spectral theory of differential operators, where it often yields the physically or geometrically meaningful self-adjoint realization of a given symmetric operator.
In this section, we recall the definition of the Friedrich extension and collect some basic properties that are crucial for the analysis of self-adjoint realizations of differential operators. We emphasize that the Friedrich extension is characterized from a variational viewpoint as the smallest self-adjoint extension (in the sense of quadratic forms) among all self-adjoint extensions bounded below by the same constant. This property makes the Friedrich extension particularly useful in applications to partial differential equations, quantum mechanics, and index theory.
Let $(H,\langle\cdot,\cdot\rangle$ be a real Hilbert space and let $A:D(A)<H\to H$ be a densely defined, symmetric operator on a Hilbert space $H$, satisfying the semiboundedness condition:
$\langle Au,u\rangle\geq m\|u\|^{2},\quad\forall\,u\in D(A),$
for some constant $m\in\mathbb{R}$. Without loss of generality, one may assume $m>0$ by replacing $A$ with $A+cI$ for a suitable shift c>mc>-m.
We define the associated quadratic form $q$ by
$q(u,v):=\langle Au,v\rangle,\quad u,v\in D(q):=D(A).$
This form is symmetric and semibounded from below by $m$. Introduce the form inner product:
$\langle u,v\rangle_{q}:=q(u,v)+(1-m)\langle u,v\rangle,$
which induces the form norm:
$\|u\|_{q}:=\sqrt{q(u,u)+(1-m)\|u\|^{2}}.$
Let $H_{q}$ denote the completion of $D(A)$ with respect to $\|\cdot\|_{q}$. The form $q$ then extends uniquely to a closed, densely defined, symmetric, and semibounded quadratic form $t_{A}$ on $H_{q}$.
By the Kato’s first representation theorem, there exists a unique self-adjoint operator $A_{F}$ on $H$, called the Friedrichs extension of $A$, such that:
    $D(A_{F})<D(t_{A})$,
    For all $u\in D(A_{F})$, there exists $f\in H$ such that
    $t_{A}(u,v)=\langle f,v\rangle,\qquad\forall\,v\in D(q),$
    and $A_{F}u:=f$.
Hence, the domain of the Friedrichs extension is given by:
$D(A_{F})=\left\{u\in D(q):\exists\,f\in H\text{ such that }t_{A}(u,v)=\langle f,v\rangle,\ \forall\,v\in D(q)\right\}.$
Remark C.1.
The key feature of the Friedrichs extension is that its domain is determined entirely by the closed quadratic form. A function $u\in D(q)$ lies in $D(A_{F})$ if and only if the map $v\mapsto t_{A}(u,v)$ is continuous with respect to the Hilbert space norm on $H$. In other words, there exists a constant $C>0$ such that
$|t_{A}(u,v)|\leq C\|v\|,\quad\forall v\in D(q).$
By the Riesz representation theorem, this defines a unique element $f\in H$, and we set $A_{F}u:=f$.
This variational characterization ensures that the Friedrichs extension is the “smallest” self-adjoint extension in the sense of form ordering. It typically corresponds to imposing Dirichlet-type boundary conditions or energy minimization constraints.
More explicitly by using the Kato’s representation theorem for closed, semibounded quadratic forms, we get that there exists a unique self-adjoint and semibounded operator $A_{F}$ on $H$ such that:
    $D\left(A_{F}^{1/2}\right)=D(t_{A})=H_{q}$
    $\langle A_{F}^{1/2}\phi,\ A_{F}^{1/2}\psi\rangle=t_{A}(\phi,\psi)$ for all $\phi,\psi\in H_{q}$.
Lemma C.2.
We assume that $A:D(A)\to H$ is a Fredholm operator. Then $A_{F}:H_{q}\to H_{q}$ is also a Fredholm operator.
Proof.
Since $A$ is a Fredholm operator, then the Friedrich extension $A_{F}$ is also a Fredholm operator. The following diagram
[…]
is commutative. Since $(A_{F}+(1-m)I)^{1/2}:D\big([A_{F}+(1-m)I]^{1/2}\big)\to H$ is a homeomorphism, we conclude that $A_{F}:H_{q}\to:H_{q}$ is a Fredholm operator. ∎
The main conclusions of the preceding discussion are encapsulated in the following result.
Theorem C.3.
[Friedrich extension] Let $A:D(A)<H\to H$ be a densely defined, symmetric and non-negative operator on a Hilbert space $H$. There is a unique self-adjoint operator $A_{F}:D(A_{F})<H\to H$ such that $D(A_{F})<D(t_{A})$ and
$t_{A}(u,v)=\langle A_{F}u,v\rangle_{H}\qquad\forall\,u\in D(A_{F})\quad v\in D(t_{A}).$
Furthermore, $A_{F}$ is a self-adjoint extension of $A$.
Such a self-adjoint extension of $A$ is called the Friedrichs extension of $A$.
Corollary C.4.
Let $A:D(A)<H\to H$ be a non-negative self-adjoint operator with dense domain $D(A)$. Then the Friedrichs extension of $A$ coincides with $A$ itself.
Remark C.5.
The Friedrichs extension provides a canonical self-adjoint realization of a symmetric and semibounded operator. In particular, if a densely defined symmetric operator $A$ is essentially self-adjoint and semibounded from below, then its closure is self-adjoint, and the Friedrichs extension coincides with this closure. Thus, in the case of semibounded and essentially self-adjoint operators, the Friedrichs extension reproduces the unique self-adjoint extension.
The following result is straightforward and the proof is left to the reader.
Lemma C.6.
Under the above notation, we get
$\ker t_{A}=\ker A_{F}.$
Proof.
The proof directly follows by a straightforward computation. ∎
Lemma C.7.
Let $H$ be a Hilbert space and let $U<V<H$ be two dense subspaces of $H$. We assume that $A_{V}:V<H\to H$ is a semi-bounded closed and symmetric operator. Let $A_{U}=L|_{U}$. Then we have
$t_{A_{U}}=t_{A_{V}}|_{D(t_{A_{U}})}.$
Proof.
Without loss of generality, we may assume that $t_{A_{V}}(u)\geq\|u\|_{H}^{2}$.
Note that $\langle A_{U}\,u,u\rangle_{H}=t_{A_{V}}(u)$ for all $u\in E$. Therefore, $D(t_{A_{V}})$ is the closure of $U$ with respect to the norm $t_{A_{U}}(\cdot)^{1/2}$, and we have
$t_{A_{U}}=t_{A_{V}}|_{D(t_{A_{U}})},$
concluding the proof. ∎
Remark C.8.
By Lemma C.7 and by the very definition of the Friedrich extension, we get that
$A^{*}|_{D(A_{F})}=A_{F}.$
So, $A_{F}$ is characterized by its domain.
We are now ready to characterize the domain of the Friedrich extension $A_{F}$ of a closed, Fredholm symmetric and semi-bounded operator in terms of the adjoint domain $D(A^{*})$ and the form domain $D(t_{A})$. We state the result below.
Lemma C.9.
Let $A\in\operatorname{\mathcal{CF}^{s}}(H)$ and we assume that $A$ is semi-bounded from below. Then, we have
$D(A^{*})\cap D(t_{A})=D(A_{F}).$
Proof.
Consider the canonical projection $\gamma:D(A^{*})\to\beta(A):=D(A^{*})/D(A)$, and observe that $D(A_{F})<D(t_{A})$. Therefore, we have
$D(A_{F})<D(t_{A})\cap D(A^{*}).$
Let $\mu:=\gamma(D(t_{A})\cap D(A^{*}))$. Then $\mu\supset\gamma(D(A_{F}))$. Since $\gamma(D(A_{F}))$ is a Lagrangian subspace of $\beta(A)$, it follows that $\mu$ is a coisotropic subspace of $\beta(A)$. By Lemma A.11, it suffices to show that $\gamma(D(A_{F}))$ is the unique Lagrangian subspace contained in $\mu$.
Let $\Lambda$ be a Lagrangian subspace of $\beta(A)$ such that $\Lambda<\mu$, and set $A_{\Lambda}:=A^{*}|_{\gamma^{-1}(\Lambda)}$. Then $A_{\Lambda}$ is a self-adjoint extension of $A$. By Lemma C.7, we have
$t_{A}=t_{A_{\Lambda}}|_{D(t_{A})}.$
Since $D(A_{\Lambda})<D(t_{A})$, it follows that
$t_{A}|_{D(A_{\Lambda})}=t_{A_{\Lambda}}|_{D(A_{\Lambda})}.$
Thus,
$t_{A}(x,y)=\langle A_{\Lambda}x,y\rangle_{H},\quad\forall x,y\in D(A_{\Lambda}).$
Because $D(A)<D(A_{\Lambda})<D(t_{A})$ and $D(A)$ is dense in the Hilbert space $D(t_{A})$, it follows that $D(A_{\Lambda})$ is also dense in $D(t_{A})$.
Hence, for each $u\in D(t_{A})$, there exists a sequence $(u_{n})_{n}<D(A_{\Lambda})$ such that $u_{n}\to u$ both in $D(t_{A})$ and in $H$. Therefore,
$t_{A}(x,y)=\langle A_{\Lambda}x,y\rangle_{H},\quad\forall x\in D(A_{\Lambda})\text{ and }y\in D(t_{A}).$
This shows that $A_{\Lambda}$ is the unique Friedrich extension of $A$. Consequently, $\gamma(D(A_{\Lambda}))$ is the unique Lagrangian subspace contained in $\mu$.
By Lemma A.11, we finally conclude that
$D(A_{F})=D(t_{A})\cap D(A^{*}),$
which completes the proof. ∎
Example C.10.
We consider the Sturm–Liouville operator $l=-\dfrac{d^{2}}{dt^{2}}$ defined on $\mathscr{\mathscr{}}C_{0}^{\infty}([0,1],\mathbb{R}^{n})$. Then, we have
$D(A_{F})=D(A^{*})\cap D(t_{A})=W^{2,2}([0,1],\mathbb{R}^{n})\cap W^{1,2}_{0}([0,1],\mathbb{R}^{n}).$
In this case, $D(A_{F})$ provides the appropriate functional setting for treating the boundary value problem with Dirichlet boundary conditions.
We refer to the general boundary conditions parametrized by the quotient space $D(A_{F})/D(A)$ as the general Dirichlet boundary conditions.
Appendix D Auxiliary Results
This section contains a number of auxiliary results used for proving the main result of the manuscript. We collect all the results in this appendix, for smoothing the reading of the manuscript.
D.1 Miscellaneous Results
Lemma D.1.
If $s\mapsto E_{s}$ and $s\mapsto F_{s}$ are two gap-continuous paths in $\operatorname{\mathcal{S}}(H)$ such that $s\mapsto E_{s}+F_{s}$ is also gap-continuous in $\operatorname{\mathcal{S}}(H)$. Then $s\mapsto E_{s}\cap F_{s}$ is a gap-continuous path in $\operatorname{\mathcal{S}}(H)$.
If $s\mapsto E_{s}\cap F_{s}$, $s\mapsto E_{s}$, $s\mapsto F_{s}$ are gap continuous and $E_{s}+F_{s}$ is closed for every $s\in[0,1]$, then the path $s\mapsto E_{s}+F_{s}$ is gap continuous in $\operatorname{\mathcal{S}}(H)$.
Proof.
For the proof of this result we refer to [BZ18, Corollary A.3.14] and references therein. ∎
Corollary D.2.
Let $E,F$ be Hilbert spaces. Let $A_{s}:E\rightarrow F,s\in[0,1]$ be a gap-continuous path of operators. Assume that im $A_{s}=F,s\in[0,1]$. Then $\operatorname{ker}A_{s}$ is a gap-continuous path in $\mathcal{S}(E)$.
Proof.
Without loss of generality, we can assume $E,F$ be subspace of $E\oplus F$. We have $\mathcal{G}\left(A_{s}\right)+E=$ $E+\operatorname{im}\left(A_{s}\right)=E\oplus F$. By Lemma D.1, $\ker A_{s}=\mathcal{G}\left(A_{s}\right)\cap E$ is a gap-continuous path in $\mathcal{S}(E)$. ∎
Lemma D.3.
Let $H$ be a Hilbert space and for $i=1,2$ we consider the symmetric and Fredholm operators $T_{i}:D(T_{i})<H\to H$. We assume that
    $D(T_{1})<D(T_{2})$
    $T_{1}=T_{2}|_{D(T_{1})}$.
Then we have
$\dim(D(T_{2})/D(T_{1}))=\mathrm{ind\,}(T_{2})-\mathrm{ind\,}(T_{1}).$
Proof.
Let $j$ be the injection $D(T_{1})\to D(T_{2})$. Then we have
$\mathrm{ind\,}(T_{1})=\mathrm{ind\,}(T_{2}\,j)=\mathrm{ind\,}(T_{2})+\mathrm{ind\,}(j)=\mathrm{ind\,}(T_{2})-\operatorname{coker}j=\mathrm{ind\,}(T_{2})-\dim(D(T_{2})/D(T_{1})).$
Lemma D.4.
Let $T\in\operatorname{\mathcal{CF}^{s}}(H)$ be densely defined, $T_{1}$ be a self-adjoint extension of $T$ and we assume that $\dim\big(D(T_{1})/D(T)\big)<+\infty$. Then, we have
$\sigma_{ess}(T_{1})=\sigma_{ess}(T)$
where $\sigma_{ess}(\#)$ denotes the essential spectrum of the operator #\#.
Proof.
We prove that $\sigma_{ess}(T_{1})\supset\sigma_{ess}(T)$ being the proof of the opposite inclusion analogous. Let $n:=\dim\big(D(T_{1})/D(T)\big)$. By Lemma 2.2, we get that
$\dim\beta(T):=\dim(D(T^{*})/D(T))=2\,n$
and
$\dim(D(T^{*})/D(T_{1}))=n.$
Assume that $\bar{\lambda}\in{\mathbb{C}}\setminus\sigma_{ess}(T_{1})$. By definition, we get $T_{1}-\bar{\lambda}I\in\operatorname{\mathcal{CF}^{sa}}(H)$. Since $\operatorname{im}(T^{*}-\bar{\lambda}I)/\operatorname{im}(T_{1}-\bar{\lambda}I)$ is finite dimensional then $\operatorname{im}(T^{*}-\bar{\lambda}I)$ is closed and it has finite codimension and son $T^{*}-\bar{\lambda}I\in\operatorname{\mathcal{CF}^{sa}}(H)$. Now, since $T-\lambda I=(T^{*}-\bar{\lambda}I)^{*}\in\operatorname{\mathcal{CF}^{sa}}(H)$, we get that
$\sigma_{ess}(T_{1})\supset\sigma_{ess}(T)$
concluding the proof. ∎
Lemma D.5.
Let $A\in\operatorname{\mathcal{CF}^{s}}(H)$ and let $\Lambda\in\operatorname{\mathrm{Lag}}(H)$. Under the notation above, the following identity holds:
$(D(t_{A}))^{t_{A_{\Lambda}}}=(D(A_{F})+D(A_{\Lambda}))\cap\ker A^{*}.$
Proof.
Since $D(A)$ is dense in the Hilbert space $D(t_{A})$ and $t_{A_{\Lambda}}$ is a bounded form on $D(t_{A_{\Lambda}})$, we have
$(D(t_{A}))^{t_{A_{\Lambda}}}=(D(A))^{t_{A_{\Lambda}}}=\{y\in D(t_{A_{\Lambda}})\mid t_{A_{\Lambda}}[x,y]=0,\ \forall x\in D(A)\}.$
Recalling the definition of the Friedrich extension, for every $x\in D(A)$ we have
$t_{A_{\Lambda}}[x,y]=\langle A_{\Lambda}x,y\rangle_{H}=\langle Ax,y\rangle_{H}.$
It follows that
$(D(t_{A}))^{t_{A_{\Lambda}}}=(\operatorname{im}A)^{\perp}\cap D(t_{A_{\Lambda}})=\ker A^{*}\cap D(t_{A_{\Lambda}})=\ker A^{*}\cap D(t_{A_{\Lambda}})\cap D(A^{*}).$
By Lemma D.6, we have
$D(t_{A_{\Lambda}})=D(A_{\Lambda})+D(t_{A}).$
Therefore,
$\ker A^{*}\cap D(t_{A_{\Lambda}})\cap D(A^{*})$ $=\ker A^{*}\cap(D(A_{\Lambda})+D(t_{A}))\cap D(A^{*})$
$=\ker A^{*}\cap(D(A_{\Lambda})+D(A^{*})\cap D(t_{A}))$
$=\ker A^{*}\cap(D(A_{\Lambda})+D(A_{F})).$
Applying Lemma C.9, the result follows. ∎
Lemma D.6.
Let $A\in\operatorname{\mathcal{CF}^{s}}(H)$ and let $\Lambda\in\operatorname{\mathrm{Lag}}(H)$. Then we have
$D(A_{\Lambda})+D(t_{A})=D(t_{A_{\Lambda}})$
Proof.
Since we have $D(A_{\Lambda})<D(t_{A_{\Lambda}})$, and $D(t_{A})<Dt_{A_{\Lambda}}$ we only need to show that
$D(A_{\Lambda})+D(t_{A})\supset D(t_{A_{\Lambda}}).$
We choose $V$ such that $D(A_{\Lambda})=V\oplus D(A)$ and since $D(t_{A})$ is closed in the Hilbert space $D(t_{A_{\Lambda}})$ and $V$ is finite dimensional, then $D(t_{A})+V$ is also closed in $D(t_{A_{\Lambda}})$. Then we have
$D(A_{\Lambda})+D(t_{A})\supset D(t_{A})+V\supset D(A)+V=D(A_{\Lambda})\Rightarrow D(t_{A})+D(A_{\Lambda})\supset\overline{D(A_{\Lambda})}=D(t_{A_{\Lambda}})$
where we denoted by $\overline{\#}$ the topological closure of subspace in $D(t_{A_{\Lambda}})$. ∎
Corollary D.7.
Let $A\in\operatorname{\mathcal{CF}^{s}}(H)$ and let $\Lambda\in\operatorname{\mathrm{Lag}}(H)$. Then we have
$D(A_{\Lambda})\cap D(t_{A})=D(A_{\Lambda})\cap D(A_{F}).$
Proof.
By Lemma C.9, it follows that
$D(A_{\Lambda})\cap D(t_{A})=D(A_{\Lambda})\cap D(A^{*})\cap D(t_{A})=D(A_{\Lambda})\cap D(A_{F}).$
The result then follows. ∎
D.2 Morse indices of forms and of Friedrich extensions
We now compare the Morse index of the operator under Dirichlet boundary conditions with that under a general self-adjoint boundary condition.
Let $A_{\Lambda}$ be a self-adjoint extension of $A$. The Morse index of $A_{\Lambda}$, denoted by $\mathrm{n_{-}}(A_{\Lambda})$, is defined as the number of negative eigenvalues (counted with multiplicities) of the operator $A_{\Lambda}$. Associated with $A_{\Lambda}$ is the closed quadratic form $t_{A_{\Lambda}}$. The Morse index of this quadratic form is defined as the maximal dimension of a subspace of $D(t_{A_{\Lambda}})$ on which $t_{A_{\Lambda}}$ is negative definite.
Lemma D.8.
Let $q$ be a quadratic form on $V$ and we assume that $\dim(V/V_{1})<+\infty$. Then $\mathrm{n_{-}}(q)-\mathrm{n_{-}}(q|_{V_{1}})\leq\dim(V/V_{1})$
Proof.
Let $W$ be the maximum negative subspace of $q$. Then $W\cap V_{1}$ is a negative subspace of $q|_{V_{1}}$. We observe that $\dim W/(W\cap V_{1})=\dim(V_{1}+W)/V_{1}\leq\dim(V/V_{1})\leq+\infty$. This concludes the proof. ∎
Proposition D.9.
Let $A_{1},A_{2}\in\mathcal{Cl}^{sa}(H)$ be two closed and self-adjoint operators. We assume that $k_{i}:=\dim D(A_{i})/(D(A_{1})\cap D(A_{2}))<+\infty$, for $i=1,2$ and
$A_{1}|_{D(A_{1})\cap D(A_{2})}=A_{2}|_{D(A_{1})\cap D(A_{2})}$
We have $|\mathrm{n_{-}}(A_{1})-\mathrm{n_{-}}(A_{2})|\leq k_{1}+k_{2}$
Proof.
Let $q_{i}$ be the quadratic form $\langle A_{i}u,u\rangle$ on $D(A_{i})$, $i=1,2$. By Lemma D.8, we have
$\mathrm{n_{-}}(q_{i})-\mathrm{n_{-}}(q_{i}|_{DA_{1}\cap DA_{2}})\leq k_{i}.$
Since $\mathrm{n_{-}}(A_{i})=\mathrm{n_{-}}(q_{i})$, then the proposition follows. ∎
Lemma D.10.
Let $Q$ be a quadratic form on linear space $V$.Let $G$ be a closed subspace of $V$ and let
[…]
where $b_{Q}$ is the bilinear form induced by $Q$ through the polarization identity and we set
$\ker Q=V^{Q}.$
We assume that $G^{QQ}=G+\ker Q$. Then the following formula holds:
$\mathrm{n_{-}}(Q|_{V})-\mathrm{n_{-}}(Q|_{G})=\mathrm{n_{-}}(Q|_{G^{Q}})+\dim(G\cap G^{Q}+\ker Q)/\ker Q.$
Proof.
We refer the interested reader to [HWY20, Theorem 3.1] for the proof. ∎
Lemma D.11.
Let $A:D(A)<H\to H$ be closed a self-adjoint operator with finite Morse index. Then, we have
$\mathrm{n_{-}}(A)=\mathrm{n_{-}}(t_{A}).$
Proof.
Let $\lambda_{1},\ldots,\lambda_{k}$ be the negative eigenvalues of $A$, and let $W=\bigoplus_{i=1}^{k}\ker(A-\lambda_{i}I)$. Then $t_{A}|_{W}$ is a negative definite quadratic form. Therefore, we have
$\mathrm{n_{-}}(t_{A})\geq\dim W=\mathrm{n_{-}}(A).$
Now, $D(t_{A})$ is a Hilbert space with inner product given by $t_{A}(\cdot,\cdot)+c\langle\cdot,\cdot\rangle_{H}$, for some sufficiently large constant $c>0$. For brevity, we denote $E:=D(t_{A})$, and let $\|\cdot\|_{E}$ be the norm induced by this inner product. Then $t_{A}$ defines a continuous inner product on $E$.
Let $V$ be an $n$-dimensional negative subspace for $t_{A}$, and let $\{v_{1},\ldots,v_{n}\}$ be a basis of $V$. Since $D(A)$ is dense in $E=D(t_{A})$, for every $\delta>0$, there exist elements $u_{1},\ldots,u_{n}\in D(A)$ such that $\|v_{i}-u_{i}\|_{E}<\delta$ for all $i$.
Because $t_{A}|_{V}$ is negative definite, the matrix $(t_{A}(v_{i},v_{j}))$ is negative definite. Moreover, since $t_{A}$ is continuous on $E$, the matrix $(t_{A}(u_{i},u_{j}))$ is also negative definite for $\delta$ sufficiently small.
Since $t_{A}(v_{i},v_{j})=\langle Av_{i},v_{j}\rangle_{H}$, by the Min–Max Theorem for self-adjoint operators, it follows that the $n$-th eigenvalue of $A$ satisfies
$\lambda_{n}\leq\max\left\{\langle Au,u\rangle_{H}\,\middle|\,u\in\mathrm{span}\{u_{1},\ldots,u_{n}\},\ \|u\|_{H}=1\right\}<0.$
Thus, we conclude that $n\leq k$, and hence $\mathrm{n_{-}}(t_{A})\leq\mathrm{n_{-}}(A)$. Combining both inequalities, we obtain $\mathrm{n_{-}}(t_{A})=\mathrm{n_{-}}(A)$, completing the proof. ∎
We are now in a position to compute the Morse index
$\mathrm{n_{-}}\left(t_{A_{\Lambda}}\big|_{(D(t_{A}))^{t_{A_{\Lambda}}}}\right).$
Lemma D.12.
Given the decomposition $D(A^{*})=D(A)\oplus U$ Let $p$ be the canonical projection to U. We have
$\mathrm{n_{-}}\left(t_{A_{\Lambda}}\big|_{(D(t_{A}))^{t_{A_{\Lambda}}}}\right)=\mathrm{n_{+}}\big[Q(p(\ker A^{*}),p(D(A_{\Lambda}));p(D(A_{F}))\big].$
Proof.
[…]
$\mathrm{n_{-}}\left(t_{A_{\Lambda}}\big|_{(D(t_{A}))^{t_{A_{\Lambda}}}}\right)=\mathrm{n_{+}}\big[Q(\ker A^{*},\Lambda;F)\big].$
By Lemma D.5,
$(D(t_{A})^{t_{A_{\Lambda}}})=\ker A^{*}\cap(D(A_{\Lambda})+D{A_{F}})=\ker A^{*}\cap U\cap(D(A_{\Lambda})+D(A_{F}))=\ker A^{*}\cap(\Lambda+F).$
Let $z_{1},z_{2}\in\ker A^{*}\cap(\Lambda+F)$, and write $z_{i}=p_{i}+q_{i}$ with $p_{i}\in\Lambda$ and $q_{i}\in F$.
Then:
$t_{A_{\Lambda}}(z_{1},z_{2})=t_{A_{\Lambda}}(p_{1},p_{2})+t_{A_{\Lambda}}(p_{1},q_{2})+t_{A_{\Lambda}}(q_{1},p_{2})+t_{A_{\Lambda}}(q_{1},q_{2}).$
Since $A_{\Lambda}$ is the Friedrich extension of itself, we have:
$t_{A_{\Lambda}}(u,v)=\langle A_{\Lambda}u,v\rangle_{H},\quad\forall\,u\in D(A_{\Lambda}),\ v\in D(t_{A_{\Lambda}}).$
Moreover, since $t_{A_{\Lambda}}|_{D(t_{A})}=t_{A}$, it follows that:
$t_{A_{\Lambda}}(u,v)=t_{A}(u,v)=\langle A_{F}u,v\rangle_{H},\quad\forall\,u,v\in D(A_{F}).$
Therefore:
$t_{A_{\Lambda}}(z_{1},z_{2})$ $=\langle A_{\Lambda}p_{1},p_{2}\rangle+\langle A_{\Lambda}p_{1},q_{2}\rangle+\langle q_{1},A_{\Lambda}p_{2}\rangle+\langle A_{F}q_{1},q_{2}\rangle$
$=\langle A^{*}p_{1},p_{2}\rangle+\langle A^{*}p_{1},q_{2}\rangle+\langle q_{1},A^{*}p_{2}\rangle+\langle A^{*}q_{1},q_{2}\rangle$
$=\langle A^{*}(p_{1}+q_{1}),p_{2}+q_{2}\rangle-\langle p_{1},A^{*}q_{2}\rangle+\langle A^{*}p_{1},q_{2}\rangle.$
Since $z_{1}=p_{1}+q_{1}\in\ker A^{*}$, we have $A^{*}(p_{1}+q_{1})=0$. Thus,
$t_{A_{\Lambda}}(z_{1},z_{2})=-\langle p_{1},A^{*}q_{2}\rangle+\langle A^{*}p_{1},q_{2}\rangle=-\omega(p_{1},q_{2}).$
Thus, we conclude that
$\mathrm{n_{-}}\left(t_{A_{\Lambda}}\big|_{(D(t_{A}))^{t_{A_{\Lambda}}}}\right)=\mathrm{n_{+}}\big[Q(\ker A^{*},\Lambda;F)\big],$
as claimed. ∎
Lemma D.13.
Let $\Lambda=\Lambda_{0}\oplus\Lambda_{D}$ and let $(A_{s})_{s\in(a,b)}\in\operatorname{\mathcal{CF}^{s}}(H)$. For each $r,s\in(a,b)$ and $r<s$, we assume that $D(A_{s})<D(A_{r})$ and that
$\bigcap_{s>a)}D(t_{A_{s,\Lambda}})=(0).$
Then the following monotonicity property on the Morse index holds:
$\mathrm{n_{-}}(A_{r,\Lambda})\geq\mathrm{n_{-}}(A_{s,\Lambda}).$
Moreover, there exists $s>a$ such that
$\mathrm{n_{-}}(A_{s,\Lambda})=0.$
Proof.
We observe that
$D({A_{\Lambda}})+D(t_{L})=D(t_{A_{\Lambda}}),$
where $A$ is the minimal operator and $A_{\Lambda}$ is the self-adjoint extension of $A$ with boundary condition $\Lambda$. Then, we have $D(t_{A_{s,\Lambda}})<D(t_{A_{r,\Lambda}})$ with $s\leq r$. We also have
$\bigcap_{s>a}D(t_{A_{s,\Lambda}})=\set{0}$.
Let $\alpha=t_{A_{b,\Lambda}}$ and $V=D(\alpha)$. Then $\alpha(u,v)=\langle A_{b}u,v\rangle_{V}$. We assume that $Z<V$ is the negative spectral subspace of $\alpha$, let $V_{n}=D(t_{A_{a+\frac{1}{n},\Lambda}})$ and let $K_{n}=V_{n}^{\alpha}$ where
[…]
Since $A$ is a bounded Fredholm operator, then $A(V_{n})$ is a closed subspace of $V$. Moreover, we have $L(V_{1})\supset L(V_{2})\cdots\supset L(V_{n})\cdots$ and so
$\bigcap_{n=1}^{+\infty}A(V_{n})=(0).$ (D.1)
Let $T_{n}$ be the projection onto $K_{n}$ and we observe that $T_{n}$ strongly converge to $I$ because of Equation (D.1). Let $\set{w_{i}}$ be a basis of $Z$. Then $\lim_{n\to+\infty}T_{n}w_{i}=w_{i}$ . Since $Z$ is the negative spectral subspace of $\alpha$, then the matrix $A]_{ij}=\alpha(w_{i},w_{j})$ is negative definite. So, for $n$ large enough, the matrix
$A_{n}]_{ij}=\alpha(T_{n}w_{i},T_{n}w_{j})$
is also negative definite. So $Z_{n}:=T_{n}(V)$ is the maximal negative subspace of $V$. Since $Z_{n}=T_{n}(H)<K_{n}$, we have $Z_{n}\perp A(V_{n})$ and by this we get that
[…]
So, $Z_{n}\cap V_{n}=\ker A_{a+1/n,\Lambda}$. By taking into account Lemma 5.7, for $n$ sufficiently large, we get that $\ker A_{a+1/n,\Lambda}=(0)$ and by this we conclude that $Z_{n}\cap V_{n}=(0)$ for $n$ large enough and $\alpha(Z_{n},V_{n})=0$. Since $Z_{n}$ is the maximal negative subspace, $V_{n}$ is a non-negative subspace. Then we can conclude that there is $s>a$.
$\mathrm{n_{-}}(A_{s})=0.$
This concludes the proof. ∎
Lemma D.14.
Let $H$ be a Hilbert space, let $(V_{n})_{n\in{\mathbb{N}}}$ be a sequence of closed (nested) subspaces such that
$V_{k+1}<V_{k}\qquad\textrm{ and }\qquad\bigcap_{n\geq 1}V_{n}=(0)$
and let $T\in\operatorname{\mathcal{BF}}(H)$ be a bounded Fredholm operator. Then for every $n\geq 1$ the subspace $T(V_{n})$ is closed and
$\bigcap_{n\geq 1}T(V_{n})=(0).$
Proof.
Since $T$ is a bounded and Fredholm operator, then $T:(\ker T)^{\perp}\to\operatorname{im}T$ is a linear isomorphism. Let $P$ be the orthogonal projection onto $(\ker T)^{\perp}$ and we observe that the following holds: $T=T|_{(\ker T)^{\perp}}P:H\to\operatorname{im}T$. Since $T|_{(\ker T)^{\perp}}$ is an isomorphism for proving the result, we only need to show that $\bigcap_{n\geq 1}P(V_{n})=\set{0}$ and that $P(V_{n})$ is closed.
Let $W=\ker T=\ker P$. Since $T$ is a Fredholm operator, then $\dim W<+\infty$. Let $Pu\in\bigcap_{n\geq 1}P(V_{n})$. Then for every $n\geq 1$, we get that $u\in V_{n}+W$. Moreover $\bigcap_{n\geq 1}(V_{n}\cap W)<\bigcap_{n\geq 1}V_{n}=(0)$. Now, since $W$ is finite dimensional, there exists $m\in{\mathbb{N}}$ such that $V_{n}\cap W=(0)$ for each $n>m$. Without loss of generality, we can assume that $V_{n}\cap W=(0)$ for $n\geq 1$. Since $V_{n+1}<V_{n}$, the decomposition of $u$ in $V_{n}+W$ is unique. This in particular implies that $u\in\bigcap_{n\geq 1}V_{n}\oplus W=W$. By this we get that $Pu=0$ and and thus $\bigcap_{n\geq 1}P(V_{n})=(0)$.
Now, since $V_{n}$ is closed and $W$ is finite dimensional, $V_{n}+W$ is closed and $P(V_{n})=(V_{n}+W)\cap W^{\perp}$ is closed, too. This concludes the proof. ∎
Lemma D.15.
Let $(V_{n})_{n\in{\mathbb{N}}}$ be a nested sequence of closed subspaces of the of Hilbert space $H$ and we assume that $\bigcap_{n\geq 1}V_{n}=(0)$. Let $B\in\operatorname{\mathcal{BF}^{sa}}(H)$ such that $\mathrm{n_{-}}(B)<\infty$ and let $q(\cdot)=\langle B\cdot,\cdot\rangle$ be the corresponding quadratic form. Then there exists $m>0$ such that
$\mathrm{n_{-}}(q|_{V_{m}})=0.$
Proof.
Let $u\in(B(V_{n}))^{\perp}$ and $v\in V_{n}$. Then we have
$b(u,v)=\langle Bu,v\rangle=\langle u,Bv\rangle=0$
where $b$ is the bilinear form associated to $q$ through the polarization identity. This, in particular, implies that for every $n\geq 1$, the following inequality holds
$\mathrm{n_{-}}(q)\geq\mathrm{n_{-}}(q|_{V_{n}+(B(V_{n}))^{\perp}})\geq\mathrm{n_{-}}(q|_{V_{n}})+\mathrm{n_{-}}(q|_{(B(V_{n}))^{\perp}}).$ (D.2)
For concluding the proof we only need to prove that there exists $m\in{\mathbb{N}}$ such that $\mathrm{n_{-}}(q)\leq\mathrm{n_{-}}(q|_{(B(V_{m}))^{\perp}})$.
Let $P_{n}$ be the orthogonal projection onto $(B(V_{n}))^{\perp}$. By Lemma D.14, we get that
$\lim_{n\to+\infty}(I-P_{n})x=0\qquad\textrm{ for each }\quad x\in H.$
We denote by $W=\operatorname{span}\{x_{1},\cdots,x_{k}\}$ be the negative spectral space of $q$ and we observe that $\lim_{n\to+\infty}b(P_{n}x_{i},P_{n}x_{j})=b(x_{i},x_{j})$. So, there exists $m>0$ such that the matrix $A_{ij}:=b(P_{m}x_{i},x_{j})$ is negative definite and by this we get that
$\mathrm{n_{-}}(q|_{(B(V_{m}))^{\perp}})\geq\mathrm{n_{-}}(q).$ (D.3)
By adding the inequalities provided at Equation (D.2) and Equation (D.3), the conclusion readily follows. ∎
Example D.16.
We assume that the SL-operator $L:D(L^{*})<L^{2}((a,b),{\mathbb{R}}^{n})\to L^{2}((a,b),{\mathbb{R}}^{n})$ is regular and that $P$ is positive definite. In particular under this assumptions the associated maximal operator $L^{*}$ is bounded from below. Given the following decomposition $D(L^{*})=D(L)\oplus U$, let $\Lambda$ be any Lagrangian subspace of $U,\rho)$ and let $\Lambda_{D}<U$ be the Dirichlet Lagrangian. It follows that
$D(L^{*})=W^{2,2}([a,b],{\mathbb{R}}^{n})\qquad\textrm{ and }\qquad D(L_{\Lambda_{D}})=W_{0}^{2,2}([a,b],{\mathbb{R}}^{n}).$
Let $E$ be an isotropic subspace of $(U,\rho)$ and we observe that for every $u,v\in D(L_{E})$, the bilinear form defined by $\langle L_{E}u,v\rangle_{L^{2}}$ is bounded and symmetric. By taking the closure of $D(L_{E})$ in $W^{1,2}([a,b],{\mathbb{R}}^{n})$, we extend this symmetric bilinear form and we denote it by $t_{L_{E}}$. So, in particular $t_{L_{E}}(\cdot,\cdot)$ is continuous on its domain denoted by $D(t_{L_{E}})$. Moreover, we have
$D(t_{L_{D}})=W_{0}^{1,2}([a,b],{\mathbb{R}}^{n})\qquad\textrm{ and }D(t_{L_{E}})=D(t_{L_{D}})+E.$
Lemma D.17.
Assume that $\Lambda$ is Lagrangian. We have $\ker t_{L_{\Lambda}}=\ker L_{\Lambda}$.
Proof.
$t_{L_{\Lambda}}(u,v)=0$ for all $v\in D(t_{L_{\Lambda}})$ is equivalent to $t_{L_{\Lambda}}(u,v)=0$ for all $v\in D(L_{\Lambda})$ since $D(L_{\Lambda})$ is dense in $D(t_{L_{\Lambda}})$. Similarly, it also equivalent to $\langle u,L_{\Lambda}v\rangle_{L^{2}}=0$. So, we have
$\ker t_{L_{\Lambda}}=D(t_{L_{\Lambda}})\cap(\operatorname{im}L_{\lambda})^{\perp}=(W_{0}^{1,2}((a,b),{\mathbb{R}}^{n})+\Lambda)\cap\ker L_{\Lambda}=\ker L_{\Lambda}.$
Lemma D.18.
$\mathrm{n_{-}}(t_{L_{\Lambda}})=\mathrm{n_{-}}(L_{\Lambda})$
Proof.
For each eigenvector $u$ of $L_{\Lambda}$ corresponding to a negative eigenvalue, we get that
$t_{L_{\Lambda}}(u,u)=(L_{\Lambda}u,u)<0.$
By this, it follows that $\mathrm{n_{-}}(L_{\Lambda})\leq\mathrm{n_{-}}(t_{L_{\Lambda}})$.
Conversely, let $V$ be a finite dimensional subspace of $D(t_{L_{\Lambda}})$ where the form $t_{L_{\Lambda}}$ is negative definite and let $(v_{i})_{i}$ be its basis. For each $\varepsilon>0$, there exists $(w_{i})_{i}\subset D(L_{\Lambda})$ such that $\|v_{i}-w_{i}\|_{W^{1,2}}<\varepsilon$. For $\varepsilon$ small enough, $t_{L_{\Lambda}}|_{\operatorname{span}\set{w_{i}}}=(L_{\Lambda}w_{i},w_{i})$ is negative definite. By this it follows that $\mathrm{n_{-}}(L_{\Lambda})\geq\mathrm{n_{-}}(t_{L_{\Lambda}})$ concluding the proof. ∎
Lemma D.19.
Let $H$ be a Hilbert space, $B\in\operatorname{\mathcal{BF}^{sa}}(H)$ and let $Q$ be a quadratic form defined by $Q(u)=\langle Bu,u\rangle$. Let $V$ be a closed subspace of $H$. Then we have
$V^{QQ}=V+\ker Q.$
Proof.
[…]
$V^{QQ}=B^{-1}((B(V))^{\perp\perp})=B^{-1}(\overline{B(V)}).$
Since $B$ is a Fredholm operator, then $B(V)$ is closed. Then we have
$V^{QQ}=B^{-1}(B(V))=V+\ker B.$